### v2

# Inheritune: Training Smaller Yet More Attentive Language Models

Abstract
Large Language Models (LLMs) have achieved remarkable performance across various natural language processing tasks, primarily due to the transformer architecture and its self-attention mechanism. However, we observe that in standard decoder-style LLMs, attention matrices degenerate to single-column for deeper layers. Layers in this state are unable to learn anything meaningful and mostly redundant; we refer to these as lazy layers. The goal of this paper is to train smaller models by eliminating this structural inefficiency without compromising performance.
Motivated by this observation, we propose Inheritune, a simple yet effective training recipe for developing smaller, high-performing language models. Smaller models trained with $\mathsf{Inheritune}$ inherit early transformer layers from a larger pre-trained model, then retrain and progressively expand until they match or exceed the performance of the larger model. We demonstrate that $\mathsf{Inheritune}$ enables the training of various sizes of GPT-2 models on datasets like OpenWebText-9B and FineWeb_Edu. Models trained with $\mathsf{Inheritune}$, despite having significantly fewer layers, match or even surpass the performance of their larger counterparts. For instance, our 16-layer GPT-2 medium variant achieves comparable performance to the standard 24-layer GPT-2 medium model. Code is available at https://github.com/sanyalsunny111/LLM-Inheritune.
1 Introduction
Large Language Models (LLMs) are built with decoder-style transformer blocks (Vaswani et al., 2017). These models are typically designed to be large, with a significant portion of their parameters dedicated to their depth, with multiple transformer blocks stacked together, building model capacity. Each block or layer in the stack refines the representations learned by the previous blocks, allowing the model to develop a nuanced understanding of the input data. As these models scale in depth and size, their performance tends to improve Kaplan et al. (2020); Hoffmann et al. (2022), benefiting from increased model capacity.
The causal self-attention (hereafter referred to as attention) mechanism is arguably the most crucial component of a transformer block. It allows models to combine tokens as a weighted linear sum of their attention scores, effectively capturing long-range dependencies and contextual relationships within text data. However, as models grow in depth, they often encounter a phenomenon known as attention degeneration, caused by a collapse in the attention rank Noci et al. (2022); Dong et al. (2021); He et al. (2023). Notably, this phenomenon has not been studied in the context of standard LLMs. A formal discussion on attention degeneration is provided in Section 2.
In this paper, we empirically analyze 24-layer GPT-2 medium and 36-layer GPT-2 large models (decoder-style LLMs) Radford et al. (2019) for attention degeneration and observe that many deeper layers in both models exhibit rank-1 attention matrices. Further investigation reveals that most of these rank-1 matrices are also single-column, i.e., their mass is concentrated in a single column. Our attention matrix analysis is presented in Figure 1. We term these deeper layers, where all attention matrices of a given layer are degenerate, as lazy layers.
Motivated by this new finding, we aim to develop performant small base language models (LMs) that utilize weights from inefficient larger base LMs. A base LM is a decoder-style model trained solely for next-token prediction without additional enhancements like instruction tuning or reinforcement learning with human feedback (RLHF). Our proposal is straightforward, we start by initializing our smaller LM (target) using the first few blocks from a large pre-trained LM (reference). We then train the target model for a specified number of steps. After this initial training, we incrementally grow the target model by adding more blocks, continuing the training process until it matches or surpasses the pre-train validation loss (also val loss) of the reference model. During the growth phase, the newly added blocks can be initialized with lazy layers of the reference LM. We refer to this simple, yet effective, training approach as $\mathsf{Inheritune}$.
(a) Rank analysis of vanilla GPT-2 Medium
(b) Matrix mass analysis of vanilla GPT-2 Medium
(c) Model initialization with later layers
(d) Rank analysis of vanilla GPT-2 Large
(e) Matrix mass analysis of vanilla GPT-2 Large
(f) Model initialization with later layers
Figure 1: Attention matrices of many deeper layers often degenerate to single-column matrices in regular decoder style LLMs, layers with fully degenerate attention fail to learn meaningful representations. We computed a single attention matrix with 100 tokens from the OpenWebText validation set with 4M tokens. Next we performed 100 runs and plotted the mean and std of the max rank and mass as a function of layers for our rank and mass analysis respectively. Figure (a and d): An analysis of a 24-layer GPT-2 medium and a 36-layer GPT-2 large shows the max rank of the attention matrices across all layers. Figure (b and e): A closer look at the same GPT-2 models also reveals that the dominant mass proportion of several attention matrices is concentrated in a single-column, particularly in deeper layers. Figure (c and f): When initializing 12-layer and 18-layer variants of the vanilla GPT-2 medium and GPT-2 large models with deeper layers (Lazy layers) exhibiting degenerated attention–their performance is comparable to models with random initialization. However, initializing models with early layers leads to significantly better generalization and convergence.
In summary, our key contributions are as follows:
  1. 1.
    Analysis of Attention Degeneration Leading to Lazy Layers. We empirically investigate attention degeneration in standard LLM settings. Our analysis shows that rank-collapsed attention matrices often exhibit single-column structures, revealing a significant structural inefficiency in the attention mechanism of standard LLMs in deeper layers (see Figure 1 and Figure 2).
  2. 2.
    Introduction of Inheritune. We propose $\mathsf{Inheritune}$ as an approach to effectively train high-performing, smaller models. This method involves inheriting a few early blocks from a larger pre-trained model and progressively growing and training the smaller model. The initialization is entirely zero-shot. We validate the effectiveness of $\mathsf{Inheritune}$ through comprehensive experiments using GPT-2 xlarge (1.5B), GPT-2 large (770M), and GPT-2 medium (355M) models, trained on the OpenWebText dataset with 9B tokens (with data repetition) and the FineWeb_edu dataset with 100B tokens (without data repetition).
  3. 3.
    Evaluation Against Multiple Baselines. Models derived using $\mathsf{Inheritune}$ consistently outperform various baselines, including much larger models trained from scratch (refer Table 1), model initialization and efficient training baselines (refer Table 2), and models trained using two knowledge distillation techniques (refer Figure 3). In settings where training tokens are not repeated, we observe similar trends (refer Figure 4).
L4 H2
L7 H3
L20 H2
L22 H4
24 layer GPT2 medium model trained from scratch.
L4 H2
L7 H3
L11 H2
L15 H1
Our 16-layer GPT2 medium model trained using $\mathsf{Inheritune}$.
Figure 2: Inheritune preserves effective attention patterns in smaller models. Comparison of attention patterns across layers (L) and heads (H) in two GPT2-medium models: (top) 24-layer model trained from scratch, (bottom) 16-layer model trained with $\mathsf{Inheritune}$. Attention maps are averaged over three randomly selected strings, with 40 tokens each from the validation. Darker colors indicate higher attention scores. $\mathsf{Inheritune}$ maintains focused attention even in deeper layers, contrasting with the uniform patterns in the standard model’s later layers.
2 Attention Degeneration in Standard Decoder-Style LLMs
Preliminaries:
A vanilla transformer-based model consists of $L$ transformer blocks (layers). The model operates on an input sequence $X\in\mathbb{R}^{T\times d}$, where $T$ denotes the sequence length (number of tokens), and $d$ represents the embedding dimension or hidden size of the model. The output of each layer $l$ is denoted as $X^{(l)}\in\mathbb{R}^{T\times d}$.
Each transformer block primarily consists of two sub-layers: a self-attention block and a position-wise feed-forward network (FFN). The self-attention mechanism enables the model to weigh the relevance of different tokens in the sequence relative to each other. Specifically, for a single attention head, the attention computation is defined as
$\text{Attention}(Q,K,V)=\underbrace{\text{softmax}\left(\frac{QK^{\top}}{\sqrt{d_{k}}}\right)}_{\textbf{Attention matrix: $A(X)$}}V$ (1)
where the queries $Q=XW_{Q}$,  keys $K=XW_{K}$,  and values $V=XW_{V}$ are linear transformations of the input $X$. Here, $W_{Q},W_{K}\in\mathbb{R}^{d\times d_{k}}$ and $W_{V}\in\mathbb{R}^{d\times d_{v}}$ are the weight matrices for the queries, keys, and values, respectively. Typically, $d_{k}=d_{v}=\frac{d}{h}$, where $h$ is the number of attention heads. In this single-head scenario, we set $d_{k}=d_{v}=d$.
The attention matrix $A(X)\in\mathbb{R}^{T\times T}$ captures the pairwise attention scores between all token positions in the sequence. The softmax is applied row-wise. The attention matrix $A(X)$ is then used to compute a weighted sum of the value vectors.
Previous research by Dong et al. (2021) and He et al. (2023) has shown that in self-attention networks (SANs) without residual connections and feed-forward networks (FFNs), the rank of an attention matrix converges to rank-1 doubly exponentially with respect to the depth of the model. This phenomenon, known as the rank collapse of attention matrices, results in a loss of expressive power as the attention mechanism attends to all tokens uniformly. Noci et al. (2022) showed that even with residual connections (without layernorm) attention matrices can still lose rank in deeper layers if the residual connections are not scaled by $1/\sqrt{L}$. Interestingly, they also linked the rank collapse to vanishing gradients of the keys and queries in deeper layers, which affects the overall trainability of the transformer-based models. However, these findings do not directly apply to the standard LLMs, as transformer blocks in these models include residual connections, layer norms, and FFNs, which are expected to mitigate both rank collapse and the vanishing gradient problem.
Approximate Rank Computation of Attention Matrices
In this paper, we deeply analyzed the structure of attention matrices to diagnose the presence of rank collapse or similar phenomena in standard transformer-based LLMs using GPT-2 models. For our first analysis, we compute the approximate rank (referred to as rank hereafter) of $A(X)$ for all attention heads within each layer. Formally, we began by computing the Singular Value Decomposition (SVD) of $A(X)=U\Sigma V^{\top}$, where the diagonal entries $\{\sigma_{i}\}_{i=1}^{T}$ of $\Sigma$ represent the singular values, quantifying the variance captured by each corresponding singular vector of $A(X)$. To determine the minimal number of singular values required to capture 90% of the total variance, we solved: $k^{*}=\min\left\{k\in\{1,2,\ldots,T\}\mid\frac{\sum_{i=1}^{k}\sigma_{i}^{2}}{\sum_{j=1}^{T}\sigma_{j}^{2}}\geq 0.90\right\}$. Here, $k^{*}$ is the approximate rank of $A(X)$ computed using the explained variance method.
Dominant Single-Column Structure in Attention Matrices
We further investigated the dominant structure of these rank-1 attention matrices and observed that, on average, many of these matrices have their mass concentrated in a single column. This intrinsic structure can be viewed as a special case of rank-1 attention matrices. To quantify this, we computed the proportion of the matrix mass contributed by each column $j$ of $A(X)$ by computing $\frac{\|A_{\cdot,j}\|_{2}^{2}}{\|A(X)\|_{F}^{2}}$, where $A_{\cdot,j}$ denotes the $j$-th column of $A(X)$, $\|A_{\cdot,j}\|_{2}$ is the $\ell_{2}$-norm of that column, and $\|A(X)\|_{F}$ is the Frobenius norm of $A(X)$. Next, to determine the minimal number of columns required to capture 90% of the total mass, we solved; $m^{*}=\min\left\{m\in\{1,2,\ldots,T\}\mid\sum_{j=1}^{m}\frac{\|A_{\cdot,j}\|_{2}^{2}}{\|A(X)\|_{F}^{2}}\geq 0.90\right\}$. Here, $m^{*}$ represents the minimal number of columns needed for the cumulative column mass ratios to reach or exceed 90%.
Degeneration of Attention Matrices in GPT-2 Models
In Figure 1, we present the layer-wise analysis of rank and mass. For this analysis, we computed $A(X)$ using 100 randomly selected samples from the validation set of OpenWebText valset with 4M tokens, each with a sequence length of $T=100$ tokens, across all attention heads within each layer. Next, we computed the rank with a 90% variance threshold, and for every layer, we chose the maximum rank across all the heads. In Figure 1(a) and 1(d) we plotted maximum rank as a function of layers for 100 runs with mean and standard deviation (std); it is quite evident that many deeper layers exhibit all rank-1 attention matrices. We highlight that this rank collapse occurs for both the GPT-2 medium and GPT-2 Large models with skip connections and FFNs, extending the findings of Dong et al. (2021) and Noci et al. (2022) to standard LLM architectures. Next, based on our mass analysis, we demonstrate that most of the rank collapsed attention matrices are also single-column matrices as depicted in Figure 1(b) and Figure 1(e). We follow the analysis protocol as discussed in the rank analysis. This degeneration of attention matrices in deeper layers provides quantitative evidence for the existence of lazy layers. Specifically, we observe that some deeper layers exhibit complete degeneration of all attention matrices across all attention heads, indicating reduced performance and less effective token mixing.
How much transferable knowledge do lazy layers hold compared to their early counterparts? To answer this question, we initialized a 12-layer GPT-2 medium variant [^1] A variant shares the same configurations as the parent model but has fewer layers. and an 18-layer variant of GPT-2 large using lazy layers extracted from pre-trained 24-layer GPT-2 medium and 36-layer GPT-2 large models. These pre-trained models are trained on the OpenWebText-9B dataset for 100K steps. We then fine-tuned these GPT-2 variants on the same dataset for an additional 10K steps. For comparison, we conducted two baseline experiments in which the GPT-2 variants were initialized either with the first few transformer blocks or with random initialization. As shown in Figures 1(c) and 1(f), models initialized with lazy layers demonstrate poor transferability, performing similarly to models with random initialization. This provides additional evidence that lazy layers with fully degenerate attention fail to learn meaningful representations.
2.1 Attention Pattern Visualization
To provide further evidence of lazy layers and provide a preview of our solution, we visualized attention patterns across various layers of a vanilla 24-layer GPT-2 medium model. Fig. 2 shows the attention patterns for both a vanilla 24-layer model trained from scratch and a 16-layer model trained using our proposed method, $\mathsf{Inheritune}$. We visualized full attention and not causal attention for the sake of better visualization. In practice, GPT-2 models compute causal attention. We computed these attention matrices using randomly selected strings from the val set of OpenWebText and took 40 tokens averaged over three runs.
Algorithm 1 Inheritune: Training Recipe for Small Language Models
1:Reference model $\mathcal{M}_{\text{ref}}$ with $k$ layers, datasets $\mathcal{D}_{\text{train}}$ and $\mathcal{D}_{\text{val}}$, steps $\mathsf{T}$
2:Initialize $\mathcal{M}_{\text{tgt}}$ with first $n=k/2$ layers from $\mathcal{M}_{\text{ref}}$
3:Train $\mathcal{M}_{\text{tgt}}$ on $\mathcal{D}_{\text{train}}$ for $\mathsf{T}$ steps
4:while $\mathcal{M}_{\text{tgt}}$ performance < $\mathcal{M}_{\text{ref}}$ performance on $\mathcal{D}_{\text{val}}$ do
5:     Grow $\mathcal{M}_{\text{tgt}}$ by inheriting additional layers
6:     Train $\mathcal{M}_{\text{tgt}}$ for $\mathsf{T}$ steps
7:end while
8:return Optimized model $\mathcal{M}_{\text{tgt}}$
In the 24-layer model trained from scratch (top row of Fig. 2), we observe a clear progression in attention patterns. The early layers (L4 and L7) exhibit structured patterns with a mix of local and global attention Gong et al. (2019); Beltagy et al. (2020); Chen et al. (2021). In contrast, the deeper layers (L20 and L22) display more uniform patterns, indicating a loss of focus. This uniformity is a hallmark of lazy layers, where the attention mechanism loses its ability to selectively focus on specific relevant tokens. In contrast, our 16-layer model trained with $\mathsf{Inheritune}$ (bottom row) demonstrates more focused and effective attention patterns, even in its later layers (L11 and L15). This striking difference suggests that our method makes the model more attentive and addresses attention degeneration, potentially leading to more efficient models in a compact size (also refer to Figure 9). We will discuss these results in more detail after introducing our method, but this preview underscores the promise of our approach.
3 Inheritune: Our Proposed Training Recipe
This section offers a detailed description of our method, key implementation considerations, and how it addresses the inefficiencies present in current architectures.
Recall that we have established the issue of attention degeneration with two motivating examples, highlighting specific inefficiencies in pre-trained LLMs. In this paper, we turn this challenge into an opportunity to create smaller base LMs that are equally performant, achieving similar or lower validation loss compared to their larger, less efficient counterparts, which we refer to as reference models. Our proposed solution builds on two key insights: (1) the early layers of deep LLMs provide effective model initialization, and (2) multiple lazy layers can be collapsed into fewer layers and re-trained to improve the model capacity.
Setup:
We split the dataset into a training set $\mathcal{D}_{\text{train}}$ and a validation subset $\mathcal{D}_{\text{val}}$. Next, we assume that there exists a pre-trained reference model $\mathcal{M}_{\text{ref}}$, comprising $k$ layers, represented by $\mathsf{W}_{\text{ref}}=\{\mathsf{W}_{0},\mathsf{W}_{1},\ldots,\mathsf{W}_{k-1}\}$ trained with $\mathcal{D}_{\text{train}}$ for $\mathsf{T}$ steps. We want to train a smaller model $\mathcal{M}_{\text{tgt}}$ with the same or better validation loss (lower is better) compared to its larger counterpart $\mathcal{M}_{\text{ref}}$.
We now present $\mathsf{Inheritune}$, our proposed training recipe for efficiently developing small base language models (LMs). $\mathsf{Inheritune}$ operates on the principle of zero-shot initialization and progressive growth. The $\mathsf{Inheritune}$ process consists of three main steps, which we present below and formalize in Algorithm 1:
  1. 1.
    Inherit: Initialize $\mathcal{M}_{\text{tgt}}$ with the first $n=k/2$ layers of $\mathcal{M}_{\text{ref}}$, including weights, prediction head, and token embeddings.
  2. 2.
    Train: Train $\mathcal{M}_{\text{tgt}}$ for $\mathsf{T}$ steps on $\mathcal{D}_{\text{train}}$ and evaluate on $\mathcal{D}_{\text{val}}$.
  3. 3.
    Grow: If needed, increase $\mathcal{M}_{\text{tgt}}$’s size and repeat steps 1-2 until the desired performance is achieved.
With our method now formally described, we turn to empirical validation. In the following sections, we present comprehensive results demonstrating $\mathsf{Inheritune}$’s effectiveness across various scenarios, including different model sizes and data regimes. In addition, we conducted an in-depth ablation study to analyze the impact of initialization on performance, providing information on the adaptability of our approach.
4 Experiments
We evaluated $\mathsf{Inheritune}$ through a series of comprehensive experiments using GPT-2 xlarge (1.5B), GPT-2 large (770M) and GPT-2 medium (355M) models, Radford et al. (2019) pre-trained on the 9B tokens OpenWebText dataset (Gokaslan & Cohen, 2019). These models are trained with data repetition, meaning data is randomly sampled with replacement during batch creation. This experimental setup is adapted from Liu et al. (2023); Sanyal et al. (2024). For evaluation, we compare model(s) trained with $\mathsf{Inheritune}$ with baselines from three key settings: a) baseline models trained from scratch with random initialization, b) baseline models trained using various zero-shot initialization techniques, and c) baseline models trained with knowledge distillation. Table 10 provides detailed specifications for all models used in our experiments. Finally, we conduct a thorough ablation study of our initialization strategy, focusing on the 16-layer GPT-2 medium variants.
We provide experimental details of our $\mathsf{Inheritune}$ training recipe using a GPT-2 large model as an example; a similar procedure was applied to train other models. Our methodology for applying $\mathsf{Inheritune}$involves the following steps:
  1. 1.
    Reference Model: We train the full 36-layer GPT-2 Large model on $\mathcal{D}_{\text{train}}$ for 100K steps and evaluate its validation loss ( log-perplexity) on $\mathcal{D}_{\text{val}}$. This establishes our benchmark validation loss.
  2. 2.
    Model initialization We initialize an 18-layer model ($n=k/2$) using the trained 36-layer model as a reference.
  3. 3.
    Training and Evaluation: We train the 18-layer model on $\mathcal{D}_{\text{train}}$ for $T$ steps and evaluate its validation loss.
  4. 4.
    Iterative Refinement: If the smaller model’s performance is inferior, we incrementally increase its size by two layers and repeat steps 2-3 until we achieve parity with the reference model’s validation loss.
Models Layers Initialization Steps Pre-train Downstream (0-shot)
Val loss ($\downarrow$) Wikitext ($\downarrow$) Lambada
GPT-2 Medium 24 rand init 100K 2.81 31.93 36.54
16 rand init 100K 2.86 33.67 34.60
16 rand init 200K 2.83
12 Ours 100K 2.87
14 Ours 100K 2.84
Final Model $\longrightarrow$ 16 Ours 100K 2.81 32.04 35.96
GPT-2 Large 36 rand init 100K 2.85 34.84 34.14
18 rand init 100K 2.97 37.63 30.97
18 rand init 200K 2.84
18 Ours 100K 2.80 35.38 34.64
GPT-2 xLarge 48 rand init 100K 2.65 25.45 39.90
24 rand init 100K 2.69 28.32 38.46
24 rand init 200K 2.62
24 Ours 100K 2.64 25.52 43.30
Table 1: Inheritune achieves superior performance with reduced model size. Comparison of $\mathsf{Inheritune}$-trained models (24-layer GPT-2 xLarge, 18-layer GPT-2 Large, 16-layer GPT-2 Medium) against full-sized counterparts and extended training baselines. Metrics include pre-training validation loss ($\downarrow$), zero-shot Wikitext ($\downarrow$) and Lambada performance. Note: GPT-2 xLarge and Large took one round of training; GPT-2 Medium took three rounds.
Baselines trained from scratch (rand init.) :
We compare our $\mathsf{Inheritune}$-derived model against much larger GPT-2 reference models trained from scratch for the same number of steps and similar-sized models trained from scratch for both the same and double the number of training steps.
Baselines trained with various model initialization and efficient training techniques.
Here, we compare our model derived using $\mathsf{Inheritune}$ to models of similar size trained with various zero-shot model initialization and efficient training techniques such as stacking, hybrid stacking, and half-width initialization. We explain these baseline training recipes using GPT-2 large and its variants as an example and apply the same process to other models.
  1. 1.
    Stacking Gong et al. (2019); J. Reddi et al. (2023) is a model initialization and efficient (stage-wise) training recipe. We trained a 9-layer GPT-2 large variant from scratch for 100K steps, then expanded the model to 18 layers by copying the weights from layers 0-8 to layers 9-17. Finally, we re-trained this new 18-layer GPT-2 large variant, using stacking initialization for an additional 100K steps.
  2. 2.
    Hybrid stacking: Hybrid stacking is stacking that utilizes a large pre-trained reference model for initialization instead of using its own pre-trained weights. We took the weights of layers 0-8 from the reference 36-layer GPT-2 large model and expanded it to an 18-layer model by copying the weights to layers 0-17. We then trained this new 18-layer GPT-2 variant for 100K steps.
  3. 3.
    Half width: We initialized the baseline GPT-2 large variant across the width dimension and preserved the entire depth. We copied the weights of the first half of the attention heads (0-9) and MLPs of the GPT-2 large reference model into the baseline GPT-2 variant with half the width but all layers.
Baselines trained with Knowledge Distillation
As a baseline, we first apply logit-based knowledge distillation (KD) Hinton et al. (2015) to train a 16-layer GPT-2 medium variant (student) initialized randomly. For the second baseline, we use a DistillBERT-style approach Sanh et al. (2019), where the 0-11 layers of the student model are initialized with each alternate block of its teacher, and the remaining 4 blocks are initialized using layers 18, 19, 20, and 21 of the teacher. Both baselines are trained for 50K steps, using a vanilla 24-layer GPT-2 medium model as the teacher (our reference model). Additional results with 12-layer GPT-2 medium variant(s) trained from scratch, using KD baseline and our recipe can be found in Figure 8.
Models Layers Recipe Steps Pre-train Val loss ($\downarrow$)
GPT-2 Medium 24 half-width 100K 3.04
16 stacking 100K 2.84
16 hybrid-stacking 100K 2.83
16 Ours 100K 2.81
GPT-2 Large 36 half-width 100K 3.06
18 stacking 100K 2.87
18 hybrid-stacking 100K 2.89
18 Ours 100K 2.80
GPT-2 xLarge 48 half-width 100K 2.77
24 stacking 100K 2.65
24 hybrid-stacking 100K 2.64
24 Ours 100K 2.64
Table 2: Inheritune outperforms baseline zero-shot initialization and efficient training techniques. Comparison of pre-training validation loss for GPT-2 xLarge, GPT-2 Large and GPT-2 Medium variants. $\mathsf{Inheritune}$-derived models consistently achieve lower loss compared to models initialized with stacking, hybrid stacking, and half-width techniques.
4.1 Results and Discussions
Models trained with Inheritune outperforms much larger models trained from scratch.
We present our main results in Table 1. Our 24-layer, 18-layer, and 16-layer variants derived using $\mathsf{Inheritune}$ from the vanilla 48-layer GPT-2 xlarge, 36-layer GPT-2 large, and 24-layer GPT-2 medium achieve comparable or lower validation losses than their full-sized counterparts when trained for the same number of steps (100K steps). Our GPT-2 xlarge and GPT-2 large variants undergo one round of $\mathsf{Inheritune}$ training, while for GPT-2 medium, we perform three rounds of training with 12, 14, and 16 layers. We also evaluated all models on two next-word prediction downstream tasks in a zero-shot setting using the Wikitext Merity et al. (2016) and Lambada Paperno et al. (2016) datasets. The downstream performance of our GPT-2 models derived using the $\mathsf{Inheritune}$ recipe matches their much larger reference models. From the convergence perspective, some prior works have made connections between over-parameterization and faster convergence Bengio et al. (2005); Vaswani et al. (2018). In supplementary Figure 5, we show that the small LMs derived with $\mathsf{Inheritune}$, although smaller compared to their reference models, still converge as fast as their large-size reference models.
Models trained with Inheritune outperform same-sized models trained from scratch.
Table 1 demonstrates that GPT-2 variants trained with $\mathsf{Inheritune}$ outperform their same-sized counterparts trained from scratch, both when trained for the same number of steps and even when trained for double the steps (200K). This result underscores the efficiency of our approach. The only exception is the 24-layer GPT-2 xlarge variant, which surpasses both our model and the full-size model when trained for 200K steps.
Models trained with Inheritune outperform all zero-shot model initialization and efficient training baselines.
In Table 2, we compare GPT-2 xlarge, GPT-2 large, and GPT-2 medium variants trained with $\mathsf{Inheritune}$ against same-sized variants trained with stacking, hybrid, and half-width initialization baselines. The half-width baseline performs poorly, revealing the limitations of naive width reduction. Although stacking and hybrid stacking demonstrate reasonable performance, they still fall short compared to $\mathsf{Inheritune}$. In all cases, $\mathsf{Inheritune}$ consistently outperforms these baselines, highlighting its effectiveness as an initialization strategy. For a detailed view of the training dynamics across all methods, refer to the training curves in supplementary Figure 6.
Figure 3: A 16-Layer GPT-2 medium variant derived using Inheritune converges faster and generalizes better than a same-sized model trained with Logit-based distillation baselines. We conducted vanilla KD Hinton et al. (2015) and DistillBERT-style KD Sanh et al. (2019) with teacher initialization, using a 24-layer GPT-2 medium as the teacher for both KD baselines.
Distillation vs Inheritune.
In Figure 3, we compare 16-layer GPT-2 medium variants derived using vanilla knowledge distillation Hinton et al. (2015) and DistillBERT-style distillation Sanh et al. (2019), which leverages teacher layers for model initialization, vanilla training from scratch and $\mathsf{Inheritune}$. Our model demonstrates faster convergence and significantly better final generalization after 50K steps. Additional distillation experiments can be found in the supplementary material.
4.2 Ablations
Layers Initialization Steps Val loss ($\downarrow$)
16 attn (w/ layernorm) 100K 2.84
16 mlp (w/ layernorm) 100K 2.85
16 attn+mlp (w/o layernorm) 100K 2.80
16 Ours (attn+mlp w/ layernorm) 100K 2.81
Table 3: Impact of initializing various sub-modules within a transformer block. We compare the validation loss of a 16-layer GPT-2 medium variant when different sets of sub-modules are initialized with weights from the first 16 layers of a 24-layer GPT-2 medium reference model. All models are trained on the OpenWebText-9B dataset. Key findings: (1) $\mathsf{Inheritune}$ initialization and attention + MLP initialization result in similar performance improvements; (2) layernorm initialization shows minimal impact. A detailed training curve is shown in Figure 7.
We conducted extensive experiments to better understand which sub-module initializations within a transformer block lead to improved generalization (in terms of validation loss) and faster convergence. For these ablations, we fixed the model to a 16-layer GPT-2 medium variant and explored three different sub-module initializations using weights from a 24-layer GPT-2 medium reference model. We initialize the transformer blocks with 1) attention ((key, query, value, and projection) and the layernorm [^2] In GPT-2 models layernorm blocks are parameterized. weights (attn w/ layernorm), 2) attention and mlp weights without the layer-norm (attn+mlp w/o layernorm), and 3) mlp weights with the layer norm (mlp w/ layernorm). We emphasize that $\mathsf{Inheritune}$ performs initialization by inheriting attention and mlp weights with the layer norm (attn+mlp w/ layernorm).
As shown in Table 3, models trained with attention and mlp weights demonstrated the best performance, regardless of the layer norm initialization. A detailed validation loss vs training steps plot is presented in supplementary Figure 7. We conclude that initializing both attention and MLP weights provides a clear advantage. Surprisingly, we also observed that initializing either the attention or mlp weights resulted in similar improvements in both convergence speed and final validation loss.
5 Training without Data Repetition
(a) GPT2-Large
(b) GPT2-Medium
Figure 4: Models derived using Inheritune without data repetition converge faster and match the final validation loss of the full-sized model despite using fewer layers. Additionally, the model trained using $\mathsf{Inheritune}$ demonstrates data efficiency, achieving a lower validation loss in fewer steps compared to its full-sized and half-sized counterparts until 80% of the training.
Are the gains we observe due to the Inheritune recipe merely a consequence of over-fitting due to data repetition?
To investigate this, we conducted additional training experiments without data repetition, following standard LLM pre-training practices as discussed in Touvron et al. (2023a); Biderman et al. (2023). Moreover, we utilized a high-quality pre-training dataset, Fineweb_edu Penedo et al. (2024), which contains 100B tokens and has been deduplicated and filtered to ensure high data quality.
We train a 32-layer GPT-2 large (668M) and a 24-layer GPT-2 medium (355M) reference model from scratch. Next, we trained two 16-layer variants: one derived from GPT-2 large and the other from GPT-2 medium, using their respective reference models following Algorithm 1. Finally, we trained the 16-layer baseline variants from scratch for comparison. All these models are trained for 100K steps. The model configurations and training hyper-parameters can be found in supplementary Table 10.
As shown in Figure 4, our GPT-2 variants trained using $\mathsf{Inheritune}$ consistently perform on par with their full-size counterparts and outperform their same-size counterparts in terms of training loss. In LLM pre-training literature, where data is not repeated, training loss has been shown to be a reliable metric Touvron et al. (2023a, b). Furthermore, we performed zero-shot downstream evaluations using lm-evaluation-harness Gao et al. (2024) on a variety of tasks, including ARC-easy (ARCE; Clark et al. (2018)), LAMBADA Paperno et al. (2016), SciQ Welbl et al. (2017), Hellaswag Zellers et al. (2019), and PIQA Bisk et al. (2020). As shown in Table 4, on average, the models demonstrate superior performance than the baseline models trained from scratch.
Models Recipe Layers ARCE PIQA SciQ Hellaswag Lambada Average
(acc) (acc) (acc) (acc norm) (acc)
GPT-2 Medium rand init 24 51.05 61.81 74.8 30.79 20.28 47.74
rand init 16 49.92 61.92 73.3 29.56 19.54 46.84
Ours 16 51.26 61.81 73.8 30.55 23 48.08
GPT-2 Large rand init 32 52.48 64.58 75.3 32.65 22.2 49.44
rand init 16 50.34 63.11 75 30.86 21.56 48.17
Ours 16 52.9 63.55 76.1 32.14 24.06 49.75
Table 4: Models trained with Inheritune outperform both their larger and same-size counterparts trained from scratch on average zero-shot downstream performance. For evaluation we choose accuracy (acc) and normalized accuracy (acc norm) metrics following Open LLM leaderboard Beeching et al. (2023). All the models are trained with FineWeb_edu.
6 Related Works
Attention degeneration has been studied in the past through the lens of attention rank collapse Dong et al. (2021) leading to representation collapse, and attention entropy collapse Zhai et al. (2023) leading training instability. This also has been studied as a theoretical setup for transformer models by Noci et al. (2022); Barbero et al. (2024). He et al. (2023) address rank collapse in self-attention networks (SANs) without residual connections or layer norms, using two model initialization techniques that enable faithful signal propagation, i.e., $\Sigma_{L}$ of $A(X^{L})$ does not collapse in deeper layers. However, this approach significantly slows down training. Noci et al. (2022) proposes scaling residual connections by $1/\sqrt{L}$, while Barbero et al. (2024) suggests that adding additional tokens to already long sequences of repeated tokens can help mitigate degeneration. In contrast to previous work, we address attention degeneration by developing smaller models that eliminate structural inefficiencies and training these models to match the performance of their larger, inefficient counterparts.
LLM training recipes and model initialization. The stacking method Gong et al. (2019); J. Reddi et al. (2023), which employs a stage-wise training strategy that uses weights from initial layers to initialize later layers, has been shown to be effective for LLM training both empirically Gong et al. (2019); J. Reddi et al. (2023); Du et al. (2024) and theoretically Agarwal et al. (2024). Knowledge distillation Hinton et al. (2015) has been very successful in training small LMs in some cases Turc et al. (2020); Sanh et al. (2019) the smaller student model is also initialized with teacher layers-though this is often done without clear explanation or intuition. Recent works in model initialization, such as Trockman & Kolter (2023), have studied synthetic attention patterns for initialization, primarily in vision settings. However, such methods have limited success in language models. Xu et al. (2024) use weight initialization for faster fine-tuning of vision models. In contrast, our proposed recipe focuses on creating a smaller model by eliminating specific structural inefficiency in lazy layers. This distinction sets our work apart in terms of both objective and methodology.
7 Conclusion
In this study, we identified a structural flaw in the attention mechanism of deep decoder-style transformers, which are currently the de-facto architectural choice for LLMs. Specifically, we discovered that the attention matrices in many deeper layers tend to lose rank and converge into single-column matrices, resulting in what we term lazy layers. This phenomenon causes the model to lose attention and leads to structural inefficiencies in LLMs.
To avoid this issue, we propose a simple training recipe $\mathsf{Inheritune}$ where we inherit a few early blocks from a larger reference model and continue training and expanding the model with more blocks until it matches the performance of the reference model. We evaluated our method using 48-layer, 36-layer, and 24-layer GPT-2 xlarge, GPT-2 large, and GPT-2 medium reference models. Through this process, we successfully developed smaller models with 24, 18, and 16 layers without losing any performance compared to their larger counterparts while utilizing the OpenWebText-9B dataset. We observed thematically similar results with GPT-2 large and GPT-2 medium models trained on FineWeb_edu without data repetition. These findings offer a novel perspective on developing efficient small base language models from existing large pre-trained models, potentially democratizing LLM pre-training for a wider community.
Acknowledgments
We thank Nived Rajaraman, for his helpful feedback and discussion during this work.
Supplementary Materials
Contents
  • A: Supplementary Experiments
  • B: Additional Experiments on Low Data Regime
  • C: Implementation Details
Appendix A Supplementary Experiments
We provide additional training plots for our main results discussed in Section 4.1 as shown in Figure 5 and Figure 6. In Figure 5 (also refer Table 1 we compare our GPT-2 variants with baseline models trained form scratch. In Figure 6 (also refer Table 2) we compare our GPT-2 variants with baseline models trained using baseline zero-shot model initialization (and also re-training) techniques.
In Figure 7, we present the training curves of models trained during ablation as discussed in Section 4.2.
Knowledge Distillation
Recall we have already discussed distillation as a baseline in Section 4.1 and associated Figure 3. We perform an additional experiment in the same setting i.e. knowledge distillation as a baseline. Here we trained GPT-2 medium variants with 12 layers (half the number of a vanilla GPT-2 medium). We trained three models. First we distilled a 24-layer GPT-2 medium (teacher) to a 12-layer GPT-2 medium variant (student) and this student is initialized with all the alternate layers of the teacher. This setting is exactly same as discussed in DistillBERT Sanh et al. (2019). Next we trained two GPT-2 medium variants one from scratch (vanilla training) and the other with $\mathsf{Inheritune}$ recipe. Model trained with our recipe beats model trained with distillation. We defer a through investigation of distillation compared to $\mathsf{Inheritune}$ to future work.
How Inheritune addresses Attention Degeneration?
Recall we have discussed attention degeneration in Section 2 and attention patterns are visualized in Figure 2. Following up on our previous discussions in Figure 9 we demonstrate that models trained with $\mathsf{Inheritune}$ has lesser lazy layers (refer Figure 9) compared to it’s larger counterpart trained form scratch. We performed rank analysis utilizing 24-layer GPT-2 medium and our 16-layer GPT-2 model trained using $\mathsf{Inheritune}$.
Recall we have previously discussed that attention degeneration is connected with vanishing gradients of keys and queries Noci et al. (2022). The vanishing gradients is caused when the norm of the gradients Bengio et al. (1994) are so small that it fails to generate meaningful back-propagation signal. Since we are training smaller models intuitively $\lVert W_{Q}\rVert$ and $\lVert W_{K}\rVert$ should be smaller compared to their larger counterparts and hence the norm of gradients in the case of smaller models derived using $\mathsf{Inheritune}$ is higher leading to better training.
(a) GPT2-Medium
(b) GPT2-Large
(c) GPT2-xLarge
Figure 5: Models derived using Inheritune converge faster and match the final validation loss of the full-sized model trained from scratch, despite being smaller. Training GPT-2 xlarge, GPT-2 large and GPT-2 medium vanilla models from scratch and our variants with OpenWebText-9B for 100K steps.
(a) GPT2-Medium
(b) GPT2-Large
(c) GPT2-xLarge
Figure 6: Models derived using Inheritune outperform three zero-shot initialization and efficient training baselines in terms of final validation loss. Our models demonstrate better convergence and generalization compared to all baselines. We trained GPT-2 xlarge, GPT-2 large and GPT-2 medium variants on OpenWebText-9B for 100K steps using baseline model initialization and efficient training techniques and our $\mathsf{Inheritune}$ training recipe.
Figure 7: Full training curves of 16-layer GPT-2 variants trained during ablations. We analyze $\mathsf{Inheritune}$ approach while initializing some specific sub-modules in transformer blocks. Here, we initialize each transformer block of a 16-layer GPT-2 medium variant with three different configurations. First, we separately initialize attention and MLPs (FFNs) submodules; second, we initialize the attention and MLP weights while randomly initializing the layer norms. Finally, we perform $\mathsf{Inheritune}$-initialize only the attention and MLP weights with all the respective layer norms.
Figure 8: A 12-Layer GPT-2 medium variant derived using Inheritune converges faster and generalizes better than a same-sized models trained from scratch and with Logit-based distillation with teacher initialization baseline. Three 12-layer GPT-2 medium variants were trained: (1) a distilled model initialized with alternate layers from a 24-layer GPT-2 medium teacher, following the DistillBERT setup Sanh et al. (2019); (2) a model trained from scratch (vanilla training); and (3) a model trained using the $\mathsf{Inheritune}$ recipe. The model trained with $\mathsf{Inheritune}$ outperforms both the distillation-based model and the one trained from scratch, demonstrating the effectiveness of our approach.
(a) Vanilla GPT2-Medium 24 layers
(b) Our GPT2-Medium 16 layer variant trained with $\mathsf{Inheritune}$
Figure 9: Rank collapse in deeper layers and its mitigation through $\mathsf{Inheritune}$. (a) Analysis of a 24-layer GPT2 medium model reveals rank-1 attention matrices in later layers, indicating rank collapse. (b) Our 16-layer GPT2 medium variant, trained with $\mathsf{Inheritune}$, demonstrates improved rank across all layers, highlighting the effectiveness of our approach in preserving expressive power.
Appendix B Developing a 1.5B Small Base LM in a Low Data Regime with Inheritune
Figure 10: Performance of our 1.5B base LM derived using 1B data with $\mathsf{Inheritune}$ on an average of 9 different datasets (left) and MMLU benchmark (right) that evaluates commonsense, truthfulness, natural language inference and language understanding. We compare our model’s performance with reference model-OpenLLamA-3B (2x size), other small base LMs of size 1B-2B parameters such as MPT-1.3B, OPT-1.3B, Pythia-1.4B (pre-trained from scratch) and ShearLLaMA-1.5B (pruned and continually trained using existing large base LM).
In this section, we aim to investigate the efficacy of $\mathsf{Inheritune}$ in a data and compute-constrained setting. We train a 1.5B parameter small base LM with only 1B tokens using a 3B parameter base LM on a single GPU (A6000) for less than half a day.
We assume the existence of a pre-trained reference model $\mathcal{M}_{\text{ref}}$, comprising $k$ layers, represented by $\mathsf{W}_{\text{ref}}=\{\mathsf{W}_{0},\mathsf{W}_{1},\ldots,\mathsf{W}_{k-1}\}$ trained with $\mathcal{D}_{\text{train}}$. However, this full training data is unavailable, and we only have a random tiny subset $\hat{\mathcal{D}}_{\text{train}}\sim\mathcal{D}_{\text{train}}$. We use OpenLLaMA-3B version 1 as the reference model pre-trained with 1T tokens from the RedPajama V1 dataset, which contains data from various domains such as common crawl, C4, Wikipedia, books, arXiv papers, GitHub, and Stack Exchange. We take 1B randomly sampled tokens [^3] https://huggingface.co/datasets/togethercomputer/RedPajama-Data-1T-Sample from the RedPajama dataset.
Training recipe.
To adapt $\mathsf{Inheritune}$ for this new setting, we perform step 1 and step 2 in Algorithm 1 without growing the model (i.e., we skip step 3). We use the first $n=13$ layers from our $k=26$ layer reference model. We call our small base LM Ours-1.5B(#tokens). We train our model with data repetition for eight epochs (each epoch uses all the 1B tokens) with a batch size of 131K tokens per batch. We use 1 A6000 GPU for less than half a day of training. The choice of training epochs is based on the analysis provided later in this paper (refer to Figure 13). We use the lit-gpt framework for training all small base LMs discussed in this paper. Further discussions on the training hyper-parameters can be found in the next Section.
Baseline models and evaluation.
We choose similarly sized (1-2B parameter) small base LMs trained with the RedPajama dataset and the reference base LM as primary baselines, as the quality of the pre-training data plays a key role in model development. We also include models OPT-1.3B Zhang et al. (2022) and Pythia-1.3B Biderman et al. (2023) as these models are pre-trained with a dataset similar to the RedPajama dataset. Table 6 lists the baseline models with their pre-training data.
In this study, we use few-shot accuracy, particularly 0-shot and 5-shot accuracy, on ten different downstream tasks to measure the quality of our 1.5B base LM. This evaluation of pre-trained base LLMs has been done in several prior works. Our evaluation methodology categorizes downstream tasks across four distinct genres: commonsense reasoning, natural language understanding, factuality, and natural language inference. We perform 0-shot evaluation for PIQA Bisk et al. (2020), BOOLQ Clark et al. (2019), WINOGRANDE Sakaguchi et al. (2020), WINOGRAD Kocijan et al. (2020), LOGIQA Liu et al. (2020), TruthfulQA Lin et al. (2022), MNLI Bowman et al. (2015), QNLI Wang et al. (2018) and WNLI Wang et al. (2018) datasets. Next, we perform a 5-shot evaluation on the massive multitask language understanding benchmark (MMLU) Hendrycks et al. (2020). We use the lm eval harness framework Gao et al. (2024) for the entire evaluation.
Model Commonsense Reasoning
Name (# train tokens) Reference Winograd PIQA Boolq WinoGrande Logiqa
OpenLLaMA-3B (1T) n/a 63.46 74.97 67.18 62.27 28.4
OPT-1.3B (300B) n/a 38.46 71.82 57.83 59.51 27.04
Pythia-1.4B (300B) n/a 36.54 70.89 63.12 56.99 27.65
MPT-1.3B (200B) n/a 63.46 71.44 50.89 58.09 28.26
Sheared LLaMA-1.3B (50B) LLaMA2-7B 36.54 73.45 62.02 58.17 27.34
Ours-1.5B (1B) OpenLLaMA-3B 50.96 56.47 61.68 51.69 25.19
Model Lang. Understanding & Inference Factuality
Name ( # train tokens) Reference MMLU(5) WNLI QNLI MNLI TruthfulQA
OpenLLaMA-3B (1T) n/a 27.21 50.7 51.3 37.3 35
OPT-1.3B (300B) n/a 24.96 42.25 51.29 35.82 38.67
Pythia-1.4B (300B) n/a 25.56 53.52 49.48 32.76 38.66
MPT-1.3B (200B) n/a 25.82 40.85 50.52 35.93 38.68
Sheared LLaMA-1.3B (50B) LLaMA2-7B 25.71 49.3 50.98 37.94 37.14
Ours-1.5B (1B) OpenLLaMA-3B 25.67 43.66 49.41 34.42 48.61
Table 5: Our 1.5B model achieves performance comparable to baseline models despite being trained with fewer tokens. Comparison of our target model ($\mathcal{M}_{\text{tgt}}$) derived using $\mathsf{Inheritune}$ with the reference model ($\mathcal{M}_{\text{ref}}$) and other baseline models of similar size when pre-trained from scratch and pre-trained with inherited weights and pruning. Although trained with fewer tokens, our model achieves performance comparable to the baseline models. We have highlighted all the scores in bold where our 1.5B model achieves at least 90% of the score compared to the reference LM or outperforms at least two of the baseline similar-size LMs. All the tasks are evaluated using 0-shot except MMLU, which is 5-shot. The models against which n/a is mentioned are trained from scratch.
B.1 Main Results in Low Data Regime
Table 5 presents a detailed performance evaluation across various tasks. Our 1.5B model, developed using $\mathsf{Inheritune}$, excels in 7 out of 10 individual tasks. It achieves a score of 90% or higher compared to the reference language model, which is twice its size and trained with 1000 times more data, or it outperforms at least two other base LMs of similar size trained with 50-300 times more data. Favorable scores are highlighted in bold.
Next, we compare our small LM with the MPT-1.3B [^4] https://huggingface.co/mosaicml/mpt-1b-redpajama-200b model trained from scratch with 200B tokens of RedPajama dataset and find that we match 97% accuracy in all nine downstream tasks and the MMLU (5-shot) score. Additionally, we compare with OPT-1.3B and Pythia-1.3B models, showing that we outperform both in the MMLU (5-shot) score and perform comparably on the other nine datasets. This comparison illustrates that having a large reference base LM and a subset of its pre-training data allows the inherited target size base LM to be trained remarkably more sample-efficiently than training from scratch. Extended discussions on comparisons with the ShearedLLaMa model, generated by pruning and continual training from LLaMA2-7B, are provided in the supplementary materials.
Ablation of Inheritune Across Different Model Sizes with 1B Tokens.
In the previous section, we considered a single choice of $n=k/2$, i.e., half the layers, for the size of the smaller model. Here, we investigate $\mathsf{Inheritune}$ with different choices of $n$, but the same 1B token dataset). All models use OpenLLAMA-3B as the large pre-trained reference model, with consistent training hyperparameters, changing only the choice of $n$.
We developed eight different submodels with $n=\{4,6,8,10,13,16,18,20\}$. Figure 11 shows the MMLU (5-shot) score as a function of $n$. As expected, the trend line is positive-sloping. The submodel with 20 layers slightly decreases performance, potentially due to data overfitting as the model size increases. The training details for all these submodels are consistent with the target 1.5B small base LM and are detailed in the appendix. A more comprehensive investigation on the choice of $n$—including varying both $n$ and the number of training tokens jointly and evaluating a broader set of tasks is left for future work.
Figure 11: Inheritune scales across multiple different model sizes. Utilizing the OpenLLaMA-3B as a reference large base LM, demonstrates that multiple performant small base LMs of target size can be crafted using $\mathsf{Inheritune}$ with just 1B training tokens. The MMLU (5-shot) as a function of the number of submodels.
B.2 Additional analysis with larger reference LMs and 50B data
We further analyze $\mathsf{Inheritune}$ to see the impact of it’s performance when more tokens are available. Initially for the main results we limited ourselves to 1B (i.e. 0.1%) tokens from the 1T pre-training data, here we use a 50B subset (i.e. 5%) of the pre-train data. Moreover we also extend this study to include larger base LMs of 7B parameters as reference models, employing OpenLLaMA-7B and LLaMA2-7B as reference models. For the purpose of this study we do not repeat the tokens from our 50B subset. As shown in Figure 12, we observe that there is clear improvement in overall MMLU (5-shot) score with more data. Additionally it is interesting to see that 1.5B (or 1.6B models) developed with $\mathsf{Inheritune}$ using larger reference models show even greater improvements when fed with 50B subset of non repetitive data (i.e fresh tokens). We present a Table 8 using Figure 12 to show the best MMLU (5-shot) scores achieved using different reference LMs. For developing our small base LMs using larger reference LMs we use $n$=7 (i.e. 7 layers). The training details are discussed in the following section.
Model Training Data (# tokens)
OpenLLaMA-3B v1(ref) RedPajama(1T)
Ours-1.5B* RedPajama (1B)
Shear-LLaMA-1.3B* RedPajama(50B)
MPT-1.3B RedPajama(200B)
Pythia-1.4B The Pile(300B)
OPT-1.3B Custom data(300B)
Table 6: Comparison of training data across baseline models. Overview of reference and baseline models, including their pre-training datasets and the number of tokens used during training. Note the significant variation in training data size, ranging from 1B to 1T tokens.
Models (# train tokens) GPU Count GPU Type Time (# days)
MPT-1.3B (200B) 440 A100 half
Pythia-1.4B (300B) 64 A100 4.6
TinyLLaMA-1.1B (3T) 16 A100 90
OPT-1.3B (300B) 992 A100
Sheared LLaMA-1.3B (50B) 16 A100
OpenLLaMA-3B (1T) 256 TPU v4 10
Our-1.5B (1B) 1 A6000 ~half
Table 7: Computational efficiency of Inheritune versus baseline models. Comparison of pre-training compute requirements for publicly available small base LMs and our $\mathsf{Inheritune}$-derived model. Metrics include GPU count, GPU type, and training duration, highlighting Inheritune’s significant reduction in computational resources.
Figure 12: Impact of reference model choice on Inheritune performance. MMLU (5-shot) scores for 1.5B base LMs derived using $\mathsf{Inheritune}$, trained on 50B unique tokens. Comparison across three reference models: OpenLLaMA-7B, LLaMA2-7B, and OpenLLaMA-3B. Results demonstrate Inheritune’s effectiveness with various large language models as references.
Figure 13: Performance of our 1.5B base LM derived using $\mathsf{Inheritune}$ based on existing OpenLLaMA-3B base model. Here we use 1B tokens and perform data repetition (epochs) during training. We further evaluate our model on an average of 9 different datasets (left) and MMLU benchmark (right).
Model (# tokens), ref MMLU(5)
Ours-1.6B (1B), LLaMA2-7B 24.27
Ours-1.5B (1B), OpenLLaMA-3B 25.67
Ours-1.5B (50B), OpenLLaMA-3B 25.71
Ours-1.6B (50B), LLaMA2-7B 26.07
Ours-1.6B (50B), OpenLLaMA-7B 26.72
Table 8: Performance comparison of models on the MMLU (5-shot) task. Our models, even when trained with fewer tokens, show competitive performance compared to benchmarks. We have highlighted the best MMLU 5-shot score in bold.
Model (# tokens) Data type MMLU (5-shot)
Ours-1.5B (1B) 10 epochs 24.95
Ours-1.5B (50B) 10B fresh 23.62
Ours-1.5B (1B) 20 epochs 25.46
Ours-1.5B (50B) 20B fresh 24.96
Table 9: MMLU (5-shot) scores of Our-1.5B small base LM derived using 1B data for multiple data repetition–10 epochs and 20 epochs compared to the same model trained without data repetition for 10B and 20B fresh tokens. We derive all the variants of Our-1.5B small base using $\mathsf{Inheritune}$ with OpenLLaMA-3B as reference model. The models featured in this table correspond to those discussed in Figures 12 and 13.
Ablations with number of epochs.
We ran ablations (refer Figure 13) to choose the total number of epochs (multiple passes over the data) and observe that repetition when training our 1.5B (or 1.6B) LM is helpful particularly for MMLU. We also observe that the for an average of all the 9 other datasets (i.e. except MMLU) peaks it’s performance at 5 epochs and then deteriorates. Some prior works have studied this phenomenon that the scaling of downstream tasks with data is not always linear Biderman et al. (2023).
To repeat or not to repeat the tokens.
Next we tackle the question – whether one should re-use 1B tokens for multiple epochs or use the same number of fresh tokens? Some prior works have recommended that if you have a reasonably large size dataset one can repeat it upto 4 epochs Muennighoff et al. (2023). In our study we observe that one can safely re-use 1B tokens upto 10-20 epochs as shown in Table 9. We emphasis that this phenomenon needs a through investigation in itself and we defer this to future work. The models discussed in Table are saved checkpoints during a single training run and not the final model unless otherwise specified.
B.3 Implications of Low Data Regime
In this section, we discuss some of the key implications of our work in low data regime.
Cheap and easy development of small base LMs.
Pre-training a small base LM of 1-2B parameters from scratch is extremely expensive. For instance mpt-1.3B base LM is pre-trained with 440 A100 GPUs for half a day, while the Pythia-1.4B base LM Biderman et al. (2023) utilized 64 A100-40GB GPUs for 4.6 days. Similarly, TinyLLaMA-1.1B model Peiyuan Zhang & Lu (2023) was pre-trained using 16 A100 GPUs for 3 months. Our 1.5B (1B data variant) LM shows competitive performance despite being trained with 1 A6000 GPU for less than 12 hours. The computational details are provided in Table 7, comparing the training resources of the baseline models listed in this paper. Typically small base LMs are finetuned for a specific task before deployment and are not used in it’s base form. With $\mathsf{Inheritune}$ we present a really easy and cheap way for developing a small base LM to be later finetuned before deployment.
Naive baseline for pre-training a scaled down variant of large base LMs.
Typically small variants of large base LMs are pre-trained using the same pre-training data Peiyuan Zhang & Lu (2023); Groeneveld et al. (2024). Our recipe introduces a new perspective of identifying sufficient depth without losing any generalization on the held out validation set. Next, we also show that even with a small fraction of pre-train data (randomly sampled) and few initial layers of the large base LM one can develop a small base LM. Therefore our $\mathsf{Inheritune}$ recipe has the potential to become the naive baseline for any pre-training pipeline aiming to develop a smaller variant of a large base LM.
Appendix C Implementation Details
C.1 Training details of GPT-2 models
For our main experiments, we focused on three sizes of GPT-2 models Radford et al. (2019): the vanilla GPT-2 xlarge with 1.5B parameters, GPT-2 large with 770M parameters and the vanilla GPT-2 medium with 355M parameters. We developed several variants of these models by adjusting the number of layers and hidden size. We trained all GPT-2 models with data repetition while using OpenWebText dataset, the trainset has 9B tokens and the validation set has 4.4M tokens. The key architectural configurations for the reference models, our models, and baseline models discussed in this paper are summarized in Table 10.
For all training runs, we used GELU activations, disabled bias terms, and removed dropout, following the nanoGPT codebase and Liu et al. (2023). We employed the AdamW optimizer with $\beta_{1}=0.90$ and $\beta_{2}=0.95$. The GPT-2 models were trained on a single node with 3 A100 GPUs (each with 40 GB of memory) using distributed data parallelism and gradient accumulation. In line with Liu et al. (2023), we scaled the attention logits inversely to the layer index across all GPT-2 models. Most hyperparameters were adapted from Liu et al. (2023), with key details provided below.
Hyper-parameter details of GPT-2 Medium and variants.
  • Batch size: 50K tokens
  • Learning rate: $3\times 10^{-4}$,
  • Warmup steps: 2K,
  • Scheduler type: cosine decay to $\frac{1}{10}$ of max learning rate,
  • Weight decay: 0.1,
  • Gradient clipping value: 1,
  • Total training steps: 100K
Hyper-parameter details of GPT-2 large and variants.
  • Batch size: 16K tokens
  • Learning rate: $2\times 10^{-4}$,
  • Warmup steps: 2K,
  • Scheduler type: cosine decayed to $1\times 10^{-5}$,
  • Weight decay: 0.1,
  • Gradient clipping value: 1,
  • Total training steps: 100K
Hyper-parameter details of GPT-2 xlarge and variants.
  • Batch size: 16K tokens
  • Learning rate: $1.5\times 10^{-4}$,
  • Warmup steps: 2K,
  • Scheduler type: cosine decayed to $1\times 10^{-5}$,
  • Weight decay: 0.1,
  • Gradient clipping value: 1,
  • Total training steps: 100K
Hyper-parameter details of knowledge distillation training.
We use the below loss for as our distillation based training loss. The validation loss is the student_loss.
[…]
  • Model: 16-layer and 12-layer GPT-2 medium variants
  • Softmax temperature: 1
  • $\alpha$: 0.6
  • Batch size: 50K tokens
  • Learning rate: $3\times 10^{-4}$,
  • Warmup steps: 2K,
  • Scheduler type: cosine decay to $\frac{1}{10}$ of max learning rate,
  • Weight decay: 0.1,
  • Gradient clipping value: 1,
  • Total training steps: 50K
Models Layers Hidden Size Heads Variant
[…]
GPT2-large(770M) 36 1280 20 Original
GPT2-large(680M) 32 1280 20 Original
GPT2-medium(355M) 24 1024 16 Original
[…]
GPT2-medium 16 512 8 half width
[…]
GPT2-large 18 1280 20 Ours
GPT2-large 16 1280 20 Ours
GPT2-medium 16 1024 16 Ours
Table 10: Overview of all the GPT2 models used in this study and their architectural configurations. The model configurations of stacking and hybrid stacking are same as our variants.
C.2 Training details of 1.5B OpenLLaMA model
Small base LMs trained with 1B data
We present our main results with Our-1.5B model trained with an existing OpenLLaMA version 1 Geng & Liu (2023) and 1 B tokens randomly sampled from 1T redpajama version1 data. The hyper-parameters related to this model is provided below. It is important to note that our claim that we only use 1 GPU for less than 12 hours to train Our-1.5 B model is specific to models derived using $\mathsf{Inheritune}$ with 1B data. Next we also train multiple sub-models as shown in Figure 11 the training details remains consistent with that of the initial model discussed earlier. However we observe that increasing the number of layers in a sub-model also increase the training time.
Hyper-parameter details of our 1.5B base LM derived using OpenLLaMA-3B as refernce LM:
  • Training tokens: 1B
  • Training epochs: 8
  • Training steps: 64K
  • Learning rate: $3\times 10^{-4}$
  • Scheduler: Cosine
  • Weight decay: 0.1
  • Optimizer: AdamW
  • Warm up steps: 1000
  • Batch size: 131K
  • GPU count: 1
  • GPU type: A6000
  • GPU hours: $\sim$ 8 hours
  • GPU hours/epoch: $\sim$ 54 minutes
Training details of small base LMs with 50B data.
We also trained our 1.5B model with larger subsets of data as shown in Figure 12.It is important to note that all the intermediate tokens until 50B are intermediate checkpoints of a single training run. Some of the key hyper-parameters of our training runs are discussed below. We have also trained three variants of small base LMs utilizing 3 different reference base LMs namely OpenLLaMA-3B, OpenLLaMA-7B and LLaMA2-7B. For target LMs developed with OpenLLaMA-3B we use $n$=13 i.e. 13 layers. For target LMs developed using reference LMs of 7B parameters we use $n$=7 i.e. 7 layers. The training hyper-parameters remains consistent across all the models trained with 50B subset of the pre-train data.
Training hyper-parameters of our target 1.5B and 1.6B small base LMs:
  • Training tokens: 50B
  • Training epochs: $\sim$1
  • Training steps: 191K
  • Learning rate: $3\times 10^{-4}$
  • Scheduler: Cosine
  • Weight decay: 0.1
  • Optimizer: AdamW
  • Warm-up steps: 1000
  • Batch size: 131K tokens
  • GPU count: 1
  • GPU type: A100
  • GPU hours: $\sim$18 hours

### v3

@@
-Inheritune: Training Smaller Yet More Attentive Language Models
+When Attention Collapses: How Degenerate Layers in LLMs Enable Smaller, Stronger Models
 Abstract
-Large Language Models (LLMs) have achieved remarkable performance across various natural language processing tasks, primarily due to the transformer architecture and its self-attention mechanism.
-However, we observe that in standard decoder-style LLMs, attention matrices degenerate to single-column for deeper layers.
-Layers in this state are unable to learn anything meaningful and mostly redundant; we refer to these as lazy layers.
-The goal of this paper is to train smaller models by eliminating this structural inefficiency without compromising performance.
-Motivated by this observation, we propose Inheritune, a simple yet effective training recipe for developing smaller, high-performing language models.
-Smaller models trained with $\mathsf{Inheritune}$ inherit early transformer layers from a larger pre-trained model, then retrain and progressively expand until they match or exceed the performance of the larger model.
-We demonstrate that $\mathsf{Inheritune}$ enables the training of various sizes of GPT-2 models on datasets like OpenWebText-9B and FineWeb_Edu.
-Models trained with $\mathsf{Inheritune}$, despite having significantly fewer layers, match or even surpass the performance of their larger counterparts.
-For instance, our 16-layer GPT-2 medium variant achieves comparable performance to the standard 24-layer GPT-2 medium model.
+Large Language Models (LLMs) leverage the transformer architecture and its self-attention mechanism to achieve remarkable performance.
+However, we identify a structural inefficiency in standard pre-trained decoder-style LLMs: attention matrices in many deeper layers frequently degenerate, often collapsing to near rank-one, single-column structures.
+These layers, which we term lazy layers, are redundant and inefficient.
+To address this, we introduce $\mathsf{Inheritune}$, a simple yet effective training recipe for developing smaller, more attentive, and high-performing language models.
+$\mathsf{Inheritune}$ initializes a compact model by inheriting the potent early layers from a larger pre-trained model, then progressively retrains and expands it.
+Our demonstration of $\mathsf{Inheritune}$’s efficacy on various models and datasets reveals that models trained with this approach, despite having significantly fewer layers, can achieve performance comparable to or even exceeding their larger counterparts.
+This approach enables model compression by design, producing language models that are both compact and high-performing.
 Code is available at https://github.com/sanyalsunny111/LLM-Inheritune.
@@
 Large Language Models (LLMs) are built with decoder-style transformer blocks (Vaswani et al., 2017).
-These models are typically designed to be large, with a significant portion of their parameters dedicated to their depth, with multiple transformer blocks stacked together, building model capacity.
+These models are typically designed to be large, with a significant portion of their parameters dedicated to their depth, with multiple transformer blocks stacked with eachother building model capacity.
 Each block or layer in the stack refines the representations learned by the previous blocks, allowing the model to develop a nuanced understanding of the input data.
@@
 It allows models to combine tokens as a weighted linear sum of their attention scores, effectively capturing long-range dependencies and contextual relationships within text data.
-However, as models grow in depth, they often encounter a phenomenon known as attention degeneration, caused by a collapse in the attention rank Noci et al.
-(2022); Dong et al.
-(2021); He et al.
-(2023).
+However, as models grow in depth, they often encounter a phenomenon known as attention degeneration caused by collapse in the attention rank ((Noci et al., 2022; Dong et al., 2021; He et al., 2023)).
 Notably, this phenomenon has not been studied in the context of standard LLMs.
@@
 (2019) for attention degeneration and observe that many deeper layers in both models exhibit rank-1 attention matrices.
-Further investigation reveals that most of these rank-1 matrices are also single-column, i.e., their mass is concentrated in a single column.
+Further investigation reveals that most of these rank-1 matrices are also single-column, i.e.
+their mass is concentrated to a single column.
 Our attention matrix analysis is presented in Figure 1.
-We term these deeper layers, where all attention matrices of a given layer are degenerate, as lazy layers.
-Motivated by this new finding, we aim to develop performant small base language models (LMs) that utilize weights from inefficient larger base LMs.
+We term these deeper layers, where all attention matrices of a given layer are degenerated, as lazy layers.
+Motivated by this new finding we aim to develop performant small base language models (LMs) utilizing weights from inefficient larger base LMs.
 A base LM is a decoder-style model trained solely for next-token prediction without additional enhancements like instruction tuning or reinforcement learning with human feedback (RLHF).
@@
 During the growth phase, the newly added blocks can be initialized with lazy layers of the reference LM.
-We refer to this simple, yet effective, training approach as $\mathsf{Inheritune}$.
-(a) Rank analysis of vanilla GPT-2 Medium
-(b) Matrix mass analysis of vanilla GPT-2 Medium
-(c) Model initialization with later layers
-(d) Rank analysis of vanilla GPT-2 Large
-(e) Matrix mass analysis of vanilla GPT-2 Large
-(f) Model initialization with later layers
-Figure 1: Attention matrices of many deeper layers often degenerate to single-column matrices in regular decoder style LLMs, layers with fully degenerate attention fail to learn meaningful representations.
-We computed a single attention matrix with 100 tokens from the OpenWebText validation set with 4M tokens.
-Next we performed 100 runs and plotted the mean and std of the max rank and mass as a function of layers for our rank and mass analysis respectively.
-Figure (a and d): An analysis of a 24-layer GPT-2 medium and a 36-layer GPT-2 large shows the max rank of the attention matrices across all layers.
-Figure (b and e): A closer look at the same GPT-2 models also reveals that the dominant mass proportion of several attention matrices is concentrated in a single-column, particularly in deeper layers.
-Figure (c and f): When initializing 12-layer and 18-layer variants of the vanilla GPT-2 medium and GPT-2 large models with deeper layers (Lazy layers) exhibiting degenerated attention–their performance is comparable to models with random initialization.
-However, initializing models with early layers leads to significantly better generalization and convergence.
+We refer to this simple yet effective training approach as $\mathsf{Inheritune}$.
+(a) GPT-2 Medium (355M) L=24, H=1024
+(b) GPT-2 Large (770M) L=36, H=1280
+(c) GPT-2 XLarge (1.5B) L=48, H=1600
+Figure 1: In decoder-style LLMs, attention matrices in deeper layers often degenerate to near rank-1, limiting their ability to learn meaningful representations.
+We compute $\text{MaxRank}^{(l)}$ (averaged over $N=100$ randomly selected sequences each with $T=100$ tokens) for each layer $l$ using the OpenWebText validation set.
+Our rank analysis of 24-layer GPT-2 medium, 36-layer GPT-2 large, and 48-layer GPT-2 xlarge models reveals that attention matrices in many deeper layers collapse to near rank-1.
 In summary, our key contributions are as follows:
@@
 We empirically investigate attention degeneration in standard LLM settings.
-Our analysis shows that rank-collapsed attention matrices often exhibit single-column structures, revealing a significant structural inefficiency in the attention mechanism of standard LLMs in deeper layers (see Figure 1 and Figure 2).
+Our analysis shows that rank-collapsed attention matrices often exhibit single-column structures, revealing a significant structural inefficiency in the attention mechanism of standard LLMs in deeper layers (refer Figure 1).
   2. 2.
@@
 The initialization is entirely zero-shot.
-We validate the effectiveness of $\mathsf{Inheritune}$ through comprehensive experiments using GPT-2 xlarge (1.5B), GPT-2 large (770M), and GPT-2 medium (355M) models, trained on the OpenWebText dataset with 9B tokens (with data repetition) and the FineWeb_edu dataset with 100B tokens (without data repetition).
+We validate the effectiveness of $\mathsf{Inheritune}$ through comprehensive experiments using GPT-2 xlarge (1.5B), GPT-2 large (770M), and GPT-2 medium (355M) models, trained on the OpenWebText and the FineWeb_edu dataset.
   3. 3.
     Evaluation Against Multiple Baselines.
-Models derived using $\mathsf{Inheritune}$ consistently outperform various baselines, including much larger models trained from scratch (refer Table 1), model initialization and efficient training baselines (refer Table 2), and models trained using two knowledge distillation techniques (refer Figure 3).
-In settings where training tokens are not repeated, we observe similar trends (refer Figure 4).
-L4 H2
-L7 H3
-L20 H2
-L22 H4
-24 layer GPT2 medium model trained from scratch.
-L4 H2
-L7 H3
-L11 H2
-L15 H1
-Our 16-layer GPT2 medium model trained using $\mathsf{Inheritune}$.
-Figure 2: Inheritune preserves effective attention patterns in smaller models.
-Comparison of attention patterns across layers (L) and heads (H) in two GPT2-medium models: (top) 24-layer model trained from scratch, (bottom) 16-layer model trained with $\mathsf{Inheritune}$.
-Attention maps are averaged over three randomly selected strings, with 40 tokens each from the validation.
-Darker colors indicate higher attention scores.
-$\mathsf{Inheritune}$ maintains focused attention even in deeper layers, contrasting with the uniform patterns in the standard model’s later layers.
-2 Attention Degeneration in Standard Decoder-Style LLMs
+Models derived using $\mathsf{Inheritune}$ consistently outperform various baselines, including much larger models trained from scratch (refer Figure 5 and Figure 6), model initialization and efficient training baselines (refer Table 1), and models trained using two knowledge distillation techniques (refer Figure 7).
+2 Attention Collapse and the Emergence of Lazy Layers in LLMs
 Preliminaries:
 A vanilla transformer-based model consists of $L$ transformer blocks (layers).
-The model operates on an input sequence $X\in\mathbb{R}^{T\times d}$, where $T$ denotes the sequence length (number of tokens), and $d$ represents the embedding dimension or hidden size of the model.
-The output of each layer $l$ is denoted as $X^{(l)}\in\mathbb{R}^{T\times d}$.
+The model operates on an input sequence $X\in\mathbb{R}^{T\times e}$, where $T$ denotes the sequence length (number of tokens), and $e$ represents the embedding dimension or model hidden size.
+The output of each layer $l$ is denoted as $X^{(l)}\in\mathbb{R}^{T\times e}$.
 Each transformer block primarily consists of two sub-layers: a self-attention block and a position-wise feed-forward network (FFN).
-The self-attention mechanism enables the model to weigh the relevance of different tokens in the sequence relative to each other.
-Specifically, for a single attention head, the attention computation is defined as
-$\text{Attention}(Q,K,V)=\underbrace{\text{softmax}\left(\frac{QK^{\top}}{\sqrt{d_{k}}}\right)}_{\textbf{Attention matrix: $A(X)$}}V$ (1)
+The self-attention mechanism enables the model to weight the relevance of different tokens in the sequence relative to each other.
+Specifically, for a single attention head, the attention computation is defined as $\text{Attention}(Q,K,V)=\underbrace{\text{softmax}\left(\frac{QK^{\top}}{\sqrt{d_{k}}}\right)}_{\textbf{Attention matrix: $A(X)$}}V$
 where the queries $Q=XW_{Q}$,  keys $K=XW_{K}$,  and values $V=XW_{V}$ are linear transformations of the input $X$.
@@
 The attention matrix $A(X)$ is then used to compute a weighted sum of the value vectors.
+Attention rank collapse refers to the phenomenon where the attention matrices $A(X)$ of individual heads in many layers of transformer-based language models lose their expressive capacity, converging towards lower effective rank structures.
+Specifically, the effective rank of attention matrices significantly reduces, often approaching rank-1, limiting the model’s ability to meaningfully differentiate between token interactions across positions in the sequence.
 Previous research by Dong et al.
@@
 (2023) has shown that in self-attention networks (SANs) without residual connections and feed-forward networks (FFNs), the rank of an attention matrix converges to rank-1 doubly exponentially with respect to the depth of the model.
-This phenomenon, known as the rank collapse of attention matrices, results in a loss of expressive power as the attention mechanism attends to all tokens uniformly.
+This phenomenon, known as rank collapse of attention matrices, results in a loss of expressive power as the attention mechanism attends to all tokens uniformly.
 Noci et al.
 (2022) showed that even with residual connections (without layernorm) attention matrices can still lose rank in deeper layers if the residual connections are not scaled by $1/\sqrt{L}$.
-Interestingly, they also linked the rank collapse to vanishing gradients of the keys and queries in deeper layers, which affects the overall trainability of the transformer-based models.
-However, these findings do not directly apply to the standard LLMs, as transformer blocks in these models include residual connections, layer norms, and FFNs, which are expected to mitigate both rank collapse and the vanishing gradient problem.
+Interestingly they also linked the rank collapse to vanishing gradients of the keys and queries in deeper layers which affects the overall trainability of the transformer based models.
+However, these findings do not directly apply to the standard LLMs, as transformer blocks in these models include residual connections, layernorms and FFNs, which are expected to mitigate both rank collapse and the vanishing gradient problem.
 Approximate Rank Computation of Attention Matrices
-In this paper, we deeply analyzed the structure of attention matrices to diagnose the presence of rank collapse or similar phenomena in standard transformer-based LLMs using GPT-2 models.
-For our first analysis, we compute the approximate rank (referred to as rank hereafter) of $A(X)$ for all attention heads within each layer.
-Formally, we began by computing the Singular Value Decomposition (SVD) of $A(X)=U\Sigma V^{\top}$, where the diagonal entries $\{\sigma_{i}\}_{i=1}^{T}$ of $\Sigma$ represent the singular values, quantifying the variance captured by each corresponding singular vector of $A(X)$.
-To determine the minimal number of singular values required to capture 90% of the total variance, we solved: $k^{*}=\min\left\{k\in\{1,2,\ldots,T\}\mid\frac{\sum_{i=1}^{k}\sigma_{i}^{2}}{\sum_{j=1}^{T}\sigma_{j}^{2}}\geq 0.90\right\}$.
-Here, $k^{*}$ is the approximate rank of $A(X)$ computed using the explained variance method.
-Dominant Single-Column Structure in Attention Matrices
-We further investigated the dominant structure of these rank-1 attention matrices and observed that, on average, many of these matrices have their mass concentrated in a single column.
-This intrinsic structure can be viewed as a special case of rank-1 attention matrices.
-To quantify this, we computed the proportion of the matrix mass contributed by each column $j$ of $A(X)$ by computing $\frac{\|A_{\cdot,j}\|_{2}^{2}}{\|A(X)\|_{F}^{2}}$, where $A_{\cdot,j}$ denotes the $j$-th column of $A(X)$, $\|A_{\cdot,j}\|_{2}$ is the $\ell_{2}$-norm of that column, and $\|A(X)\|_{F}$ is the Frobenius norm of $A(X)$.
-Next, to determine the minimal number of columns required to capture 90% of the total mass, we solved; $m^{*}=\min\left\{m\in\{1,2,\ldots,T\}\mid\sum_{j=1}^{m}\frac{\|A_{\cdot,j}\|_{2}^{2}}{\|A(X)\|_{F}^{2}}\geq 0.90\right\}$.
-Here, $m^{*}$ represents the minimal number of columns needed for the cumulative column mass ratios to reach or exceed 90%.
-Degeneration of Attention Matrices in GPT-2 Models
-In Figure 1, we present the layer-wise analysis of rank and mass.
-For this analysis, we computed $A(X)$ using 100 randomly selected samples from the validation set of OpenWebText valset with 4M tokens, each with a sequence length of $T=100$ tokens, across all attention heads within each layer.
-Next, we computed the rank with a 90% variance threshold, and for every layer, we chose the maximum rank across all the heads.
-In Figure 1(a) and 1(d) we plotted maximum rank as a function of layers for 100 runs with mean and standard deviation (std); it is quite evident that many deeper layers exhibit all rank-1 attention matrices.
-We highlight that this rank collapse occurs for both the GPT-2 medium and GPT-2 Large models with skip connections and FFNs, extending the findings of Dong et al.
+To assess the presence and severity of rank collapse within standard decoder style transformer architectures (e.g., GPT-2, LLaMA etc.), we utilize singular value decomposition (SVD) for each attention matrix $A(X)=U\Sigma V^{\top},$ where $\Sigma$ is a diagonal matrix containing singular values $\sigma_{1}\geq\sigma_{2}\geq\dots\geq\sigma_{T}\geq 0$.
+The approximate rank (referred to as rank hereafter) of an attention matrix, parameterized by a variance threshold $\tau$, is formally computed as:
+$k^{*}=\min\left\{k\in\{1,2,\dots,T\}\mid\frac{\sum_{i=1}^{k}\sigma_{i}^{2}}{\sum_{j=1}^{T}\sigma_{j}^{2}}\geq\tau\right\},$
+where $\tau\in(0,1)$ represents the proportion of variance that must be captured by the top $k$ singular values.
+A lower value of $k^{*}$ indicates stronger rank collapse.
+In this work, we set $\tau=0.90$.
+A lower $k^{*}$ indicates stronger rank collapse.
+In Figure 1, we present the layer-wise analysis of rank of GPT-2 models.
+For this analysis, we computed $A(X)$ using $N=100$ sequences selected at random from the validation set of OpenWebText with 4M tokens, each with a sequence length of $T=100$ tokens across all attention heads within each layer.
+We then define the average approximate rank for each head and layer as $\text{Rank}^{(h,l)}=\frac{1}{N}\sum_{n=1}^{N}k^{*}_{n,h,l}$.
+Subsequently, we aggregate this metric per layer by taking the maximum rank across heads: $\mathrm{MaxRank}^{(l)}=\max_{h}\{\mathrm{Rank}^{(h,l)}\}$.
+As demonstrated in Figure 1, $\text{MaxRank}^{(l)}$ reveals that many deeper layers exhibit attention matrices that are predominantly near rank-1.
+We highlight that this rank collapse occurs in GPT-2 medium, Large, and xLarge models—which are standard architectures incorporating skip connections and FFNs—thereby extending the findings of Dong et al.
 (2021) and Noci et al.
-(2022) to standard LLM architectures.
-Next, based on our mass analysis, we demonstrate that most of the rank collapsed attention matrices are also single-column matrices as depicted in Figure 1(b) and Figure 1(e).
-We follow the analysis protocol as discussed in the rank analysis.
-This degeneration of attention matrices in deeper layers provides quantitative evidence for the existence of lazy layers.
-Specifically, we observe that some deeper layers exhibit complete degeneration of all attention matrices across all attention heads, indicating reduced performance and less effective token mixing.
-How much transferable knowledge do lazy layers hold compared to their early counterparts?
-To answer this question, we initialized a 12-layer GPT-2 medium variant [^1] A variant shares the same configurations as the parent model but has fewer layers.
-and an 18-layer variant of GPT-2 large using lazy layers extracted from pre-trained 24-layer GPT-2 medium and 36-layer GPT-2 large models.
-These pre-trained models are trained on the OpenWebText-9B dataset for 100K steps.
-We then fine-tuned these GPT-2 variants on the same dataset for an additional 10K steps.
-For comparison, we conducted two baseline experiments in which the GPT-2 variants were initialized either with the first few transformer blocks or with random initialization.
-As shown in Figures 1(c) and 1(f), models initialized with lazy layers demonstrate poor transferability, performing similarly to models with random initialization.
-This provides additional evidence that lazy layers with fully degenerate attention fail to learn meaningful representations.
-2.1 Attention Pattern Visualization
-To provide further evidence of lazy layers and provide a preview of our solution, we visualized attention patterns across various layers of a vanilla 24-layer GPT-2 medium model.
-Fig.
-2 shows the attention patterns for both a vanilla 24-layer model trained from scratch and a 16-layer model trained using our proposed method, $\mathsf{Inheritune}$.
-We visualized full attention and not causal attention for the sake of better visualization.
-In practice, GPT-2 models compute causal attention.
-We computed these attention matrices using randomly selected strings from the val set of OpenWebText and took 40 tokens averaged over three runs.
+(2022) to these widely-used LLMs.
+The rank analysis is quite robust to the dataset in Figure 10, we performed the same analysis using the same models on held out data from Fineweb_edu dataset (Penedo et al., 2024).
+Overall, the degeneration of attention matrices in deeper layers provides quantitative evidence for the existence of lazy layers.
+Specifically, we observe that some deeper layers exhibit a near-complete rank collapse of attention matrices across all heads, suggesting potentially reduced representational capacity and less effective token mixing in these layers.
+2.1 The Functional Ineffectiveness of Lazy Layers
+Having identified lazy layers, we investigate their practical utility: Do these structurally degenerated layers retain transferable knowledge, or are they functionally impaired?
+Our experiments suggest the latter.
+(a) Rank profile (GPT-2 Small 12L).
+(b) Performance of 4L variants.
+Figure 2: Higher-rank layers transfer better.
+(Left, 2(a)) Layer-wise $\mathrm{MaxRank}^{(l)}$ of a pre-trained 12L GPT-2 Small.
+(Right, 2(b)) Validation loss of 4L variants initialized with potent blocks (AvgRank $\approx 8.4-9.5$) vs.
+a lazy block (AvgRank $\approx 1.2$) or random weights, after 100K steps.
+Lazy block initialization mirrors random.
+In the first set of experiments, we trained a vanilla GPT2-small (125M) model with 12 layers for 100K steps on the OpenWebText dataset.
+We then performed the rank analysis described earlier, with results presented in Figure 2.
+Specifically, we aggregated the approximate ranks over groups of contiguous layers using $\text{AvgRank}=\frac{1}{L}\sum_{l=1}^{L}\text{MaxRank}^{(l)}$, where $L$ is the number of layers in each group.
+Subsequently, we trained three GPT2-small variants [^1] A variant shares the same configurations as the parent model but has fewer layers.
+for 100K steps, each initialized with a different contiguous block of four layers from the trained vanilla GPT2-small model: (a) layers 1–4, with $\text{AvgRank}=8.40$; (b) layers 5–8, with $\text{AvgRank}=9.48$; and (c) layers 9–12, with $\text{AvgRank}=1.22$.
+The last model is initialized with lazy layers.
+For comparison, we also trained another GPT2-small variant with random initialization for 100K steps.
+All models were trained on the OpenWebText dataset.
+As shown in Figure 2, the model initialized with layers from the vanilla GPT2-small model having higher AvgRank demonstrated the best performance (i.e., lowest final validation loss).
+Additionally, we observed that the model initialized with lazy layers performed very similarly to the model with random initialization suggesting that lazy layers contain minimal transferable knowledge.
+The key results are also summarized in Table 4.
+(a) GPT2-medium (12-layer) variants.
+(b) GPT2-large (16-layer) variants.
+Figure 3: When initializing 12-layer and 16-layer variants of GPT2-medium and GPT2-large with deeper (lazy) layers showing degenerated attention, performance is comparable to random initialization.
+In contrast, early-layer initialization leads to significantly better convergence and generalization.
+For the second set of experiments we utilized larger models namely GPT-2 medium and GPT-2 large both similarly trained for 100K steps using OpenWebText dataset.
+Here we initialized a 12-layer GPT-2 medium variant and an 18-layer variant of GPT-2 large using lazy layers extracted from pre-trained 24-layer GPT-2 medium and 36-layer GPT-2 large models.
+We then trained these GPT-2 variants on the same dataset for 10K steps.
+For comparison, we conducted two baseline experiments where the GPT-2 variants were initialized either with the first half of transformer layers (potent layers with high AvgRank) and with random initialization.
+As shown in Figure 3, models initialized with lazy layers demonstrate poor transferability, performing similarly to models with random initialization.
+This provide additional evidence that lazy layers with fully degenerated attention, fails to learn meaningful representations.
+3 Inheritune: Our Proposed Training Recipe
+This section provides a detailed description of our method, key implementation considerations, and how it addresses the inefficiencies present in current architectures.
+  1. 1.
+    Inherit: Initialize $\mathcal{M}_{\text{tgt}}$ with the first $l$ out of $L$ layers of $\mathcal{M}_{\text{ref}}$, including prediction head, and token embeddings.
+  2. 2.
+    Train: Train $\mathcal{M}_{\text{tgt}}$ for $\mathsf{T}$ steps on $\mathcal{D}_{\text{train}}$ and evaluate on $\mathcal{D}_{\text{val}}$.
+  3. 3.
+    Grow: If needed, increase $\mathcal{M}_{\text{tgt}}$’s size by adding next few contiguous layers and repeat steps 1-2 until desired performance is achieved.
+As previously established, we have identified the problem of attention degeneration and its connection to lazy layers, highlighting specific inefficiencies in pre-trained LLMs.
+In this work, we transform this challenge into an opportunity to create smaller base language models, which we refer to as target models $\mathcal{M}_{\text{tgt}}$, that achieve comparable performance with similar or lower validation loss compared to their larger, less efficient counterparts, which we term reference models $\mathcal{M}_{\text{ref}}$.
 Algorithm 1 Inheritune: Training Recipe for Small Language Models
-1:Reference model $\mathcal{M}_{\text{ref}}$ with $k$ layers, datasets $\mathcal{D}_{\text{train}}$ and $\mathcal{D}_{\text{val}}$, steps $\mathsf{T}$
-2:Initialize $\mathcal{M}_{\text{tgt}}$ with first $n=k/2$ layers from $\mathcal{M}_{\text{ref}}$
-3:Train $\mathcal{M}_{\text{tgt}}$ on $\mathcal{D}_{\text{train}}$ for $\mathsf{T}$ steps
-4:while $\mathcal{M}_{\text{tgt}}$ performance < $\mathcal{M}_{\text{ref}}$ performance on $\mathcal{D}_{\text{val}}$ do
-5:     Grow $\mathcal{M}_{\text{tgt}}$ by inheriting additional layers
-6:     Train $\mathcal{M}_{\text{tgt}}$ for $\mathsf{T}$ steps
-7:end while
-8:return Optimized model $\mathcal{M}_{\text{tgt}}$
-In the 24-layer model trained from scratch (top row of Fig.
-2), we observe a clear progression in attention patterns.
-The early layers (L4 and L7) exhibit structured patterns with a mix of local and global attention Gong et al.
-(2019); Beltagy et al.
-(2020); Chen et al.
-(2021).
-In contrast, the deeper layers (L20 and L22) display more uniform patterns, indicating a loss of focus.
-This uniformity is a hallmark of lazy layers, where the attention mechanism loses its ability to selectively focus on specific relevant tokens.
-In contrast, our 16-layer model trained with $\mathsf{Inheritune}$ (bottom row) demonstrates more focused and effective attention patterns, even in its later layers (L11 and L15).
-This striking difference suggests that our method makes the model more attentive and addresses attention degeneration, potentially leading to more efficient models in a compact size (also refer to Figure 9).
-We will discuss these results in more detail after introducing our method, but this preview underscores the promise of our approach.
-3 Inheritune: Our Proposed Training Recipe
-This section offers a detailed description of our method, key implementation considerations, and how it addresses the inefficiencies present in current architectures.
-Recall that we have established the issue of attention degeneration with two motivating examples, highlighting specific inefficiencies in pre-trained LLMs.
-In this paper, we turn this challenge into an opportunity to create smaller base LMs that are equally performant, achieving similar or lower validation loss compared to their larger, less efficient counterparts, which we refer to as reference models.
-Our proposed solution builds on two key insights: (1) the early layers of deep LLMs provide effective model initialization, and (2) multiple lazy layers can be collapsed into fewer layers and re-trained to improve the model capacity.
+1:Reference model $\mathcal{M}_{\text{ref}}$ with $L$ layers, datasets $\mathcal{D}_{\text{train}}$ and $\mathcal{D}_{\text{val}}$, steps $\mathsf{T}$
+2:Copy embedding layer and LM head from $\mathcal{M}_{\text{ref}}$ to $\mathcal{M}_{\text{tgt}}$
+3:Select $l$ early contiguous layers from $\mathcal{M}_{\text{ref}}$ with high AvgRank
+4:Initialize $\mathcal{M}_{\text{tgt}}$ with selected layers between embeddings and LM head
+5:Train $\mathcal{M}_{\text{tgt}}$ on $\mathcal{D}_{\text{train}}$ for $\mathsf{T}$ steps
+6:while $\mathcal{M}_{\text{tgt}}$ performance < $\mathcal{M}_{\text{ref}}$ performance on $\mathcal{D}_{\text{val}}$ do
+7:     Grow $\mathcal{M}_{\text{tgt}}$ by inheriting additional layers
+8:     Train $\mathcal{M}_{\text{tgt}}$ for $\mathsf{T}$ steps
+9:end while
+10:return Optimized model $\mathcal{M}_{\text{tgt}}$
+Our proposed solution builds on two key insights: (1) the early layers of deep LLMs contain a higher concentration of potent layers with high AvgRank values, making them suitable for effective model initialization, and (2) lazy layers can be identified, removed, or utilized in smaller numbers, then subsequently re-trained to improve overall model capacity.
+Figure 4: Overview of the $\mathsf{Inheritune}$ training recipe using a 24-Layer GPT-2 medium model example.
+A smaller target model is initialized using early layers from a larger, pre-trained reference model.
+The target model goes multiple rounds of training while inheriting more early layers until it matches the reference model.
+The intensity of the red color in layers correlates with $\mathrm{MaxRank}^{(l)}$.
 Setup:
 We split the dataset into a training set $\mathcal{D}_{\text{train}}$ and a validation subset $\mathcal{D}_{\text{val}}$.
-Next, we assume that there exists a pre-trained reference model $\mathcal{M}_{\text{ref}}$, comprising $k$ layers, represented by $\mathsf{W}_{\text{ref}}=\{\mathsf{W}_{0},\mathsf{W}_{1},\ldots,\mathsf{W}_{k-1}\}$ trained with $\mathcal{D}_{\text{train}}$ for $\mathsf{T}$ steps.
+Next, we assume that there exists a pre-trained reference model $\mathcal{M}_{\text{ref}}$, comprising $L$ layers, represented by $\mathsf{W}_{\text{ref}}=\{\mathsf{W}_{0},\mathsf{W}_{1},\ldots,\mathsf{W}_{L-1}\}$ trained with $\mathcal{D}_{\text{train}}$ for $\mathsf{T}$ steps.
 We want to train a smaller model $\mathcal{M}_{\text{tgt}}$ with the same or better validation loss (lower is better) compared to its larger counterpart $\mathcal{M}_{\text{ref}}$.
@@
 The $\mathsf{Inheritune}$ process consists of three main steps, which we present below and formalize in Algorithm 1:
-  1. 1.
-    Inherit: Initialize $\mathcal{M}_{\text{tgt}}$ with the first $n=k/2$ layers of $\mathcal{M}_{\text{ref}}$, including weights, prediction head, and token embeddings.
-  2. 2.
-    Train: Train $\mathcal{M}_{\text{tgt}}$ for $\mathsf{T}$ steps on $\mathcal{D}_{\text{train}}$ and evaluate on $\mathcal{D}_{\text{val}}$.
-  3. 3.
-    Grow: If needed, increase $\mathcal{M}_{\text{tgt}}$’s size and repeat steps 1-2 until the desired performance is achieved.
 With our method now formally described, we turn to empirical validation.
 In the following sections, we present comprehensive results demonstrating $\mathsf{Inheritune}$’s effectiveness across various scenarios, including different model sizes and data regimes.
-In addition, we conducted an in-depth ablation study to analyze the impact of initialization on performance, providing information on the adaptability of our approach.
+In addition, we conducted an in-depth ablation study to analyze the impact of initialization on performance, providing insights into the adaptability of our approach.
 4 Experiments
-We evaluated $\mathsf{Inheritune}$ through a series of comprehensive experiments using GPT-2 xlarge (1.5B), GPT-2 large (770M) and GPT-2 medium (355M) models, Radford et al.
-(2019) pre-trained on the 9B tokens OpenWebText dataset (Gokaslan & Cohen, 2019).
-These models are trained with data repetition, meaning data is randomly sampled with replacement during batch creation.
-This experimental setup is adapted from Liu et al.
+We evaluate $\mathsf{Inheritune}$ through a comprehensive set of experiments using several GPT-2 models: 48-layer GPT-2 xlarge (1.5B), 36-layer GPT-2 large (770M), 32-layer GPT-2 large (668M) and 24-layer GPT-2 medium (355M) models Radford et al.
+(2019).
+Table 11 provides detailed specifications of all model configurations used.
+We use two training datasets in our experiments: OpenWebText Gokaslan & Cohen (2019) and Fineweb_edu Penedo et al.
+(2024).
+Our experimental setup closely follows prior works Liu et al.
 (2023); Sanyal et al.
 (2024).
-For evaluation, we compare model(s) trained with $\mathsf{Inheritune}$ with baselines from three key settings: a) baseline models trained from scratch with random initialization, b) baseline models trained using various zero-shot initialization techniques, and c) baseline models trained with knowledge distillation.
-Table 10 provides detailed specifications for all models used in our experiments.
-Finally, we conduct a thorough ablation study of our initialization strategy, focusing on the 16-layer GPT-2 medium variants.
-We provide experimental details of our $\mathsf{Inheritune}$ training recipe using a GPT-2 large model as an example; a similar procedure was applied to train other models.
-Our methodology for applying $\mathsf{Inheritune}$involves the following steps:
+For models trained on OpenWebText, we report validation loss (log perplexity) on a held-out set.
+For those trained on Fineweb-Edu, we report training loss (also log perplexity).
+In addition, we conduct zero-shot downstream evaluations using the lm-evaluation-harness Gao et al.
+(2024) across five standard tasks: ARC-easy (ARCE; Clark et al.
+(2018)), LAMBADA Paperno et al.
+(2016), SciQ Welbl et al.
+(2017), HellaSwag Zellers et al.
+(2019), and PIQA Bisk et al.
+(2020).
+We also perform a thorough ablation study of our initialization strategy, focusing on 16-layer GPT-2 medium variant.
+We provide experimental details our $\mathsf{Inheritune}$ training recipe using a GPT-2 medium model as an example; similar procedure was applied to train other models.
+A visualization of the training recipe is presented in Figure 4.
+Our recipe for applying $\mathsf{Inheritune}$ involves the following steps.
   1. 1.
-    Reference Model: We train the full 36-layer GPT-2 Large model on $\mathcal{D}_{\text{train}}$ for 100K steps and evaluate its validation loss ( log-perplexity) on $\mathcal{D}_{\text{val}}$.
+    Reference Model: We train the full 24-layer GPT-2 medium model on $\mathcal{D}_{\text{train}}$ for 100K steps and evaluate its validation loss ( log-perplexity) on $\mathcal{D}_{\text{val}}$.
 This establishes our benchmark validation loss.
   2. 2.
-    Model initialization We initialize an 18-layer model ($n=k/2$) using the trained 36-layer model as a reference.
+    Model initialization: We initialize an 12-layer model ($l=L/2$) using the reference model.
   3. 3.
-    Training and Evaluation: We train the 18-layer model on $\mathcal{D}_{\text{train}}$ for $T$ steps and evaluate its validation loss.
+    Training and Evaluation: We train the 12-layer model on $\mathcal{D}_{\text{train}}$ for $\mathsf{T}$=100K steps and evaluate its validation loss.
   4. 4.
-    Iterative Refinement: If the smaller model’s performance is inferior, we incrementally increase its size by two layers and repeat steps 2-3 until we achieve parity with the reference model’s validation loss.
-Models Layers Initialization Steps Pre-train Downstream (0-shot)
-Val loss ($\downarrow$) Wikitext ($\downarrow$) Lambada
-GPT-2 Medium 24 rand init 100K 2.81 31.93 36.54
-16 rand init 100K 2.86 33.67 34.60
-16 rand init 200K 2.83
-12 Ours 100K 2.87
-14 Ours 100K 2.84
-Final Model $\longrightarrow$ 16 Ours 100K 2.81 32.04 35.96
-GPT-2 Large 36 rand init 100K 2.85 34.84 34.14
-18 rand init 100K 2.97 37.63 30.97
-18 rand init 200K 2.84
-18 Ours 100K 2.80 35.38 34.64
-GPT-2 xLarge 48 rand init 100K 2.65 25.45 39.90
-24 rand init 100K 2.69 28.32 38.46
-24 rand init 200K 2.62
-24 Ours 100K 2.64 25.52 43.30
-Table 1: Inheritune achieves superior performance with reduced model size.
-Comparison of $\mathsf{Inheritune}$-trained models (24-layer GPT-2 xLarge, 18-layer GPT-2 Large, 16-layer GPT-2 Medium) against full-sized counterparts and extended training baselines.
-Metrics include pre-training validation loss ($\downarrow$), zero-shot Wikitext ($\downarrow$) and Lambada performance.
-Note: GPT-2 xLarge and Large took one round of training; GPT-2 Medium took three rounds.
-Baselines trained from scratch (rand init.) :
-We compare our $\mathsf{Inheritune}$-derived model against much larger GPT-2 reference models trained from scratch for the same number of steps and similar-sized models trained from scratch for both the same and double the number of training steps.
-Baselines trained with various model initialization and efficient training techniques.
-Here, we compare our model derived using $\mathsf{Inheritune}$ to models of similar size trained with various zero-shot model initialization and efficient training techniques such as stacking, hybrid stacking, and half-width initialization.
-We explain these baseline training recipes using GPT-2 large and its variants as an example and apply the same process to other models.
-  1. 1.
-    Stacking Gong et al.
+    Iterative Refinement: If the smaller model’s performance is inferior to the reference model, then we incrementally increase its size by two layers and repeat steps 2-3 until we achieve parity with the reference model’s validation loss.
+We choose $l=L/2$ as the starting point and increase the model size by two layers in each round across all our experiments, aiming to minimize the number of training rounds.
+In principle, $\mathsf{Inheritune}$ should generalize to other hyperparameter choices as well.
+(a) GPT2-Medium
+(b) GPT2-Large
+(c) GPT2-xLarge
+Figure 5: Models derived using Inheritune converge faster and match the final validation loss of the full-sized model, despite having much fewer layers.
+Comparison of $\mathsf{Inheritune}$-trained models (24-layer GPT-2 xLarge variant, 18-layer GPT-2 Large variant, 16-layer GPT-2 Medium variant) against their full-sized counterparts and same sized variants trained from scratch.
+All models are trained for 100K steps using OpenWebText data with data repetition.
+(a) GPT2-Medium
+(b) GPT2-Large
+Figure 6: Models derived using $\mathsf{Inheritune}$ converge faster and match the final validation loss and average downstream performance (Perf.) across 5 datasets of the full-sized model, while outperforming same-sized models.
+(Plots) Validation loss curves for our 16-layer variants versus full-sized models.
+(Table) Zero-shot downstream performance comparison.
+All models are trained for 100K steps using FineWeb_edu data without data repetition.
+Models Recipe Layers Perf. Medium rand init 24 47.74 rand init 16 46.84 Ours 16 48.08 Large rand init 32 49.44 rand init 16 48.17 Ours 16 49.75
+Baselines.
+We compare GPT-2 model variants (variants are models with lesser layers compared to their vanilla configuration) trained using $\mathsf{Inheritune}$ against two types of baselines: (1) larger reference models with more layers, and (2) similar-sized models with the same number of layers trained from scratch (randomly initialized).
+Additionally, we compare $\mathsf{Inheritune}$ against efficient training baselines that leverage pre-trained model initialization.
+These include stacking Gong et al.
 (2019); J.
 Reddi et al.
-(2023) is a model initialization and efficient (stage-wise) training recipe.
-We trained a 9-layer GPT-2 large variant from scratch for 100K steps, then expanded the model to 18 layers by copying the weights from layers 0-8 to layers 9-17.
-Finally, we re-trained this new 18-layer GPT-2 large variant, using stacking initialization for an additional 100K steps.
-  2. 2.
-    Hybrid stacking: Hybrid stacking is stacking that utilizes a large pre-trained reference model for initialization instead of using its own pre-trained weights.
-We took the weights of layers 0-8 from the reference 36-layer GPT-2 large model and expanded it to an 18-layer model by copying the weights to layers 0-17.
-We then trained this new 18-layer GPT-2 variant for 100K steps.
-  3. 3.
-    Half width: We initialized the baseline GPT-2 large variant across the width dimension and preserved the entire depth.
-We copied the weights of the first half of the attention heads (0-9) and MLPs of the GPT-2 large reference model into the baseline GPT-2 variant with half the width but all layers.
-Baselines trained with Knowledge Distillation
-As a baseline, we first apply logit-based knowledge distillation (KD) Hinton et al.
-(2015) to train a 16-layer GPT-2 medium variant (student) initialized randomly.
-For the second baseline, we use a DistillBERT-style approach Sanh et al.
-(2019), where the 0-11 layers of the student model are initialized with each alternate block of its teacher, and the remaining 4 blocks are initialized using layers 18, 19, 20, and 21 of the teacher.
-Both baselines are trained for 50K steps, using a vanilla 24-layer GPT-2 medium model as the teacher (our reference model).
-Additional results with 12-layer GPT-2 medium variant(s) trained from scratch, using KD baseline and our recipe can be found in Figure 8.
+(2023), hybrid stacking (which initializes the model using layers from the reference model), and half-width initialization (where all layers are retained, but the hidden dimension and number of attention heads are halved and initialized from the reference model).
+Finally we briefly compare $\mathsf{Inheritune}$ with knowledge distillation (Hinton et al., 2015).
+Detailed descriptions of these baselines are provided in the supplementary material Section B.
 Models Layers Recipe Steps Pre-train Val loss ($\downarrow$)
@@
 24 Ours 100K 2.64
-Table 2: Inheritune outperforms baseline zero-shot initialization and efficient training techniques.
+Table 1: $\mathsf{Inheritune}$ outperforms baseline zero-shot initialization and efficient training techniques.
 Comparison of pre-training validation loss for GPT-2 xLarge, GPT-2 Large and GPT-2 Medium variants.
@@
 4.1 Results and Discussions
-Models trained with Inheritune outperforms much larger models trained from scratch.
-We present our main results in Table 1.
-Our 24-layer, 18-layer, and 16-layer variants derived using $\mathsf{Inheritune}$ from the vanilla 48-layer GPT-2 xlarge, 36-layer GPT-2 large, and 24-layer GPT-2 medium achieve comparable or lower validation losses than their full-sized counterparts when trained for the same number of steps (100K steps).
-Our GPT-2 xlarge and GPT-2 large variants undergo one round of $\mathsf{Inheritune}$ training, while for GPT-2 medium, we perform three rounds of training with 12, 14, and 16 layers.
-We also evaluated all models on two next-word prediction downstream tasks in a zero-shot setting using the Wikitext Merity et al.
-(2016) and Lambada Paperno et al.
-(2016) datasets.
-The downstream performance of our GPT-2 models derived using the $\mathsf{Inheritune}$ recipe matches their much larger reference models.
-From the convergence perspective, some prior works have made connections between over-parameterization and faster convergence Bengio et al.
+Models trained with $\mathsf{Inheritune}$ outperform much larger models trained from scratch.
+We present our main results in Figure 5.
+The 24-layer, 18-layer, and 16-layer variants derived using $\mathsf{Inheritune}$ from the vanilla 48-layer GPT-2 xlarge, 36-layer GPT-2 large, and 24-layer GPT-2 medium, respectively, achieve comparable or lower validation losses than both their full-sized counterparts and same-sized models trained from scratch, when trained for the same number of steps (100K).
+Our GPT-2 xlarge and GPT-2 large variants require a single round of $\mathsf{Inheritune}$ training, while the GPT-2 medium variant undergoes three rounds with 12-, 14-, and 16-layer configurations.
+From a convergence perspective, prior work has linked overparameterization to faster convergence Bengio et al.
 (2005); Vaswani et al.
 (2018).
-In supplementary Figure 5, we show that the small LMs derived with $\mathsf{Inheritune}$, although smaller compared to their reference models, still converge as fast as their large-size reference models.
-Models trained with Inheritune outperform same-sized models trained from scratch.
-Table 1 demonstrates that GPT-2 variants trained with $\mathsf{Inheritune}$ outperform their same-sized counterparts trained from scratch, both when trained for the same number of steps and even when trained for double the steps (200K).
-This result underscores the efficiency of our approach.
-The only exception is the 24-layer GPT-2 xlarge variant, which surpasses both our model and the full-size model when trained for 200K steps.
-Models trained with Inheritune outperform all zero-shot model initialization and efficient training baselines.
+Interestingly, we find that smaller models derived using $\mathsf{Inheritune}$ converge just as fast as their larger counterparts.
+A detailed summary of these results is provided in the appendix Table 3.
+We conducted additional training experiments without data repetition, following standard LLM pre-training practices as discussed in Touvron et al.
+(2023); Biderman et al.
+(2023) using the high-quality Fineweb_edu dataset.
+We trained a 32-layer GPT-2 large (668M) and a 24-layer GPT-2 medium (355M) reference model from scratch.
+Next, we trained two 16-layer variants: one derived from GPT-2 large and the other from GPT-2 medium, using their respective reference models following $\mathsf{Inheritune}$.
+For comparison, we also trained 16-layer baseline models from scratch.
+All models were trained for 100K steps.
+We observe thematically consistent results: as shown in Figure 6, the 16-layer variants trained with $\mathsf{Inheritune}$ consistently match the performance of their full-sized counterparts and outperform same-sized baselines, both in terms of training loss and zero-shot downstream evaluation.
+Downstream results are averaged over five tasks (see table within Figure 6); the complete breakdown is provided in Table 5 placed in appendix.
+Model configurations and training hyper-parameters are detailed in the supplementary material (refer Section F).
+Figure 7: A 16-layer GPT-2 medium variant derived using $\mathsf{Inheritune}$ converges faster and generalizes better than a same-sized model trained with logit-based distillation baselines.
+Models trained with Inheritune outperform all zero-shot model initialization baselines.
 In Table 1, we compare GPT-2 xlarge, GPT-2 large, and GPT-2 medium variants trained with $\mathsf{Inheritune}$ against same-sized variants trained with stacking, hybrid, and half-width initialization baselines.
 The half-width baseline performs poorly, revealing the limitations of naive width reduction.
-Although stacking and hybrid stacking demonstrate reasonable performance, they still fall short compared to $\mathsf{Inheritune}$.
-In all cases, $\mathsf{Inheritune}$ consistently outperforms these baselines, highlighting its effectiveness as an initialization strategy.
+While stacking and hybrid stacking demonstrate reasonable performance, they still fall short compared to $\mathsf{Inheritune}$.
+Across all cases, $\mathsf{Inheritune}$ consistently outperforms these baselines, highlighting its effectiveness as an initialization strategy.
 For a detailed view of the training dynamics across all methods, refer to the training curves in supplementary Figure 14.
-Figure 3: A 16-Layer GPT-2 medium variant derived using Inheritune converges faster and generalizes better than a same-sized model trained with Logit-based distillation baselines.
-We conducted vanilla KD Hinton et al.
-(2015) and DistillBERT-style KD Sanh et al.
-(2019) with teacher initialization, using a 24-layer GPT-2 medium as the teacher for both KD baselines.
 Distillation vs Inheritune.
@@
 (2019), which leverages teacher layers for model initialization, vanilla training from scratch and $\mathsf{Inheritune}$.
-Our model demonstrates faster convergence and significantly better final generalization after 50K steps.
-Additional distillation experiments can be found in the supplementary material.
+Our model demonstrates faster convergence and significantly better final generalization after 50K training steps.
+Additional distillation experiments can be found in the supplementary materials.
 4.2 Ablations
@@
 Table 2: Impact of initializing various sub-modules within a transformer block.
-We compare the validation loss of a 16-layer GPT-2 medium variant when different sets of sub-modules are initialized with weights from the first 16 layers of a 24-layer GPT-2 medium reference model.
-All models are trained on the OpenWebText-9B dataset.
+We compare validation loss of a 16-layer GPT-2 medium variant when different sets of sub-modules are initialized with weights from the first 16 layers of a 24-layer GPT-2 medium reference model.
+All models are trained on the OpenWebText dataset.
 Key findings: (1) $\mathsf{Inheritune}$ initialization and attention + MLP initialization result in similar performance improvements; (2) layernorm initialization shows minimal impact.
@@
 Surprisingly, we also observed that initializing either the attention or mlp weights resulted in similar improvements in both convergence speed and final validation loss.
-5 Training without Data Repetition
-(a) GPT2-Large
-(b) GPT2-Medium
-Figure 4: Models derived using Inheritune without data repetition converge faster and match the final validation loss of the full-sized model despite using fewer layers.
-Additionally, the model trained using $\mathsf{Inheritune}$ demonstrates data efficiency, achieving a lower validation loss in fewer steps compared to its full-sized and half-sized counterparts until 80% of the training.
-Are the gains we observe due to the Inheritune recipe merely a consequence of over-fitting due to data repetition?
-To investigate this, we conducted additional training experiments without data repetition, following standard LLM pre-training practices as discussed in Touvron et al.
-(2023a); Biderman et al.
-(2023).
-Moreover, we utilized a high-quality pre-training dataset, Fineweb_edu Penedo et al.
-(2024), which contains 100B tokens and has been deduplicated and filtered to ensure high data quality.
-We train a 32-layer GPT-2 large (668M) and a 24-layer GPT-2 medium (355M) reference model from scratch.
-Next, we trained two 16-layer variants: one derived from GPT-2 large and the other from GPT-2 medium, using their respective reference models following Algorithm 1.
-Finally, we trained the 16-layer baseline variants from scratch for comparison.
-All these models are trained for 100K steps.
-The model configurations and training hyper-parameters can be found in supplementary Table 10.
-As shown in Figure 4, our GPT-2 variants trained using $\mathsf{Inheritune}$ consistently perform on par with their full-size counterparts and outperform their same-size counterparts in terms of training loss.
-In LLM pre-training literature, where data is not repeated, training loss has been shown to be a reliable metric Touvron et al.
-(2023a, b).
-Furthermore, we performed zero-shot downstream evaluations using lm-evaluation-harness Gao et al.
-(2024) on a variety of tasks, including ARC-easy (ARCE; Clark et al.
-(2018)), LAMBADA Paperno et al.
-(2016), SciQ Welbl et al.
-(2017), Hellaswag Zellers et al.
-(2019), and PIQA Bisk et al.
-(2020).
-As shown in Table 4, on average, the models demonstrate superior performance than the baseline models trained from scratch.
-Models Recipe Layers ARCE PIQA SciQ Hellaswag Lambada Average
-(acc) (acc) (acc) (acc norm) (acc)
-GPT-2 Medium rand init 24 51.05 61.81 74.8 30.79 20.28 47.74
-rand init 16 49.92 61.92 73.3 29.56 19.54 46.84
-Ours 16 51.26 61.81 73.8 30.55 23 48.08
-GPT-2 Large rand init 32 52.48 64.58 75.3 32.65 22.2 49.44
-rand init 16 50.34 63.11 75 30.86 21.56 48.17
-Ours 16 52.9 63.55 76.1 32.14 24.06 49.75
-Table 4: Models trained with Inheritune outperform both their larger and same-size counterparts trained from scratch on average zero-shot downstream performance.
-For evaluation we choose accuracy (acc) and normalized accuracy (acc norm) metrics following Open LLM leaderboard Beeching et al.
-(2023).
-All the models are trained with FineWeb_edu.
 5 Related Works
@@
 (2023) leading training instability.
-This also has been studied as a theoretical setup for transformer models by Noci et al.
+This also has been studied is a theoretical setup for transformer models by Noci et al.
 (2022); Barbero et al.
 (2024).
-He et al.
-(2023) address rank collapse in self-attention networks (SANs) without residual connections or layer norms, using two model initialization techniques that enable faithful signal propagation, i.e., $\Sigma_{L}$ of $A(X^{L})$ does not collapse in deeper layers.
+Recently He et al.
+(2023) address rank collapse in self-attention networks (SANs) without residual connections or layer norms, using two model initialization techniques that enable faithful signal propagation—i.e., $\Sigma_{L}$ of $A(X^{L})$ does not collapse in deeper layers.
 However, this approach significantly slows down training.
@@
 (2022) proposes scaling residual connections by $1/\sqrt{L}$, while Barbero et al.
-(2024) suggests that adding additional tokens to already long sequences of repeated tokens can help mitigate degeneration.
-In contrast to previous work, we address attention degeneration by developing smaller models that eliminate structural inefficiencies and training these models to match the performance of their larger, inefficient counterparts.
+(2024) suggest that adding additional tokens to already long sequences of repeated tokens can help mitigate degeneration.
+In contrast to prior works, we address attention degeneration by developing smaller models that eliminate structural inefficiencies and training these models to match the performance of their larger, inefficient counterparts.
 LLM training recipes and model initialization.
@@
 Reddi et al.
-(2023), which employs a stage-wise training strategy that uses weights from initial layers to initialize later layers, has been shown to be effective for LLM training both empirically Gong et al.
+(2023) employs a stage-wise training strategy that uses weights from initial layers to initialize later layers has been shown to be effective for LLM training both empirically Gong et al.
 (2019); J.
@@
 (2020); Sanh et al.
-(2019) the smaller student model is also initialized with teacher layers-though this is often done without clear explanation or intuition.
+(2019) the smaller student model is also initialized with teacher layers, though this is often done without clear explanation or intuition.
 Recent works in model initialization, such as Trockman & Kolter (2023), have studied synthetic attention patterns for initialization, primarily in vision settings.
@@
 (2024) use weight initialization for faster fine-tuning of vision models.
-In contrast, our proposed recipe focuses on creating a smaller model by eliminating specific structural inefficiency in lazy layers.
+In contrast, our proposed recipe focuses on creating smaller model by eliminating specific structural inefficiency in lazy layers.
 This distinction sets our work apart in terms of both objective and methodology.
 6 Conclusion
-In this study, we identified a structural flaw in the attention mechanism of deep decoder-style transformers, which are currently the de-facto architectural choice for LLMs.
-Specifically, we discovered that the attention matrices in many deeper layers tend to lose rank and converge into single-column matrices, resulting in what we term lazy layers.
-This phenomenon causes the model to lose attention and leads to structural inefficiencies in LLMs.
-To avoid this issue, we propose a simple training recipe $\mathsf{Inheritune}$ where we inherit a few early blocks from a larger reference model and continue training and expanding the model with more blocks until it matches the performance of the reference model.
-We evaluated our method using 48-layer, 36-layer, and 24-layer GPT-2 xlarge, GPT-2 large, and GPT-2 medium reference models.
-Through this process, we successfully developed smaller models with 24, 18, and 16 layers without losing any performance compared to their larger counterparts while utilizing the OpenWebText-9B dataset.
-We observed thematically similar results with GPT-2 large and GPT-2 medium models trained on FineWeb_edu without data repetition.
-These findings offer a novel perspective on developing efficient small base language models from existing large pre-trained models, potentially democratizing LLM pre-training for a wider community.
-Acknowledgments
-We thank Nived Rajaraman, for his helpful feedback and discussion during this work.
+In this paper, we identified a structural flaw in the attention mechanism of deep decoder-style LLMs, where many deeper layers tend to lose rank and converge into near rank-1 matrices.
+To address this, we propose $\mathsf{Inheritune}$, to train smaller models that inherits early blocks from a larger model and expands the architecture gradually, matching the performance of the reference model.
+We validated $\mathsf{Inheritune}$ on GPT-2 models of varying sizes, achieving efficient smaller models without performance loss on the OpenWebText and FineWeb_edu datasets.
 Supplementary Materials
 Contents
-  • A: Supplementary Experiments
-  • B: Additional Experiments on Low Data Regime
-  • C: Implementation Details
-Appendix A Supplementary Experiments
-We provide additional training plots for our main results discussed in Section 4.1 as shown in Figure 5 and Figure 6.
+  • A: Extended Discussion on Attention Collapse
+  • B: Baselines
+  • C: Supplementary Experiments
+  • D: Extended Discussion
+  • E: Additional Experiments on Low Data Regime
+  • F: Implementation Details
+Appendix A Extended Discussion on Attention Collapse
+A.1 Attention Mass Analysis
+Dominant Single-Column Structure in Attention Matrices.
+We further investigated the dominant structure of the rank-1 attention matrices and observed that, on an average, many of these matrices have their mass concentrated in a single column.
+This intrinsic structure can be viewed as a special case of rank-1 attention matrices.
+To quantify this, we computed the proportion of the matrix mass contributed by each column $j$ of $A(X)$ by computing $\frac{\|A_{\cdot,j}\|_{2}^{2}}{\|A(X)\|_{F}^{2}}$, where $A_{\cdot,j}$ denotes the $j$-th column of $A(X)$, $\|A_{\cdot,j}\|_{2}$ is the $\ell_{2}$-norm of that column, and $\|A(X)\|_{F}$ is the Frobenius norm of $A(X)$.
+Next we determine the minimal number of columns required to capture $\eta$ proportion of the total mass, formally computed as;
+$m^{*}=\min\left\{m\in\{1,2,\ldots,T\}\mid\sum_{j=1}^{m}\frac{\|A_{\cdot,j}\|_{2}^{2}}{\|A(X)\|_{F}^{2}}\geq\eta\right\},$
+Here $\eta\in(0,1)$ represents the cumulative column mass threshold.
+In this work, we set $\eta=0.90$.
+A lower value of $m^{*}$ implies a stronger concentration of the attention matrix mass within fewer columns, reinforcing the phenomenon attention collapse.
+This analysis provides additional quantitative evidence highlighting the reduced representational capability of attention matrices in deeper transformer layers, further supporting the identification of lazy layers.
+In Figure 8, we present the layer-wise analysis of the attention matrix mass concentration in GPT-2 models.
+For this analysis, (similar to the rank analysis), we computed $A(X)$ using $N=100$ sequences selected at random from the validation set of OpenWebText (4M tokens), each with a sequence length of $T=100$ tokens across all attention heads within each layer.
+We define the average minimal column count $m$ required to capture 90% of the attention matrix mass for each head and layer as: $m^{(h,l)}=\frac{1}{N}\sum_{n=1}^{N}m^{*}_{n,h,l}$.Subsequently, we aggregate this metric per layer by taking the average across all heads: $\text{AvgMass}^{(l)}=\frac{1}{H}\sum_{h^{\prime}=1}^{H}m^{(h^{\prime},l)}$.
+We observe that many of the rank-collapsed attention matrices in deeper layers exhibit single-column attention structures, as measured by the $\text{AvgMass}^{(l)}$ criterion.
+As shown in Figure 9, we performed a mass analysis on contemporary billion-parameter LLaMA models (Geng & Liu, 2023) and observed a similar pattern of attention degradation in the deeper layers.
+This provides concrete evidence that the phenomenon persists across a broad range of architectures at the billion-parameter scale.
+(a) GPT-2 Medium (355M) L=24, H=1024
+(b) GPT-2 Large (770M) L=36, H=1280
+(c) GPT-2 xLarge (1.5B) L=48, H=1600
+Figure 8: In decoder-style LLMs, attention matrices in deeper layers often degenerate to near single column matrices, which is a special case of near rank-1.
+We compute $\text{AvgMass}^{(l)}$ (averaged over $N=100$ randomly selected sequences each with $T=100$ tokens) for each layer $l$ using the OpenWebText validation set.
+Our mass analysis of 24-layer GPT-2 medium, 36-layer GPT-2 large, and 48-layer GPT-2 xlarge models (L:layer, H:hidden size) reveals that attention matrices in many deeper layers collapse to single column matrices on an average.
+(a) OpenLLaMA-3B
+(b) OpenLLaMA-7B
+(c) OpenLLaMA-13B
+Figure 9: The overall mass of attention matrices in billion-scale LLMs, pre-trained on trillions of tokens, tends to concentrate in fewer columns.
+This phenomenon becomes increasingly pronounced as the model size grows.
+We computed attention matrices using 100 tokens from a random subset of RedPajama with 1B tokens.
+Next, we performed 100 runs and plotted the mean and standard deviation of the mass as a function of layers for our mass analysis, respectively.
+We followed the same procedure as discussed in Section 2.
+Pre-trained checkpoints of OpenLLaMA-3B, OpenLLaMA-7B, and OpenLLaMA-13B (Geng & Liu, 2023), trained on 1T tokens from the RedPajama dataset Computer, 2023, were utilized.
+Overall, we observed that 90 % of the total mass of the attention matrices resides in fewer columns, with many attention matrices in the OpenLLaMA-13B model being single-column.
+This observation aligns closely with our analysis in Figure 1.
+A.2 Additional Analysis on Attention Rank Collapse.
+To investigate the robustness of our analysis, we performed a rank analysis on GPT-2 models trained on OpenWebText (as shown in Figure 1) using a validation set from FineWeb-Edu—a new dataset on which the models were not trained (see Figure 10).
+It is clearly observed that attention consistently loses rank, often collapsing to single-column structures across all models, thereby reinforcing the robustness of our findings.
+Next, we conducted a rank analysis on a contemporary LLaMA-3 model with 8B parameters.
+We compute $\text{Rank}^{(h,l)}=\frac{1}{N}\sum_{n=1}^{N}k^{*}_{n,h,l}$, where $N=100$ sequences are sampled from a subset of the C4 dataset (Raffel et al., 2019).
+As shown in Figure 11, we observe that nearly 50% of the attention heads (500 out of 1024 across all layers) are close to near rank-1, highlighted in red.
+This presents an interesting case: in very large modern architectures such as LLaMA-3 8B, while there may not be entire lazy layers, a substantial number of heads within many layers exhibit degeneracy.
+A.3 Attention Pattern Visualization
+To provide further evidence of lazy layers, we visualized attention patterns across various layers of a vanilla 24-layer GPT-2 medium model.
+Fig.
+12 shows the attention patterns for both a vanilla 24-layer model trained from scratch and a 16-layer model trained using our proposed method, $\mathsf{Inheritune}$.
+Note just for the sake of better visualization we visualized full attention and not causal attention, in practice GPT-2 models computes causal attention.
+We computed these attention matrices using randomly selected strings from the val set of OpenWebText and took 40 tokens averaged over 3 runs.
+In the 24-layer model trained from scratch (top row of Fig.
+12), we observe a clear progression in attention patterns.
+The early layers (L4 and L7) exhibit structured patterns with a mix of local and global attention Gong et al.
+(2019); Beltagy et al.
+(2020); Chen et al.
+(2021).
+In contrast, the deeper layers (L20 and L22) display more uniform patterns, indicating a loss of focus.
+This uniformity is a hallmark of lazy layers, where the attention mechanism loses its ability to selectively focus on specific relevant tokens.
+In contrast, our 16-layer model trained with $\mathsf{Inheritune}$ (bottom row) demonstrates more focused and effective attention patterns, even in its later layers (L11 and L15).
+This striking difference suggests that our method makes model more attentive and addresses attention degeneration, potentially leading to more efficient models in compact size (also refer Figure 17 and Figure 18).
+Additionally, in Figure 13, we present attention patterns from a 48-layer GPT-2 xLarge model, comparing attention maps from a potent (non-lazy) and a lazy layer.
+(a) Rank analysis of GPT-2 Medium
+(b) Rank analysis of GPT-2 Large
+(c) Rank analysis of GPT-2 XLarge
+(d) Matrix mass analysis of GPT-2 Medium
+(e) Matrix mass analysis of GPT-2 Large
+(f) Matrix mass analysis of GPT-2 XLarge
+Figure 10: In standard decoder-style LLMs, attention matrices in deeper layers often degenerate into single-column matrices, leading to layers with fully degenerated attention that fail to learn meaningful representations.
+All models were trained on the OpenWebText dataset, and both rank and mass analyses were conducted using the Fineweb_edu validation set, following the same procedure described in Figure 1.
+This further demonstrates the robustness of our analysis, as we reach the same conclusion using different evaluation datasets.
+Figure 11: Rank analysis of LLaMA-3 8B reveals that nearly half of the attention heads exhibit rank collapse.
+We analyze the LLaMA-3 8B model, which contains 32 heads per layer (32 $\times$ 32), using the rank metric defined in Section 2.
+The results are visualized as a heatmap of head index vs.
+layer index.
+Potent (non-collapsed) heads are shown in varying shades of green, where higher intensity indicates higher rank, while rank-collapsed heads (near rank-1) are highlighted in red.
+Approximately 50% of all attention heads exhibit rank collapse, indicating widespread degeneracy.
+L4 H2
+L7 H3
+L20 H2
+L22 H4
+24 layer GPT2 medium model trained from scratch.
+L4 H2
+L7 H3
+L11 H2
+L15 H1
+Our 16-layer GPT2 medium model trained using $\mathsf{Inheritune}$.
+Figure 12: $\mathsf{Inheritune}$ preserves effective attention patterns in smaller models.
+Comparison of attention patterns across layers (L) and heads (H) in two GPT2-medium models: (top) 24-layer model trained from scratch, (bottom) 16-layer model trained with $\mathsf{Inheritune}$.
+Attention maps are averaged over three randomly selected string, with 40 tokens each from the validation.
+Darker colors indicate higher attention scores.
+$\mathsf{Inheritune}$ maintains focused attention even in deeper layers, contrasting with the uniform patterns in the standard model’s deeper layers.
+L8 H16
+L8 H17
+L8 H18
+L8 H19
+L8 H20
+A Potent (non-lazy) layer of a pre-trained GPT2 xLarge 48 layer model.
+L30 H16
+L30 H17
+L30 H18
+L30 H19
+L30 H20
+A lazy layer of a pre-trained GPT2 xLarge 48 layer model.
+Figure 13: Visualization of attention patterns in lazy and non-lazy layers of a vanilla GPT-2 xLarge model with 48 layers.
+The top row displays attention patterns for various heads (H) in layer (L) 8, while the bottom row shows patterns for layer (L) 30.
+Appendix B Baselines
+We compare $\mathsf{Inheritune}$ against several baselines.
+While some baseline methods are illustrated using GPT-2 Large or medium (for the knowledge distillation baseline) as an example, the same methodology is consistently applied across all model variants.
+Baselines trained from scratch (rand init.) :
+We compare our $\mathsf{Inheritune}$-derived model against much larger GPT-2 reference models trained from scratch for the same number of steps and similar-sized models trained from scratch for both the same and double the number of training steps.
+Baselines trained with various model initialization and efficient training techniques.
+Here we compare our model derived using $\mathsf{Inheritune}$, to similar sized models trained with various zeroshot model initialization and effcient training techniques such as stacking, hybrid stacking, and half-width initialization.
+We explain these baseline training recipes using GPT-2 large and its variants as an example and apply the same process for other models.
+Stacking Gong et al.
+(2019); J.
+Reddi et al.
+(2023) is a model initialization and efficient (stage-wise) training recipe.
+We train a 9-layer GPT-2 large variant from scratch for 100K steps, then expanded the model to 18 layers by copying the weights from layers 0-8 to layers 9-17.
+Finally we re-trained this new 18-layer GPT-2 large variant, using stacking initialization for an additional 100K steps.
+Hybrid stacking: Hybrid stacking is stacking but utilizes a large pre-trained reference model for initialization instead of using its own pre-trained weights.
+We took the weights of layers 0-8 from the reference 36-layer GPT-2 large model and expanded it to a 18-layer model by copying the weights to layers 0-17.
+We then trained this new 18-layer GPT-2 variant for 100K steps.
+Half width: We initialized the baseline GPT-2 large variant across the width dimension and preserved the entire depth.
+We copied the weights of the first half the attention heads (0-9) and MLPs of the GPT-2 large reference model into baseline GPT-2 variant with half the width but all layers.
+Baselines trained with Knowledge Distillation
+As a baseline, we first apply logit-based knowledge distillation Hinton et al.
+(2015) to train a 16-layer GPT-2 medium variant (student) initialized randomly.
+For the second baseline, we use a DistillBERT-style approach Sanh et al.
+(2019), where the student model 0-11 layers are initialized with every alternate block of its teacher, and the remaining 4 blocks are initialized using layers 18, 19, 20, and 21 of the teacher.
+Both baselines are trained for 50K steps, using a vanilla 24-layer GPT-2 medium model as the teacher (our reference model).
+Appendix C Supplementary Experiments
+We provide additional training data and plots for our main results discussed in Section 4.1 as shown in Table 3 and Figure 14.
 In Figure 5 (also refer Table 3 we compare our GPT-2 variants with baseline models trained form scratch.
@@
 Recall we have discussed attention degeneration in Section 2 and attention patterns are visualized in Figure 12.
-Following up on our previous discussions in Figure 9 we demonstrate that models trained with $\mathsf{Inheritune}$ has lesser lazy layers (refer Figure 9) compared to it’s larger counterpart trained form scratch.
-We performed rank analysis utilizing 24-layer GPT-2 medium and our 16-layer GPT-2 model trained using $\mathsf{Inheritune}$.
+Following up on our previous discussions in Figure 17 and Figure 18 we demonstrate that models trained with $\mathsf{Inheritune}$ has lesser lazy layers compared to it’s larger counterpart trained form scratch.
+We performed rank analysis for Figure 17 utilizing vanilla 24-layer GPT-2 medium and our 16-layer GPT-2 variant trained using $\mathsf{Inheritune}$.
+Additionally, we performed rank analysis for Figure 18 with a vanilla 48-layer GPT2 xLarge and a 24-layer GPT2 xLarge variant trained using $\mathsf{Inheritune}$.
 Recall we have previously discussed that attention degeneration is connected with vanishing gradients of keys and queries Noci et al.
@@
 Since we are training smaller models intuitively $\lVert W_{Q}\rVert$ and $\lVert W_{K}\rVert$ should be smaller compared to their larger counterparts and hence the norm of gradients in the case of smaller models derived using $\mathsf{Inheritune}$ is higher leading to better training.
-(a) GPT2-Medium
-(b) GPT2-Large
-(c) GPT2-xLarge
-Figure 5: Models derived using Inheritune converge faster and match the final validation loss of the full-sized model trained from scratch, despite being smaller.
-Training GPT-2 xlarge, GPT-2 large and GPT-2 medium vanilla models from scratch and our variants with OpenWebText-9B for 100K steps.
+Models Layers Initialization Steps Pre-train Val loss ($\downarrow$)
+GPT-2 Medium 24 rand init 100K 2.81
+16 rand init 100K 2.86
+12 Ours 100K 2.87
+14 Ours 100K 2.84
+Final Model $\longrightarrow$ 16 Ours 100K 2.81
+GPT-2 Large 36 rand init 100K 2.85
+18 rand init 100K 2.97
+18 Ours 100K 2.80
+GPT-2 xLarge 48 rand init 100K 2.65
+24 rand init 100K 2.69
+24 Ours 100K 2.64
+Table 3: Inheritune achieves superior performance with reduced model size.
+Comparison of $\mathsf{Inheritune}$-trained models (24-layer GPT-2 xLarge, 18-layer GPT-2 Large, and 16-layer GPT-2 Medium) against full-sized counterparts.
+All models are trained for 100K steps, and reported using validation loss on the OpenWebText validation set.
+Note: GPT-2 Large and xLarge use one round of $\mathsf{Inheritune}$; GPT-2 Medium uses three rounds.
 (a) GPT2-Medium
@@
 Figure 17: Rank collapse in deeper layers and its mitigation through $\mathsf{Inheritune}$.
-(a) Analysis of a 24-layer GPT2 medium model reveals rank-1 attention matrices in later layers, indicating rank collapse.
-(b) Our 16-layer GPT2 medium variant, trained with $\mathsf{Inheritune}$, demonstrates improved rank across all layers, highlighting the effectiveness of our approach in preserving expressive power.
-Appendix B Developing a 1.5B Small Base LM in a Low Data Regime with Inheritune
+The maximum (max) rank across all attention heads for each layer is plotted, following the methodology in Fig.
+1 (a) Analysis of a 24-layer GPT2 medium model reveals rank-1 attention matrices in later layers (those beyond the halfway point), indicating rank collapse.
+Specifically, 3 out of the last 12 later layers exhibit rank-1 attention matrices (mean rank accross all the 100 runs).
+(b) Our 16-layer GPT2 medium variant, trained with $\mathsf{Inheritune}$, demonstrates improved rank across all layers, highlighting the effectiveness of our approach.
+Notably, none of the later layers in our 16-layer variant exhibit rank-1 attention matrices.
+(a) Vanilla GPT2-xLarge 48 layers
+(b) Our GPT2-xLarge 24 layer variant trained with $\mathsf{Inheritune}$
+Figure 18: Rank collapse worsens for larger LLMs, $\mathsf{Inheritune}$ helps to mitigate rank collapse.
+The maximum (max) rank across all attention heads for each layer is plotted, following the methodology in Fig.
+1 (a) Analysis of a 48-layer GPT2 xLarge model reveals rank-1 attention matrices in later layers (those beyond the halfway point), indicating rank collapse.
+Specifically, 22 out of the last 24 later layers exhibit rank-1 attention matrices (mean rank across all the 100 runs).
+(b) Our 24-layer GPT2 xLarge variant, trained with $\mathsf{Inheritune}$, demonstrates improved rank across all layers, highlighting the effectiveness of our approach.
+Notably, 2 out of 12 of the later layers in our 24-layer variant exhibit rank-1 attention matrices.
+Appendix D Extended Discussion
+D.1 Discussion about Attention Sink
+The term "attention sink" Xiao et al.
+(2024) refers to the phenomenon where the first token in a sequence receives disproportionately high attention scores compared to other tokens in the attention maps.
+While there is some connection with $\mathsf{Inheritune}$, as we have also observed that many attention matrices are not only rank-1 but also single-column (with all attention scores concentrated on the first token), this connection has not been explicitly established in Xiao et al.
+(2024) with respect to rank-1 behavior or poor training of later layers.
+In contrast, as illustrated in Figure 1, we compute the maximum rank of all attention matrices within a layer.
+For instance, consider a layer where only 2 out of 5 attention heads exhibit attention sink behavior.
+This does not make the layer lazy, as attention is computed as a concatenation of activations across all heads.
+A lazy layer, however, has all 5 out of 5 attention heads fully degenerated, with their attention matrices being rank-1.
+We have provided evidence that such lazy layers are indicative of poorly trained layers.
+Layers Initialization Avg max ranks Val Loss ($\downarrow$)
+4 rand n/a 3.25
+4 1-4 layers from vanilla GPT2 8.40 3.22
+4 5-8 layers from vanilla GPT2 9.48 3.19
+4 9-12 layers (lazy layers) from GPT2 1.22 3.23
+Table 4: Impact of initialization strategies on GPT2-small variants.
+We analyzed the rank characteristics of a vanilla GPT2-small model (125M, 12 layers) trained on OpenWebText for 100K steps.
+Four-layer GPT2-small variants were initialized using the first 4 layers [1–4], middle 4 layers [5–8], last 4 layers [9–12], or with random initialization, and then trained for 100K steps on OpenWebText.
+Models initialized with the last 4 layers performed similarly to random initialization, while those initialized with layers exhibiting higher average max ranks achieved the best validation loss, regardless of proximity to the embedding layer.
+The training plots and rank analysis are provided in Figure 2.
+Models Recipe Layers ARCE PIQA SciQ Hellaswag Lambada Average
+(acc) (acc) (acc) (acc norm) (acc)
+GPT-2 Medium rand init 24 51.05 61.81 74.8 30.79 20.28 47.74
+rand init 16 49.92 61.92 73.3 29.56 19.54 46.84
+Ours 16 51.26 61.81 73.8 30.55 23 48.08
+GPT-2 Large rand init 32 52.48 64.58 75.3 32.65 22.2 49.44
+rand init 16 50.34 63.11 75 30.86 21.56 48.17
+Ours 16 52.9 63.55 76.1 32.14 24.06 49.75
+Table 5: Models trained with Inheritune outperforms both their larger and same-size counterparts trained from scratch on average zero-shot downstream performance.
+For evaluation we choose accuracy (acc) and normalized accuracy (acc norm) metrics following Open LLM leaderboard Beeching et al.
+(2023).
+All the models are trained with FineWeb_edu.
+Appendix E Developing a 1.5B Small Base LM in a Low Data Regime with Inheritune
 Figure 19: Performance of our 1.5B base LM derived using 1B data with $\mathsf{Inheritune}$ on an average of 9 different datasets (left) and MMLU benchmark (right) that evaluates commonsense, truthfulness, natural language inference and language understanding.
@@
 The models against which n/a is mentioned are trained from scratch.
-B.1 Main Results in Low Data Regime
+E.1 Main Results in Low Data Regime
 Table 6 presents a detailed performance evaluation across various tasks.
@@
 The MMLU (5-shot) as a function of the number of submodels.
-B.2 Additional analysis with larger reference LMs and 50B data
+E.2 Additional analysis with larger reference LMs and 50B data
 We further analyze $\mathsf{Inheritune}$ to see the impact of it’s performance when more tokens are available.
@@
 We derive all the variants of Our-1.5B small base using $\mathsf{Inheritune}$ with OpenLLaMA-3B as reference model.
-The models featured in this table correspond to those discussed in Figures 12 and 13.
+The models featured in this table correspond to those discussed in Figures 21 and 22.
 Ablations with number of epochs.
@@
 The models discussed in Table are saved checkpoints during a single training run and not the final model unless otherwise specified.
-B.3 Implications of Low Data Regime
+E.3 Implications of Low Data Regime
 In this section, we discuss some of the key implications of our work in low data regime.
@@
 Therefore our $\mathsf{Inheritune}$ recipe has the potential to become the naive baseline for any pre-training pipeline aiming to develop a smaller variant of a large base LM.
-Appendix C Implementation Details
-C.1 Training details of GPT-2 models
+Appendix F Implementation Details
+F.1 Training details of GPT-2 models
 For our main experiments, we focused on three sizes of GPT-2 models Radford et al.
@@
 Table 11: Overview of all the GPT2 models used in this study and their architectural configurations. The model configurations of stacking and hybrid stacking are same as our variants.
-C.2 Training details of 1.5B OpenLLaMA model
+F.2 Training details of 1.5B OpenLLaMA model
 Small base LMs trained with 1B data

### v4

@@
 Abstract
-Large Language Models (LLMs) leverage the transformer architecture and its self-attention mechanism to achieve remarkable performance.
-However, we identify a structural inefficiency in standard pre-trained decoder-style LLMs: attention matrices in many deeper layers frequently degenerate, often collapsing to near rank-one, single-column structures.
-These layers, which we term lazy layers, are redundant and inefficient.
-To address this, we introduce $\mathsf{Inheritune}$, a simple yet effective training recipe for developing smaller, more attentive, and high-performing language models.
-$\mathsf{Inheritune}$ initializes a compact model by inheriting the potent early layers from a larger pre-trained model, then progressively retrains and expands it.
-Our demonstration of $\mathsf{Inheritune}$’s efficacy on various models and datasets reveals that models trained with this approach, despite having significantly fewer layers, can achieve performance comparable to or even exceeding their larger counterparts.
-This approach enables model compression by design, producing language models that are both compact and high-performing.
+Large Language Models (LLMs) are known for their performance, but we uncover a significant structural inefficiency: a phenomenon we term attention collapse.
+In many pre-trained decoder-style LLMs, the attention matrices in deeper layers degenerate, collapsing to near rank-one structures.
+These underutilized layers, which we call lazy layers, are redundant and impair model efficiency.
+To address this, we introduce $\mathsf{Inheritune}$, a simple yet powerful training recipe designed to build smaller, stronger language models.
+Inheritune initializes a compact model by inheriting the potent early layers from a larger pre-trained model and then progressively trains and expands it.
+Our experiments on various models, including the GPT-2 family, demonstrate that models trained with $\mathsf{Inheritune}$ can match or even surpass the performance of their larger counterparts, despite having significantly fewer layers.
+This work presents a novel path toward model compression by design, enabling the creation of compact, yet highly performant language models.
 Code is available at https://github.com/sanyalsunny111/LLM-Inheritune.
 1 Introduction
-Large Language Models (LLMs) are built with decoder-style transformer blocks (Vaswani et al., 2017).
-These models are typically designed to be large, with a significant portion of their parameters dedicated to their depth, with multiple transformer blocks stacked with eachother building model capacity.
+Large Language Models (LLMs) are composed of stacks of decoder-style transformer blocks (Vaswani et al., 2017).
+As the model grows in size, the model capacity and performance typically improve (Kaplan et al., 2020; Hoffmann et al., 2022).
+A substantial fraction of the total parameters is devoted towards adding more transformer blocks to increase the depth.
 Each block or layer in the stack refines the representations learned by the previous blocks, allowing the model to develop a nuanced understanding of the input data.
-As these models scale in depth and size, their performance tends to improve Kaplan et al.
-(2020); Hoffmann et al.
-(2022), benefiting from increased model capacity.
-The causal self-attention (hereafter referred to as attention) mechanism is arguably the most crucial component of a transformer block.
-It allows models to combine tokens as a weighted linear sum of their attention scores, effectively capturing long-range dependencies and contextual relationships within text data.
-However, as models grow in depth, they often encounter a phenomenon known as attention degeneration caused by collapse in the attention rank ((Noci et al., 2022; Dong et al., 2021; He et al., 2023)).
-Notably, this phenomenon has not been studied in the context of standard LLMs.
-A formal discussion on attention degeneration is provided in Section 2.
-In this paper, we empirically analyze 24-layer GPT-2 medium and 36-layer GPT-2 large models (decoder-style LLMs) Radford et al.
-(2019) for attention degeneration and observe that many deeper layers in both models exhibit rank-1 attention matrices.
-Further investigation reveals that most of these rank-1 matrices are also single-column, i.e.
-their mass is concentrated to a single column.
-Our attention matrix analysis is presented in Figure 1.
-We term these deeper layers, where all attention matrices of a given layer are degenerated, as lazy layers.
-Motivated by this new finding we aim to develop performant small base language models (LMs) utilizing weights from inefficient larger base LMs.
+A transformer block primarily consists of a self-attention module and a feed-forward network (FFN, also referred to as an MLP).
+Among these, the causal self-attention mechanism (hereafter referred to as attention) is arguably the most critical component.
+It enables the model to combine token embeddings as a weighted linear sum of attention scores, effectively capturing long-range dependencies and contextual relationships within text.
+However, as models become deeper, they often exhibit a phenomenon known as attention degeneration, characterized by a collapse in the rank of the attention matrices.
+While prior studies have analyzed rank collapse in simplified transformer settings (Dong et al., 2021; Noci et al., 2022; He et al., 2023), this phenomenon has not, to our knowledge, been systematically explored in standard decoder-only LLMs.
+A formal discussion of attention degeneration is provided in Section 2.
+In this paper, we conduct a detailed empirical analysis of attention degeneration in the GPT-2 family of LLMs (Radford et al., 2019), including GPT-2 Medium (355M), GPT-2 Large (770M), and GPT-2 XLarge (1.5B).
+Our analysis reveals that many deeper layers in these models exhibit predominantly rank-1 attention matrices across most attention heads within a layer.
+This suggests that the attention mechanism loses its discriminative ability among tokens and instead performs near-uniform averaging across the sequence.
+We refer to layers in which all attention matrices degenerate to near rank-1 as lazy layers (a more formal definition is provided in Definition G.1).
+In the supplementary material, we further extend this analysis to billion-sized LLMs, including LLaMA-3 8B (refer Figure 11), Falcon-7B (refer Figure 21), OLMo-1B, Cerebras GPT 2.7B and LLaMA-3 3B (refer Figure 22), to highlight that attention collapse persists even in several modern architectures.
+Motivated by the new finding through our novel analysis we aim to develop performant small base language models (LMs) utilizing weights from inefficient larger base LMs without losing pre-train performance (measured by train/validation loss).
 A base LM is a decoder-style model trained solely for next-token prediction without additional enhancements like instruction tuning or reinforcement learning with human feedback (RLHF).
@@
 We refer to this simple yet effective training approach as $\mathsf{Inheritune}$.
-(a) GPT-2 Medium (355M) L=24, H=1024
-(b) GPT-2 Large (770M) L=36, H=1280
-(c) GPT-2 XLarge (1.5B) L=48, H=1600
+(a) GPT-2 Medium
+(b) GPT-2 Large
+(c) GPT-2 XLarge
 Figure 1: In decoder-style LLMs, attention matrices in deeper layers often degenerate to near rank-1, limiting their ability to learn meaningful representations.
@@
   1. 1.
-    Analysis of Attention Degeneration Leading to Lazy Layers.
-We empirically investigate attention degeneration in standard LLM settings.
-Our analysis shows that rank-collapsed attention matrices often exhibit single-column structures, revealing a significant structural inefficiency in the attention mechanism of standard LLMs in deeper layers (refer Figure 1).
+    Novel analysis of attention degeneration in standard decoder LLMs.
+We empirically investigate attention degeneration in standard decoder style LLMs.
+Our analysis shows that rank-collapse in attention matrices, revealing a significant structural inefficiency in the attention mechanism of standard LLMs in deeper layers (refer Figure 1).
+This degeneration gives rise to what we refer to as lazy layers.
   2. 2.
-    Introduction of Inheritune.
-We propose $\mathsf{Inheritune}$ as an approach to effectively train high-performing, smaller models.
-This method involves inheriting a few early blocks from a larger pre-trained model and progressively growing and training the smaller model.
-The initialization is entirely zero-shot.
-We validate the effectiveness of $\mathsf{Inheritune}$ through comprehensive experiments using GPT-2 xlarge (1.5B), GPT-2 large (770M), and GPT-2 medium (355M) models, trained on the OpenWebText and the FineWeb_edu dataset.
+    Introduction of our training recipe Inheritune.
+Building on our analysis we observe that deep LMs often fail to fully utilize their effective depth.
+To address this inefficiency, we propose $\mathsf{Inheritune}$— a simple yet effective training recipe for developing smaller LMs without losing pre-training performance.
+This method involves inheriting a few early blocks from a much larger reference model and progressively growing and training the smaller model.
+We validate the effectiveness of $\mathsf{Inheritune}$ through extensive experiments on GPT-2 XLarge (1.5B), GPT-2 Large (770M), and GPT-2 Medium (355M) models, trained primarily on the OpenWebText dataset and additionally on FineWeb data.
   3. 3.
     Evaluation against multiple baselines.
-Models derived using $\mathsf{Inheritune}$ consistently outperform various baselines, including much larger models trained from scratch (refer Figure 5 and Figure 6), model initialization and efficient training baselines (refer Table 1), and models trained using two knowledge distillation techniques (refer Figure 7).
+Models trained using $\mathsf{Inheritune}$ consistently outperform a wide range of baselines, including much larger models trained from scratch (see Figure 5 and Figure 7), as well as same sized models trained from scratch for twice as many steps (extended training; see Figure 6).
+We further compare against warm-started baselines (models initialized with pre-trained weights rather than random initialization  (Ash and Adams, 2020)) as shown in Table 2.
 2 Attention Collapse and the Emergence of Lazy Layers in LLMs
@@
 The output of each layer $l$ is denoted as $X^{(l)}\in\mathbb{R}^{T\times e}$.
-Each transformer block primarily consists of two sub-layers: a self-attention block and a position-wise feed-forward network (FFN).
+Each transformer block primarily consists of two modules: a self-attention block and a feed-forward network (FFN).
 The self-attention mechanism enables the model to weight the relevance of different tokens in the sequence relative to each other.
-Specifically, for a single attention head, the attention computation is defined as $\text{Attention}(Q,K,V)=\underbrace{\text{softmax}\left(\frac{QK^{\top}}{\sqrt{d_{k}}}\right)}_{\textbf{Attention matrix: $A(X)$}}V$
+Specifically, for a single attention head, the attention computation is defined as equation 1.
+$\text{Attention}(Q,K,V)=\underbrace{\text{softmax}\left(\frac{QK^{\top}}{\sqrt{d_{k}}}\right)}_{\textbf{Attention matrix: $A(X)$}}V$ (1)
 where the queries $Q=XW_{Q}$,  keys $K=XW_{K}$,  and values $V=XW_{V}$ are linear transformations of the input $X$.
@@
 In this work, we set $\tau=0.90$.
-A lower $k^{*}$ indicates stronger rank collapse.
 In Figure 1, we present the layer-wise analysis of rank of GPT-2 models.
@@
 As demonstrated in Figure 1, $\text{MaxRank}^{(l)}$ reveals that many deeper layers exhibit attention matrices that are predominantly near rank-1.
-We highlight that this rank collapse occurs in GPT-2 medium, Large, and xLarge models—which are standard architectures incorporating skip connections and FFNs—thereby extending the findings of Dong et al.
+We highlight that this rank collapse occurs in GPT-2 Medium, Large, and XLarge models, which are widely used modern LLMs thereby extending the limited findings of Dong et al.
 (2021) and Noci et al.
-(2022) to these widely-used LLMs.
-The rank analysis is quite robust to the dataset in Figure 10, we performed the same analysis using the same models on held out data from Fineweb_edu dataset (Penedo et al., 2024).
+(2022).
+We further visualize attention matrices from representative potent and lazy layers of GPT-2 XLarge in Figure 12.
 Overall, the degeneration of attention matrices in deeper layers provides quantitative evidence for the existence of lazy layers.
 Specifically, we observe that some deeper layers exhibit a near-complete rank collapse of attention matrices across all heads, suggesting potentially reduced representational capacity and less effective token mixing in these layers.
+We provide an extended discussion of attention collapse in Appendix B.
+We analyze five modern LLMs for attention collapse (also refer Appendix G).
+We further assess the robustness of our findings by conducting collapse analyses on different datasets and by using an alternative metric, namely the mass of attention matrices.
+Finally, we perform an ablation over a range of values of $\tau$ (refer Appendix 4.3) to study the sensitivity of the rank-based analysis with respect to $\tau$.
 2.1 The Functional Ineffectiveness of Lazy Layers
@@
 Our experiments suggest the latter.
-(a) Rank profile (GPT-2 Small 12L).
-(b) Performance of 4L variants.
-Figure 2: Higher-rank layers transfer better.
+(a) Rank analysis of GPT-2 Small (12 layer).
+(b) Performance of 4 layer GPT-2 small variants.
+Figure 2: Higher-rank (potent) layers transfer better.
 (Left, 2(a)) Layer-wise $\mathrm{MaxRank}^{(l)}$ of a pre-trained 12L GPT-2 Small.
-(Right, 2(b)) Validation loss of 4L variants initialized with potent blocks (AvgRank $\approx 8.4-9.5$) vs.
-a lazy block (AvgRank $\approx 1.2$) or random weights, after 100K steps.
-Lazy block initialization mirrors random.
-In the first set of experiments, we trained a vanilla GPT2-small (125M) model with 12 layers for 100K steps on the OpenWebText dataset.
+(Right, 2(b)) Validation loss of 4 layer variants initialized with potent layers (AvgRank $\approx 8.4-9.5$) vs.
+lazy layers (AvgRank $\approx 1.2$) and random weights, after training for 100K steps.
+Models initialized with lazy layers mirrors the model with random initialization.
+Training curves are smoothed for visual clarity.
+In the first set of experiments, we trained a vanilla GPT-2 small (125M) model with 12 layers for 100K steps on the OpenWebText dataset.
 We then performed the rank analysis described earlier, with results presented in Figure 2.
 Specifically, we aggregated the approximate ranks over groups of contiguous layers using $\text{AvgRank}=\frac{1}{L}\sum_{l=1}^{L}\text{MaxRank}^{(l)}$, where $L$ is the number of layers in each group.
-Subsequently, we trained three GPT2-small variants [^1] A variant shares the same configurations as the parent model but has fewer layers.
-for 100K steps, each initialized with a different contiguous block of four layers from the trained vanilla GPT2-small model: (a) layers 1–4, with $\text{AvgRank}=8.40$; (b) layers 5–8, with $\text{AvgRank}=9.48$; and (c) layers 9–12, with $\text{AvgRank}=1.22$.
+Subsequently, we trained three GPT-2 small variants [^1] A variant shares the same configurations as the reference model but has fewer layers.
+for 100K steps, each initialized with a different contiguous block of four layers from the trained vanilla GPT-2 small model: (a) layers 1–4, with $\text{AvgRank}=8.40$; (b) layers 5–8, with $\text{AvgRank}=9.48$; and (c) layers 9–12, with $\text{AvgRank}=1.22$.
 The last model is initialized with lazy layers.
-For comparison, we also trained another GPT2-small variant with random initialization for 100K steps.
+For comparison, we also trained another GPT-2 small variant with random initialization for 100K steps.
 All models were trained on the OpenWebText dataset.
-As shown in Figure 2, the model initialized with layers from the vanilla GPT2-small model having higher AvgRank demonstrated the best performance (i.e., lowest final validation loss).
+As shown in Figure 2, the model initialized with layers from the vanilla GPT-2 small model having higher AvgRank demonstrated the best performance (i.e., lowest final validation loss).
 Additionally, we observed that the model initialized with lazy layers performed very similarly to the model with random initialization suggesting that lazy layers contain minimal transferable knowledge.
-The key results are also summarized in Table 4.
-(a) GPT2-medium (12-layer) variants.
-(b) GPT2-large (16-layer) variants.
-Figure 3: When initializing 12-layer and 16-layer variants of GPT2-medium and GPT2-large with deeper (lazy) layers showing degenerated attention, performance is comparable to random initialization.
-In contrast, early-layer initialization leads to significantly better convergence and generalization.
-For the second set of experiments we utilized larger models namely GPT-2 medium and GPT-2 large both similarly trained for 100K steps using OpenWebText dataset.
+The results are also summarized in Table 4.
+(a) GPT-2 Medium (12 layer) variants.
+(b) GPT-2 Large (18-layer) variants.
+Figure 3: Initializing 12-layer and 18-layer variants of GPT-2 Medium and GPT-2 Large with deeper lazy layers exhibiting degenerated attention results in performance comparable to random initialization.
+In contrast, initializing with early (high-rank) potent layers leads to substantially better convergence and generalization.
+Training curves are smoothed for visual clarity.
+For the second set of experiments we utilized larger models namely GPT-2 Medium and GPT-2 Large both similarly trained for 100K steps using OpenWebText.
 Here we initialized a 12-layer GPT-2 Medium variant and an 18-layer variant of GPT-2 Large using lazy layers extracted from pre-trained 24-layer GPT-2 Medium and 36-layer GPT-2 Large models.
@@
 As shown in Figure 3, models initialized with lazy layers demonstrate poor transferability, performing similarly to models with random initialization.
-This provide additional evidence that lazy layers with fully degenerated attention, fails to learn meaningful representations.
+This provides additional evidence that lazy layers with fully degenerated attention, fails to learn meaningful representations.
+Theoretical Analysis.
+Finally, we analyze the implications of attention rank collapse on model training from a theoretical perspective (refer Section H).
+Our key insight is that rank-collapsed attention head(s) impede learning by inducing vanishing gradients, effectively suppressing updates to the associated $W_{Q}$ and $W_{K}$.
 3 Inheritune: Our Proposed Training Recipe
-This section provides a detailed description of our method, key implementation considerations, and how it addresses the inefficiencies present in current architectures.
-  1. 1.
-    Inherit: Initialize $\mathcal{M}_{\text{tgt}}$ with the first $l$ out of $L$ layers of $\mathcal{M}_{\text{ref}}$, including prediction head, and token embeddings.
-  2. 2.
-    Train: Train $\mathcal{M}_{\text{tgt}}$ for $\mathsf{T}$ steps on $\mathcal{D}_{\text{train}}$ and evaluate on $\mathcal{D}_{\text{val}}$.
-  3. 3.
-    Grow: If needed, increase $\mathcal{M}_{\text{tgt}}$’s size by adding next few contiguous layers and repeat steps 1-2 until desired performance is achieved.
-As previously established, we have identified the problem of attention degeneration and its connection to lazy layers, highlighting specific inefficiencies in pre-trained LLMs.
-In this work, we transform this challenge into an opportunity to create smaller base language models, which we refer to as target models $\mathcal{M}_{\text{tgt}}$, that achieve comparable performance with similar or lower validation loss compared to their larger, less efficient counterparts, which we term reference models $\mathcal{M}_{\text{ref}}$.
 Algorithm 1 Inheritune: Training Recipe for Small Language Models
@@
 5:Train $\mathcal{M}_{\text{tgt}}$ on $\mathcal{D}_{\text{train}}$ for $\mathsf{T}$ steps
-6:while $\mathcal{M}_{\text{tgt}}$ performance < $\mathcal{M}_{\text{ref}}$ performance on $\mathcal{D}_{\text{val}}$ do
+6:while $\mathcal{M}_{\text{tgt}}$ performance << $\mathcal{M}_{\text{ref}}$ performance on $\mathcal{D}_{\text{val}}$ do
 7:  Grow $\mathcal{M}_{\text{tgt}}$ by inheriting additional layers
@@
 10:return Optimized model $\mathcal{M}_{\text{tgt}}$
+This section provides a detailed description of our method, key implementation considerations, and how it addresses the inefficiencies present in current architectures.
+As previously established, we have identified the problem of attention degeneration and its connection to lazy layers, highlighting specific inefficiencies in pre-trained LLMs.
+In this work, we transform this challenge into an opportunity to create smaller base language models, which we refer to as target models $\mathcal{M}_{\text{tgt}}$, that achieve comparable performance with similar or lower validation loss compared to their larger, less efficient counterparts, which we term reference models $\mathcal{M}_{\text{ref}}$.
 Our proposed solution builds on two key insights: (1) the early layers of deep LLMs contain a higher concentration of potent layers with high AvgRank values, making them suitable for effective model initialization, and (2) lazy layers can be identified, removed, or utilized in smaller numbers, then subsequently re-trained to improve overall model capacity.
-Figure 4: Overview of the $\mathsf{Inheritune}$ training recipe using a 24-Layer GPT-2 medium model example.
+Figure 4: Overview of the Inheritune training recipe using a 24-Layer GPT-2 Medium model example.
 A smaller target model is initialized using early layers from a larger, pre-trained reference model.
-The target model goes multiple rounds of training while inheriting more early layers until it matches the reference model.
+The target model goes multiple rounds of training while inheriting contiguous layers until it matches/outperforms the reference model.
 The intensity of the red color in layers correlates with $\mathrm{MaxRank}^{(l)}$.
@@
 The $\mathsf{Inheritune}$ process consists of three main steps, which we present below and formalize in Algorithm 1:
+  1. 1.
+    Inherit: Initialize $\mathcal{M}_{\text{tgt}}$ with the first $l$ out of $L$ layers of $\mathcal{M}_{\text{ref}}$, including prediction head, and token embeddings.
+  2. 2.
+    Train: Train $\mathcal{M}_{\text{tgt}}$ for $\mathsf{T}$ steps on $\mathcal{D}_{\text{train}}$ and evaluate on $\mathcal{D}_{\text{val}}$.
+  3. 3.
+    Grow: If needed, increase $\mathcal{M}_{\text{tgt}}$’s size by adding the next few contiguous layers and repeat steps 1-2 until desired performance is achieved.
 With our method now formally described, we turn to empirical validation.
@@
 4 Experiments
-We evaluate $\mathsf{Inheritune}$ through a comprehensive set of experiments using several GPT-2 models: 48-layer GPT-2 xlarge (1.5B), 36-layer GPT-2 large (770M), 32-layer GPT-2 large (668M) and 24-layer GPT-2 medium (355M) models Radford et al.
-(2019).
-Table 11 provides detailed specifications of all model configurations used.
-We use two training datasets in our experiments: OpenWebText Gokaslan & Cohen (2019) and Fineweb_edu Penedo et al.
-(2024).
-Our experimental setup closely follows prior works Liu et al.
-(2023); Sanyal et al.
-(2024).
-For models trained on OpenWebText, we report validation loss (log perplexity) on a held-out set.
-For those trained on Fineweb-Edu, we report training loss (also log perplexity).
-In addition, we conduct zero-shot downstream evaluations using the lm-evaluation-harness Gao et al.
-(2024) across five standard tasks: ARC-easy (ARCE; Clark et al.
-(2018)), LAMBADA Paperno et al.
-(2016), SciQ Welbl et al.
-(2017), HellaSwag Zellers et al.
-(2019), and PIQA Bisk et al.
-(2020).
-We also perform a thorough ablation study of our initialization strategy, focusing on 16-layer GPT-2 medium variant.
-We provide experimental details our $\mathsf{Inheritune}$ training recipe using a GPT-2 medium model as an example; similar procedure was applied to train other models.
+We evaluate $\mathsf{Inheritune}$ through a comprehensive set of experiments using several GPT-2 models: a 48-layer GPT-2 XLarge (1.5B), a 36-layer GPT-2 Large (770M), a 32-layer GPT-2 Large (668M), and a 24-layer GPT-2 Medium (355M) (Radford et al., 2019).
+Table 6 provides detailed specifications of all model configurations used in our experiments.
+We use two training datasets: OpenWebText Gokaslan and Cohen (2019) with 10B tokens and FineWeb (education subset) (Penedo et al., 2024) also with 10B tokens.
+Our experimental setup closely follows prior work (Liu et al., 2023; Sanyal et al., 2024).
+For models trained on OpenWebText, we report validation loss (log perplexity), while for models trained on FineWeb, we report training loss (also log perplexity).
+Additionally, for models trained on FineWeb, we conduct zero-shot downstream evaluations using the lm-evaluation-harness (Gao et al., 2024) across five standard benchmarks: ARC-Easy (ARCE; Clark et al., 2018), LAMBADA (Paperno et al., 2016), SciQ (Welbl et al., 2017), HellaSwag (Zellers et al., 2019), and PIQA (Bisk et al., 2020).
+Finally, we perform a detailed ablation study by initializing each submodule within a transformer block in isolation and training it for 100K steps to identify which component contributes most to performance.
+All training curves are smoothed for visual clarity.
+(a) GPT-2 Medium
+(b) GPT-2 Large
+(c) GPT-2 XLarge
+Figure 5: Models derived using Inheritune converge faster and match the final validation loss of the full-sized model, despite having much fewer layers.
+Comparison of $\mathsf{Inheritune}$ trained models (24-layer GPT-2 XLarge variant, 18-layer GPT-2 Large variant, 16-layer GPT-2 Medium variant) against their full-sized counterparts and same sized variants trained from scratch.
+All models are trained for 100K steps using OpenWebText data.
+(a) GPT-2 Medium
+(b) GPT-2 Large
+(c) GPT-2 XLarge
+Figure 6: Models trained with Inheritune match the validation loss of same sized models trained from scratch for twice as many steps.
+We compare $\mathsf{Inheritune}$-trained models (24-layer GPT-2 XLarge, 18-layer GPT-2 Large, and 16-layer GPT-2 Medium) against their same-sized counterparts trained from scratch for twice the number of steps.
+$\mathsf{Inheritune}$ models are trained for 100K steps, while the baseline models (trained from scratch) are trained for 200K steps, on OpenWebText data.
+We provide experimental details of our proposed training recipe $\mathsf{Inheritune}$ using a GPT-2 Medium model as an example; similar procedure was applied to train other models.
 A visualization of the training recipe is presented in Figure 4.
@@
   1. 1.
-    Reference Model: We train the full 24-layer GPT-2 medium model on $\mathcal{D}_{\text{train}}$ for 100K steps and evaluate its validation loss ( log-perplexity) on $\mathcal{D}_{\text{val}}$.
+    Reference Model: We train a vanilla 24-layer GPT-2 Medium model (reference model) on $\mathcal{D}_{\text{train}}$ for 100K steps and evaluate its validation loss ( log perplexity) on $\mathcal{D}_{\text{val}}$.
 This establishes our benchmark validation loss.
@@
   4. 4.
-    Iterative Refinement: If the smaller model’s performance is inferior to the reference model, then we incrementally increase its size by two layers and repeat steps 2-3 until we achieve parity with the reference model’s validation loss.
+    Iterative Refinement: If the smaller model’s performance is inferior to the reference model, then we incrementally increase its size by adding additional layers and repeat steps 2-3 until we achieve parity with the reference model’s validation loss.
 We choose $l=L/2$ as the starting point and increase the model size by two layers in each round across all our experiments, aiming to minimize the number of training rounds.
 In principle, $\mathsf{Inheritune}$ should generalize to other hyperparameter choices as well.
-(a) GPT2-Medium
-(b) GPT2-Large
-(c) GPT2-xLarge
-Figure 5: Models derived using Inheritune converge faster and match the final validation loss of the full-sized model, despite having much fewer layers.
-Comparison of $\mathsf{Inheritune}$-trained models (24-layer GPT-2 xLarge variant, 18-layer GPT-2 Large variant, 16-layer GPT-2 Medium variant) against their full-sized counterparts and same sized variants trained from scratch.
-All models are trained for 100K steps using OpenWebText data with data repetition.
-(a) GPT2-Medium
-(b) GPT2-Large
-Figure 6: Models derived using $\mathsf{Inheritune}$ converge faster and match the final validation loss and average downstream performance (Perf.) across 5 datasets of the full-sized model, while outperforming same-sized models.
-(Plots) Validation loss curves for our 16-layer variants versus full-sized models.
-(Table) Zero-shot downstream performance comparison.
-All models are trained for 100K steps using FineWeb_edu data without data repetition.
-Models Recipe Layers Perf. Medium rand init 24 47.74 rand init 16 46.84 Ours 16 48.08 Large rand init 32 49.44 rand init 16 48.17 Ours 16 49.75
-Baselines.
-We compare GPT-2 model variants (variants are models with lesser layers compared to their vanilla configuration) trained using $\mathsf{Inheritune}$ against two types of baselines: (1) larger reference models with more layers, and (2) similar-sized models with the same number of layers trained from scratch (randomly initialized).
-Additionally, we compare $\mathsf{Inheritune}$ against efficient training baselines that leverage pre-trained model initialization.
-These include stacking Gong et al.
-(2019); J.
-Reddi et al.
-(2023), hybrid stacking (which initializes the model using layers from the reference model), and half-width initialization (where all layers are retained, but the hidden dimension and number of attention heads are halved and initialized from the reference model).
-Finally we briefly compare $\mathsf{Inheritune}$ with knowledge distillation (Hinton et al., 2015).
-Detailed descriptions of these baselines are provided in the supplementary material Section B.
+(a) GPT-2 Large
+(b) GPT-2 Medium
+Figure 7: Models derived using Inheritune without data repetition converge faster and match the final validation loss of the full-sized model despite using lesser layers.
+Additionally, the model trained using $\mathsf{Inheritune}$ demonstrates data efficiency, achieving a lower validation loss in fewer steps compared to its full-sized and half-sized counterparts.
+Models Recipe Layers ARCE PIQA SciQ Hellaswag Lambada Average
+GPT-2 Medium rand init 24 51.05 61.81 74.8 30.79 20.28 47.74
+rand init 16 49.92 61.92 73.3 29.56 19.54 46.84
+Ours 16 51.26 61.81 73.8 30.55 23 48.08
+GPT-2 Large rand init 32 52.48 64.58 75.3 32.65 22.2 49.44
+rand init 16 50.34 63.11 75 30.86 21.56 48.17
+Ours 16 52.9 63.55 76.1 32.14 24.06 49.75
+Table 1: Models trained with Inheritune achieve comparable average zero-shot downstream performance to their larger reference models and surpass same-sized counterparts trained from scratch.
+Downstream evaluations are performed on models pre-trained with the FineWeb dataset (see Figure 7).
+Performance is measured using accuracy (acc) and normalized accuracy (acc-norm) metrics, following the Open LLM Leaderboard protocol Beeching et al.
+(2023).
+We have highlighted the best average scores in bold.
+Baseline-I.
+We compare GPT-2 model variants (i.e., models with fewer layers than their vanilla configurations) trained using $\mathsf{Inheritune}$ against the following baselines:
+  1. 1.
+    Larger reference models with more layers, trained from scratch (random initialization).
+  2. 2.
+    Same-sized models with the same number of layers, trained from scratch.
+  3. 3.
+    Extended training baselines, same-sized models trained from scratch for twice as many steps compared to models trained with $\mathsf{Inheritune}$.
+Baseline-II.
+Additionally, we compare $\mathsf{Inheritune}$ against warm-started baselines (Ash and Adams, 2020).
+Warm starting in neural network training refers to initializing a model’s parameters with weights from a previously trained model, rather than starting from random initialization (a “cold start”).
+This setup enables a fair comparison with our approach, which also leverages prior learned representations.
+The warm-started baselines include stacking (Gong et al., 2019; J.
+Reddi et al., 2023), hybrid stacking (which initializes model layers directly from a larger reference model), and half-width (which retains all layers but reduces both the hidden dimension and number of attention heads by half, using weights from the reference model).
+Finally, we briefly compare $\mathsf{Inheritune}$ with knowledge distillation (Hinton et al., 2015); results are provided in the supplementary material (see Figure 17).
+Detailed descriptions of all baselines are presented in Section D of the supplementary material.
 Models Layers Recipe Steps Pre-train Val loss ($\downarrow$)
@@
 24 Ours 100K 2.64
-Table 1: $\mathsf{Inheritune}$ outperforms baseline zero-shot initialization and efficient training techniques.
-Comparison of pre-training validation loss for GPT-2 xLarge, GPT-2 Large and GPT-2 Medium variants.
-$\mathsf{Inheritune}$-derived models consistently achieve lower loss compared to models initialized with stacking, hybrid stacking, and half-width techniques.
+Table 2: Inheritune outperforms warm-started baselines (Baseline-II).
+Comparison of pre-training validation loss for GPT-2 XLarge, GPT-2 Large, and GPT-2 Medium variants trained with Stacking, Hybrid Stacking, Half-Width, and $\mathsf{Inheritune}$ recipes.
+All baselines are warm-started, i.e., initialized with pre-trained weights rather than random initialization.
+Across model scales, $\mathsf{Inheritune}$ consistently achieves lower validation loss than the majority of warm-started baselines.
+The lowest validation loss (lower is better) is highlighted in bold.
 4.1 Results and Discussions
-Models trained with $\mathsf{Inheritune}$ outperform much larger models trained from scratch.
+Models trained with Inheritune outperform both larger and same sized models trained from scratch.
 We present our main results in Figure 5.
@@
 Our GPT-2 XLarge and GPT-2 Large variants require a single round of $\mathsf{Inheritune}$ training, while the GPT-2 Medium variant undergoes three rounds with 12-, 14-, and 16-layer configurations.
-From a convergence perspective, prior work has linked overparameterization to faster convergence Bengio et al.
-(2005); Vaswani et al.
-(2018).
-Interestingly, we find that smaller models derived using $\mathsf{Inheritune}$ converge just as fast as their larger counterparts.
-A detailed summary of these results is provided in the appendix Table 3.
-We conducted additional training experiments without data repetition, following standard LLM pre-training practices as discussed in Touvron et al.
-(2023); Biderman et al.
-(2023) using the high-quality Fineweb_edu dataset.
-We trained a 32-layer GPT-2 large (668M) and a 24-layer GPT-2 medium (355M) reference model from scratch.
+Furthermore, as shown in Figure 6, models trained with $\mathsf{Inheritune}$ reach the same validation loss as same-sized models trained from scratch in approximately half the number of training steps.
+Moreover, for the GPT-2 Medium and Large variants, $\mathsf{Inheritune}$ achieves a strictly lower loss floor that same-sized models fail to reach even when trained for twice as many steps.
+A tabular summary of these results is provided in Appendix Table 5.
+We conducted additional training experiments, using a high-quality training data namely Fineweb.
+We trained a custom 32-layer GPT-2 Large (668M) and a 24-layer GPT-2 Medium (355M) reference model from scratch.
 Next, we trained two 16-layer variants: one derived from GPT-2 Large and the other from GPT-2 Medium, using their respective reference models following $\mathsf{Inheritune}$.
 For comparison, we also trained 16-layer baseline models from scratch.
-All models were trained for 100K steps.
+All models were trained for 100K steps, and training loss was used to evaluate pre-training performance.
 We observe thematically consistent results: as shown in Figure 7, the 16-layer variants trained with $\mathsf{Inheritune}$ consistently match the performance of their full-sized counterparts and outperform same-sized baselines, both in terms of training loss and zero-shot downstream evaluation.
-Downstream results are averaged over five tasks (see table within Figure 6); the complete breakdown is provided in Table 5 placed in appendix.
+Downstream results are provided in Table 1.
 Model configurations and training hyper-parameters are detailed in the supplementary material (refer Section F).
-Figure 7: A 16-layer GPT-2 medium variant derived using $\mathsf{Inheritune}$ converges faster and generalizes better than a same-sized model trained with logit-based distillation baselines.
-Models trained with Inheritune outperform all zero-shot model initialization baselines.
-In Table 1, we compare GPT-2 xlarge, GPT-2 large, and GPT-2 medium variants trained with $\mathsf{Inheritune}$ against same-sized variants trained with stacking, hybrid, and half-width initialization baselines.
+Models trained with Inheritune outperform all warm started baselines.
+In Table 2, we compare GPT-2 XLarge, GPT-2 Large, and GPT-2 Medium variants trained with $\mathsf{Inheritune}$ against same-sized variants trained with stacking, hybrid stacking, and half-width baselines.
 The half-width baseline performs poorly, revealing the limitations of naive width reduction.
 While stacking and hybrid stacking demonstrate reasonable performance, they still fall short compared to $\mathsf{Inheritune}$.
-Across all cases, $\mathsf{Inheritune}$ consistently outperforms these baselines, highlighting its effectiveness as an initialization strategy.
-For a detailed view of the training dynamics across all methods, refer to the training curves in supplementary Figure 14.
-Distillation vs Inheritune.
-In Figure 7, we compare 16-layer GPT-2 medium variants derived using vanilla knowledge distillation Hinton et al.
-(2015) and DistillBERT-style distillation Sanh et al.
-(2019), which leverages teacher layers for model initialization, vanilla training from scratch and $\mathsf{Inheritune}$.
-Our model demonstrates faster convergence and significantly better final generalization after 50K training steps.
-Additional distillation experiments can be found in the supplementary materials.
-4.2 Ablations
-Layers Initialization Steps Val loss ($\downarrow$)
-16 attn (w/ layernorm) 100K 2.84
-16 mlp (w/ layernorm) 100K 2.85
-16 attn+mlp (w/o layernorm) 100K 2.80
-16 Ours (attn+mlp w/ layernorm) 100K 2.81
+Across all cases, $\mathsf{Inheritune}$ consistently outperforms these baselines, highlighting its effectiveness as an initialization strategy, with a single exception in the GPT-2 XLarge case where it matches one baseline.
+For a detailed view of the training curves across all methods, refer to the training curves in supplementary Figure 18.
+4.2 Inheritune Mitigates Attention Collapse
+We attribute the success of $\mathsf{Inheritune}$, to its ability to mitigate attention collapse, thereby leading to fewer lazy layers after training.
+In Figures 8 we juxtapose the attention rank patterns of the vanilla and $\mathsf{Inheritune}$-trained GPT-2 Medium.
+Notably, none of the GPT-2 Medium variants exhibit lazy layers.
+A similar analysis is conducted with GPT-2 XLarge (refer supplementary Figure 13).
+The corresponding attention patterns for GPT-2 Medium, shown in Figure 9, further corroborate our observation.
+The attention patterns for both a vanilla 24-layer model trained from scratch and a 16-layer model trained using our proposed method, $\mathsf{Inheritune}$.
+Note just for the sake of better visualization we visualized full attention and not causal attention, in practice GPT-2 models compute causal attention.
+We computed these attention matrices using randomly selected strings from the validation set of OpenWebText and took 40 tokens averaged over 3 runs.
+In the 24-layer model trained from scratch (top row of Figure  9), we observe a clear progression in attention patterns.
+The early layers (L4 and L7) exhibit dense structured attention patterns.
+In contrast, the deeper layers (L20 and L22) display more uniform patterns, indicating a loss of focus (attention).
+This uniformity is a hallmark of lazy layers, where the attention mechanism loses its ability to selectively focus on specific relevant tokens.
+In contrast, our 16-layer model trained with $\mathsf{Inheritune}$ (bottom row) demonstrates more focused and effective attention patterns, even in its later layers (L11 and L15).
+This striking difference suggests that our method makes the model more attentive and addresses attention collapse, potentially leading to more efficient models in compact size.
+Figure 8: Rank collapse in deeper layers and its mitigation through Inheritune.
+The maximum (max) rank across all attention heads for each layer is plotted, following the methodology in Figure 1.
+Analysis of a 24-layer GPT-2 medium model reveals rank-1 attention matrices in later layers (those beyond the halfway point), indicating rank collapse.
+Specifically, 3 out of the last 12 later layers exhibit rank-1 attention matrices (mean rank across all the 100 runs).
+Our 16-layer GPT-2 medium variant, trained with $\mathsf{Inheritune}$, demonstrates improved rank across all layers, highlighting the effectiveness of our approach.
+Notably, none of the later layers in our 16-layer variant exhibit rank-1 attention matrices.
+L4 H2
+L7 H3
+L20 H2
+L22 H4
+Vanilla 24 layer GPT-2 Medium
+L4 H2
+L7 H3
+L11 H2
+L15 H1
+GPT-2 Medium 16 layer variant (Ours)
+Figure 9: Inheritune preserves effective attention patterns.
+Comparison of attention patterns across layers (L) and heads (H) in two GPT-2 medium models: (top) vanilla 24-layer model trained from scratch, (bottom) 16-layer variant trained with $\mathsf{Inheritune}$.
+Attention maps are averaged over three randomly selected strings, with 40 tokens each from the validation.
+Darker colors indicate higher attention scores.
+$\mathsf{Inheritune}$ maintains focused attention even in deeper layers, contrasting with the uniform patterns in the vanilla model’s deeper layers.
+4.3 Ablations and Limitations
+We conducted extensive experiments to better understand which sub-module initializations within a transformer block lead to improved generalization (in terms of validation loss) and faster convergence.
+For these ablations, we fixed the model to a 16-layer GPT-2 Medium variant and explored three different sub-module initializations using weights from a 24-layer GPT-2 Medium reference model.
+We initialize the transformer blocks with 1) Attention block (key, query, value, and projection weights) along with the layernorm [^2] In GPT-2 models layernorm blocks are parameterized., 2) Attention and MLP (FFN) weights without the layer-norm weights, and 3) MLP block weights along with the layernorm.
+We note that $\mathsf{Inheritune}$ performs initialization by inheriting the entire layer weight i.e.
+attention, MLP along with the layernorm weights.
+As shown in Table 3, models initialized with both Attention and MLP weights achieve the best performance, irrespective of the LayerNorm initialization.
+A detailed validation loss versus training steps plot is provided in the supplementary Figure 19.
+These results suggest that jointly initializing the Attention and MLP submodules offers a clear advantage over initializing either component alone.
+Interestingly, we also find that initializing only the Attention or only the MLP weights yields comparable improvements in both convergence speed and final validation loss.
+Sensitivity Analysis for attention rank computation and lazy layers.
+Next, we analyze the sensitivity of the approximate rank to the variance threshold $\tau$.
+Following the rank computation methodology described in Section 2, we perform rank analysis on the full GPT-2 Medium and GPT-2 Large models using a randomly sampled subset of the OpenWebText’s validation set, varying $\tau\in\{0.8,0.85,0.9,0.95\}$.
+As shown in Figure 10, the $\mathrm{MaxRank}(l)$ (y-axis) remains highly stable for the lazy layers across all values of $\tau$ for both the models.
+Limitations of our work.
+Our analysis primarily focuses on pre-layernorm (Pre-LN) architectures.
+By default, we initialize training at $l=L/2$, i.e.
+from the midpoint of the model where early layers have already developed strong representations to minimize the total number of training rounds.
+Figure 20 presents an ablation study showing training curves across three rounds performed following our proposed method.
+Notably, by the third round ($R=3$), models trained with $\mathsf{Inheritune}$ match the validation loss of their reference counterpart (Full model).
+Despite its effectiveness, $\mathsf{Inheritune}$ is computationally expensive, as it may require multiple rounds of training during the growth phase.
+Lastly, our current analysis focuses exclusively on attention submodules; extending this framework to feedforward (MLP) layers remains an important direction for future research.
+Layers Initialization Steps Pre-train Val loss ($\downarrow$)
+16 Attention 100K 2.84
+16 MLP 100K 2.85
+16 Attention and MLP* 100K 2.80
+16 Ours 100K 2.81
 Table 3: Impact of initializing various sub-modules within a transformer block.
 We compare validation loss of a 16-layer GPT-2 Medium variant when different sets of sub-modules are initialized with weights from the first 16 layers of a 24-layer GPT-2 Medium reference model.
-All models are trained on the OpenWebText dataset.
-Key findings: (1) $\mathsf{Inheritune}$ initialization and attention + MLP initialization result in similar performance improvements; (2) layernorm initialization shows minimal impact.
-A detailed training curve is shown in Figure 15.
-We conducted extensive experiments to better understand which sub-module initializations within a transformer block lead to improved generalization (in terms of validation loss) and faster convergence.
-For these ablations, we fixed the model to a 16-layer GPT-2 medium variant and explored three different sub-module initializations using weights from a 24-layer GPT-2 medium reference model.
-We initialize the transformer blocks with 1) attention ((key, query, value, and projection) and the layernorm [^2] In GPT-2 models layernorm blocks are parameterized.
-weights (attn w/ layernorm), 2) attention and mlp weights without the layer-norm (attn+mlp w/o layernorm), and 3) mlp weights with the layer norm (mlp w/ layernorm).
-We emphasize that $\mathsf{Inheritune}$ performs initialization by inheriting attention and mlp weights with the layer norm (attn+mlp w/ layernorm).
-As shown in Table 2, models trained with attention and mlp weights demonstrated the best performance, regardless of the layer norm initialization.
-A detailed validation loss vs training steps plot is presented in supplementary Figure 15.
-We conclude that initializing both attention and MLP weights provides a clear advantage.
-Surprisingly, we also observed that initializing either the attention or mlp weights resulted in similar improvements in both convergence speed and final validation loss.
+Submodules without the * marker also include layernorm weights.
+All models are trained with the OpenWebText dataset.
+Key findings: (1) $\mathsf{Inheritune}$ initialization and Attention and MLP initialization result in similar performance improvements; (2) layernorm initialization shows minimal impact.
+The training curves of corresponding models are presented in Figure 19.
+The lowest and our corresponding validation losses (lower is better) are highlighted in bold.
+(a) $\tau$=0.80
+(b) $\tau$=0.85
+(c) $\tau$=0.90
+(d) $\tau$=0.95
+(e) $\tau$=0.80
+(f) $\tau$=0.85
+(g) $\tau$=0.90
+(h) $\tau$=0.95
+Figure 10: Lazy layers remain robust under variations of $\tau$.
+To study the sensitivity of attention rank to the variance–ratio threshold $\tau\in\{0.8,0.85,0.9,0.95\}$, we visualize $\mathrm{MaxRank}(l)$ as a function of layer index for (a–d) the GPT-2 Medium full model with 24 layers (top row) and (e–h) the GPT-2 Large full model with 36 layers (bottom row).
 5 Related Works
@@
 (2022); Barbero et al.
+(2024); Wu et al.
 (2024).
 Recently He et al.
-(2023) address rank collapse in self-attention networks (SANs) without residual connections or layer norms, using two model initialization techniques that enable faithful signal propagation—i.e., $\Sigma_{L}$ of $A(X^{L})$ does not collapse in deeper layers.
+(2023) address rank collapse in self-attention networks (SANs) without residual connections or layer norms, using two different model initialization techniques that enable faithful signal propagation—i.e., $\Sigma_{L}$ of $A(X^{L})$ does not collapse in deeper layers.
 However, this approach significantly slows down training.
@@
 (2022) proposes scaling residual connections by $1/\sqrt{L}$, while Barbero et al.
-(2024) suggest that adding additional tokens to already long sequences of repeated tokens can help mitigate degeneration.
+(2024) suggest that adding additional tokens to already long sequences of repeated tokens can help mitigate collapse.
 In contrast to prior works, we address attention degeneration by developing smaller models that eliminate structural inefficiencies and training these models to match the performance of their larger, inefficient counterparts.
@@
 Knowledge distillation Hinton et al.
-(2015) has been very successful in training small LMs in some cases Turc et al.
+(2015) has been very successful in training small LMs.
+In some cases Turc et al.
 (2020); Sanh et al.
 (2019) the smaller student model is also initialized with teacher layers, though this is often done without clear explanation or intuition.
-Recent works in model initialization, such as Trockman & Kolter (2023), have studied synthetic attention patterns for initialization, primarily in vision settings.
+Recent works in model initialization, such as Trockman and Kolter (2023), have studied synthetic attention patterns for initialization, primarily in vision settings.
 However, such methods have limited success in language models.
@@
 6 Conclusion
-In this paper, we identified a structural flaw in the attention mechanism of deep decoder-style LLMs, where many deeper layers tend to lose rank and converge into near rank-1 matrices.
-To address this, we propose $\mathsf{Inheritune}$, to train smaller models that inherits early blocks from a larger model and expands the architecture gradually, matching the performance of the reference model.
-We validated $\mathsf{Inheritune}$ on GPT-2 models of varying sizes, achieving efficient smaller models without performance loss on the OpenWebText and FineWeb_edu datasets.
+In this work, we identify a structural inefficiency in deep decoder-style LLMs, which we term attention collapse, where attention matrices in deeper layers often degenerate into near rank-one structures, rendering these layers ineffective.
+These ineffective layers referred to as lazy layers contribute little to the model’s representational power.
+To address this, we introduce $\mathsf{Inheritune}$, a multi-stage training recipe that initializes a smaller model using a few potent early layers from a larger pre-trained model and then progressively trains and expands it through multiple rounds.
+Our experiments demonstrate that models trained with $\mathsf{Inheritune}$ can match or even surpass the performance of their larger counterparts despite having significantly fewer layers.
+By mitigating attention collapse, our approach produces compact and highly performant models, offering a new path toward designing smaller, more attentive architectures from the ground up.
+Acknowledgments
+This research was supported by NSF Grant 2217069 and by UT Austin’s Center for Generative AI and the Machine Learning Lab.
+We thank Nived Rajaraman for helpful discussions and the anonymous reviewers for their valuable feedback.
 Supplementary Materials
 Contents
-  • A: Extended Discussion on Attention Collapse
-  • B: Baselines
-  • C: Supplementary Experiments
-  • D: Extended Discussion
-  • E: Additional Experiments on Low Data Regime
-  • F: Implementation Details
-Appendix A Extended Discussion on Attention Collapse
-A.1 Attention Mass Analysis
-Dominant Single-Column Structure in Attention Matrices.
+  • A: Frequently Asked Questions
+  • B: Extended Discussion on Attention Collapse
+  • C: Understanding Attention Degradation using Attention Mass Analysis
+  • D: Baselines
+  • E: Supplementary Experiments
+  • F: Model Configurations and Training Details
+  • G: Additional Experiments and Discussions
+  • H: Theoretical Analysis
+Appendix A Frequently Asked Questions
+A.1 Is your method not just pruning?
+$\mathsf{Inheritune}$ is a stage-wise efficient training recipe that addresses a structural issue in decoder-style transformer blocks—Attention collapse—which we consistently observed across multiple models.
+Unlike pruning, Inheritune includes a growth phase where the model is expanded until it outperforms the reference model (refer Algorithm 1 and Figure 20).
+Pruning doesn’t always require re-training (Ma et al., 2023), whereas $\mathsf{Inheritune}$ may need multiple rounds of re-training.
+To the best of our knowledge, no pruning method has explicitly studied or resolved attention collapse in LLMs.
+Our method has closer proximity to efficient training recipes employing model initialization (J.
+Reddi et al., 2023; Du et al., 2024) or warm starting (Ash and Adams, 2020).
+A.2 Is the comparison with baseline models trained from scratch unfair since Inheritune uses weights from pre-trained models?
+In Baseline-II (refer Section 4) we have also compared our method with warm started baselines which also uses pre-trained model weights for model initialization for fair comparisons.
+For Baseline-I, we include much larger reference models as well as same-sized models trained for twice as many steps.
+Remarkably, $\mathsf{Inheritune}$ still outperforms both.
+We believe these findings are novel and reveal a new axis for scaling.
+A.3 The attention collapse analysis is not holistic?
+TL;DR.
+We analyzed the phenomenon of attention collapse across four datasets and four different model architectures, with model sizes ranging from millions to billions of parameters.
+We evaluated attention patterns and analyzed the phenomenon of attention collapse using four datasets: OpenWebText, FineWeb, RedPajama, and C4.
+Our analysis used two complementary metrics namely, approximate rank and approximate mass to quantify the structure of attention matrices.
+We conducted attention collapse analysis on a range of models, including GPT-2 (Medium 355M, Large 770M, XLarge 1.5B), LLaMA-3 (3B and 8B models ), OLMo 1B, Cerebras-GPT 2.7B, Falcon 7B and LLaMA-1 (OpenLLaMA 3B, 7B, 13B) which features a notably different architecture.
+The evaluation was performed on 10K tokens (100 samples $\times$ 100 tokens each), and we also provide visualizations of the resulting attention patterns.
+A.4 What is the connection between Attention collapse and Attention sinks?
+The term Attention sink (Xiao et al., 2024) refers to a phenomenon where a specific token in a sequence receives disproportionately high attention scores compared to other tokens in the attention map.
+In our analysis of Attention collapse, we also observed sink-like behavior for certain tokens across all attention maps (see Figure 9 and Figure 12).
+However, unlike typical attention sinks, we found that beyond the sink token, no meaningful attention structure remains: all other tokens receive nearly uniform attention scores.
+We further connect this behavior to the emergence of lazy layers.
+Therefore our analysis has unique insights compared to attention sinks.
+Appendix B Extended Discussion on Attention Collapse
+B.1 Attention Collapse in LLaMA-3 models
+We conducted a rank analysis on a contemporary LLaMA-3 base model with 8B parameters.
+We compute $\text{Rank}^{(h,l)}=\frac{1}{N}\sum_{n=1}^{N}k^{*}_{n,h,l}$, where $N=100$ sequences are sampled from a subset of the C4 dataset (Raffel et al., 2019).
+As shown in Figure 11, we observe that nearly 50% of the attention heads (500 out of 1024 across all layers) are close to near rank-1, highlighted in red.
+This presents an interesting case: in very large modern architectures such as LLaMA-3 8B, while there may not be entire lazy layers, a substantial number of heads within many layers exhibit degeneracy.
+A different pattern of attention collapse compared to GPT-2 models can be attributed to the architectural differences between these models.
+Figure 11: Rank analysis of LLaMA-3 8B reveals that nearly half of the attention heads exhibit rank collapse.
+We analyze the LLaMA-3 8B model, which contains 32 heads per layer (32 $\times$ 32), using the rank metric defined in Section 2.
+The results are visualized as a heatmap of head index vs.
+layer index.
+Potent (non-collapsed) heads are shown in varying shades of green, where higher intensity indicates higher rank, while rank-collapsed heads (near rank-1) are highlighted in red.
+Approximately 50% of all attention heads exhibit rank collapse, indicating widespread degeneracy.
+B.2 Attention Pattern Visualization of Potent and Lazy Layers
+Following the analysis in Section 2 and Figure 1, we present the attention patterns of two representative layers from the GPT-2 XL model: a lazy layer (Layer 30) and a potent layer (Layer 8), as shown in Figure 12.
+The attention patterns of these layers exhibit distinctly different behaviors.
+In particular, the lazy layer demonstrates a clear collapse, where attention concentrates almost exclusively on the first token.
+Next, following the discussions in Section 4.2, we visualize the rank–layer relationship for the GPT-2 XLarge model, juxtaposing a vanilla model with a $\mathsf{Inheritune}$-trained model containing half as many layers.
+Although both models achieve similar validation loss, the $\mathsf{Inheritune}$-trained model exhibits significantly fewer lazy layers compared to the vanilla counterpart.
+L8 H16
+L8 H17
+L8 H18
+L8 H19
+L8 H20
+A Potent layer of vanilla GPT-2 XLarge.
+L30 H16
+L30 H17
+L30 H18
+L30 H19
+L30 H20
+A lazy layer of vanilla GPT-2 XLarge.
+Figure 12: Visualization of attention patterns in lazy and non-lazy layers of a vanilla GPT-2 XLarge model with 48 layers.
+The top row displays attention patterns for various heads (H) in layer (L) 8, while the bottom row shows patterns for layer (L) 30.
+For visual clarity, we display the full attention maps; however, attention in GPT-2 models is inherently causal.
+Figure 13: Rank collapse worsens for larger LLMs, Inheritune helps to mitigate rank collapse.
+The maximum (max) rank across all attention heads for each layer is plotted, following the methodology in Figure 1.
+We analyze a 48-layer GPT-2 XLarge model which reveals rank-1 attention matrices in later layers (those beyond the halfway point), indicating rank collapse.
+Specifically, 22 out of the last 24 later layers exhibit rank-1 attention matrices (mean rank across all the 100 runs).
+Next, Our 24-layer GPT-2 XLarge variant, trained with $\mathsf{Inheritune}$, demonstrates improved rank across many layers, highlighting the effectiveness of our approach.
+Notably, 2 out of 12 of the later layers in our 24-layer variant exhibit rank-1 attention matrices.
+Appendix C Understanding Attention Degradation using Attention Mass Analysis
+In this paper, we have analyzed the attention degradation phenomenon primarily using a single metric-rank of the attention matrices (see Section 2).
+In this section, we aim to explore another thematically related metric to further investigate the nature of attention degradation.
 We further investigated the dominant structure of the rank-1 attention matrices and observed that, on an average, many of these matrices have their mass concentrated in a single column.
@@
 In Figure 14, we present the layer-wise analysis of the attention matrix mass concentration in GPT-2 models.
-For this analysis, (similar to the rank analysis), we computed $A(X)$ using $N=100$ sequences selected at random from the validation set of OpenWebText (4M tokens), each with a sequence length of $T=100$ tokens across all attention heads within each layer.
+For this analysis, (similar to the rank analysis), we computed $A(X)$ using $N=100$ sequences selected at random from the validation set of OpenWebText (4.4M tokens), each with a sequence length of $T=100$ tokens across all attention heads within each layer.
 We define the average minimal column count $m$ required to capture 90% of the attention matrix mass for each head and layer as: $m^{(h,l)}=\frac{1}{N}\sum_{n=1}^{N}m^{*}_{n,h,l}$.Subsequently, we aggregate this metric per layer by taking the average across all heads: $\text{AvgMass}^{(l)}=\frac{1}{H}\sum_{h^{\prime}=1}^{H}m^{(h^{\prime},l)}$.
 We observe that many of the rank-collapsed attention matrices in deeper layers exhibit single-column attention structures, as measured by the $\text{AvgMass}^{(l)}$ criterion.
-As shown in Figure 9, we performed a mass analysis on contemporary billion-parameter LLaMA models (Geng & Liu, 2023) and observed a similar pattern of attention degradation in the deeper layers.
-This provides concrete evidence that the phenomenon persists across a broad range of architectures at the billion-parameter scale.
-(a) GPT-2 Medium (355M) L=24, H=1024
-(b) GPT-2 Large (770M) L=36, H=1280
-(c) GPT-2 xLarge (1.5B) L=48, H=1600
+As shown in Figure 15, we performed a mass analysis on contemporary billion-parameter OpenLLaMA models (Geng and Liu, 2023) and observed a similar pattern of attention degradation in the deeper layers.
+This provides concrete evidence that the phenomenon persists across a broad range of architectures and also at the billion-parameter scale.
+(a) GPT-2 Medium
+(b) GPT-2 Large
+(c) GPT-2 XLarge
 Figure 14: In decoder-style LLMs, attention matrices in deeper layers often degenerate to near single column matrices, which is a special case of near rank-1.
@@
 We followed the same procedure as discussed in Section 2.
-Pre-trained checkpoints of OpenLLaMA-3B, OpenLLaMA-7B, and OpenLLaMA-13B (Geng & Liu, 2023), trained on 1T tokens from the RedPajama dataset Computer, 2023, were utilized.
+Pre-trained checkpoints of OpenLLaMA-3B, OpenLLaMA-7B, and OpenLLaMA-13B (Geng and Liu, 2023), trained on 1T tokens from the RedPajama dataset Computer, 2023, were utilized.
 Overall, we observed that 90 % of the total mass of the attention matrices resides in fewer columns, with many attention matrices in the OpenLLaMA-13B model being single-column.
 This observation aligns closely with our analysis in Figure 1.
-A.2 Additional Analysis on Attention Rank Collapse.
-To investigate the robustness of our analysis, we performed a rank analysis on GPT-2 models trained on OpenWebText (as shown in Figure 1) using a validation set from FineWeb-Edu—a new dataset on which the models were not trained (see Figure 10).
-It is clearly observed that attention consistently loses rank, often collapsing to single-column structures across all models, thereby reinforcing the robustness of our findings.
-Next, we conducted a rank analysis on a contemporary LLaMA-3 model with 8B parameters.
-We compute $\text{Rank}^{(h,l)}=\frac{1}{N}\sum_{n=1}^{N}k^{*}_{n,h,l}$, where $N=100$ sequences are sampled from a subset of the C4 dataset (Raffel et al., 2019).
-As shown in Figure 11, we observe that nearly 50% of the attention heads (500 out of 1024 across all layers) are close to near rank-1, highlighted in red.
-This presents an interesting case: in very large modern architectures such as LLaMA-3 8B, while there may not be entire lazy layers, a substantial number of heads within many layers exhibit degeneracy.
-A.3 Attention Pattern Visualization
-To provide further evidence of lazy layers, we visualized attention patterns across various layers of a vanilla 24-layer GPT-2 medium model.
-Fig.
-12 shows the attention patterns for both a vanilla 24-layer model trained from scratch and a 16-layer model trained using our proposed method, $\mathsf{Inheritune}$.
-Note just for the sake of better visualization we visualized full attention and not causal attention, in practice GPT-2 models computes causal attention.
-We computed these attention matrices using randomly selected strings from the val set of OpenWebText and took 40 tokens averaged over 3 runs.
-In the 24-layer model trained from scratch (top row of Fig.
-12), we observe a clear progression in attention patterns.
-The early layers (L4 and L7) exhibit structured patterns with a mix of local and global attention Gong et al.
-(2019); Beltagy et al.
-(2020); Chen et al.
-(2021).
-In contrast, the deeper layers (L20 and L22) display more uniform patterns, indicating a loss of focus.
-This uniformity is a hallmark of lazy layers, where the attention mechanism loses its ability to selectively focus on specific relevant tokens.
-In contrast, our 16-layer model trained with $\mathsf{Inheritune}$ (bottom row) demonstrates more focused and effective attention patterns, even in its later layers (L11 and L15).
-This striking difference suggests that our method makes model more attentive and addresses attention degeneration, potentially leading to more efficient models in compact size (also refer Figure 17 and Figure 18).
-Additionally, in Figure 13, we present attention patterns from a 48-layer GPT-2 xLarge model, comparing attention maps from a potent (non-lazy) and a lazy layer.
 (a) Rank analysis of GPT-2 Medium
@@
 (c) Rank analysis of GPT-2 XLarge
-(d) Matrix mass analysis of GPT-2 Medium
-(e) Matrix mass analysis of GPT-2 Large
-(f) Matrix mass analysis of GPT-2 XLarge
+(d) Mass analysis of GPT-2 Medium
+(e) Mass analysis of GPT-2 Large
+(f) Mass analysis of GPT-2 XLarge
 Figure 16: In standard decoder-style LLMs, attention matrices in deeper layers often degenerate into single-column matrices, leading to layers with fully degenerated attention that fail to learn meaningful representations.
-All models were trained on the OpenWebText dataset, and both rank and mass analyses were conducted using the Fineweb_edu validation set, following the same procedure described in Figure 1.
+All models were trained on the OpenWebText dataset, and both rank and mass analyses were conducted using the FineWeb validation set, following the same procedure described in Figure 1.
 This further demonstrates the robustness of our analysis, as we reach the same conclusion using different evaluation datasets.
-Figure 11: Rank analysis of LLaMA-3 8B reveals that nearly half of the attention heads exhibit rank collapse.
-We analyze the LLaMA-3 8B model, which contains 32 heads per layer (32 $\times$ 32), using the rank metric defined in Section 2.
-The results are visualized as a heatmap of head index vs.
-layer index.
-Potent (non-collapsed) heads are shown in varying shades of green, where higher intensity indicates higher rank, while rank-collapsed heads (near rank-1) are highlighted in red.
-Approximately 50% of all attention heads exhibit rank collapse, indicating widespread degeneracy.
-L4 H2
-L7 H3
-L20 H2
-L22 H4
-24 layer GPT2 medium model trained from scratch.
-L4 H2
-L7 H3
-L11 H2
-L15 H1
-Our 16-layer GPT2 medium model trained using $\mathsf{Inheritune}$.
-Figure 12: $\mathsf{Inheritune}$ preserves effective attention patterns in smaller models.
-Comparison of attention patterns across layers (L) and heads (H) in two GPT2-medium models: (top) 24-layer model trained from scratch, (bottom) 16-layer model trained with $\mathsf{Inheritune}$.
-Attention maps are averaged over three randomly selected string, with 40 tokens each from the validation.
-Darker colors indicate higher attention scores.
-$\mathsf{Inheritune}$ maintains focused attention even in deeper layers, contrasting with the uniform patterns in the standard model’s deeper layers.
-L8 H16
-L8 H17
-L8 H18
-L8 H19
-L8 H20
-A Potent (non-lazy) layer of a pre-trained GPT2 xLarge 48 layer model.
-L30 H16
-L30 H17
-L30 H18
-L30 H19
-L30 H20
-A lazy layer of a pre-trained GPT2 xLarge 48 layer model.
-Figure 13: Visualization of attention patterns in lazy and non-lazy layers of a vanilla GPT-2 xLarge model with 48 layers.
-The top row displays attention patterns for various heads (H) in layer (L) 8, while the bottom row shows patterns for layer (L) 30.
-Appendix B Baselines
+C.1 Data Robustness of Attention Degeneration Analysis
+In Section 2 (Figure 1), we performed a rank analysis on three pre-trained GPT-2 models—Medium, Large, and XLarge using the validation set of the OpenWebText dataset, whose training split was originally used for pre-training these models.
+Here, we evaluate the data robustness of our analysis by repeating the same procedure on a validation set from FineWeb, a newer and distinct dataset.
+Except for the dataset substitution, all experimental steps remain identical to those described earlier.
+The results in Figure 16 consistently show that attention tends to lose rank, particularly in deeper layers, often collapsing into near single-column structures across all models.
+These findings further reinforce the robustness and generality of our observations.
+Moreover, for the LLaMA-3 model (Figure 11) and the OpenLLaMA models (Figure 15), we used publicly available model weights and conducted our analyses on off-the-shelf datasets that were not part of the models’ original training corpora.
+Appendix D Baselines
 We compare $\mathsf{Inheritune}$ against several baselines.
 While some baseline methods are illustrated using GPT-2 Large or medium (for the knowledge distillation baseline) as an example, the same methodology is consistently applied across all model variants.
-Baselines trained from scratch (rand init.) :
+Baselines trained from scratch (random initializations) :
 We compare our $\mathsf{Inheritune}$-derived model against much larger GPT-2 reference models trained from scratch for the same number of steps and similar-sized models trained from scratch for both the same and double the number of training steps.
-Baselines trained with various model initialization and efficient training techniques.
-Here we compare our model derived using $\mathsf{Inheritune}$, to similar sized models trained with various zeroshot model initialization and effcient training techniques such as stacking, hybrid stacking, and half-width initialization.
+Baselines trained with warm started training methods.
+Here we compare our model derived using $\mathsf{Inheritune}$, to similar sized models trained with various model initialization and effcient training techniques which requires model to be initialized with trained weights such as stacking, hybrid stacking, and half-width.
 We explain these baseline training recipes using GPT-2 Large and its variants as an example and apply the same process for other models.
@@
 Both baselines are trained for 50K steps, using a vanilla 24-layer GPT-2 Medium model as the teacher (our reference model).
-Appendix C Supplementary Experiments
-We provide additional training data and plots for our main results discussed in Section 4.1 as shown in Table 3 and Figure 14.
-In Figure 5 (also refer Table 3 we compare our GPT-2 variants with baseline models trained form scratch.
-In Figure 14 (also refer Table 1) we compare our GPT-2 variants with baseline models trained using baseline zero-shot model initialization (and also re-training) techniques.
-In Figure 15, we present the training curves of models trained during ablation as discussed in Section 4.2.
-Knowledge Distillation
-Recall we have already discussed distillation as a baseline in Section 4.1 and associated Figure 7.
-We perform an additional experiment in the same setting i.e.
-knowledge distillation as a baseline.
-Here we trained GPT-2 medium variants with 12 layers (half the number of a vanilla GPT-2 medium).
-We trained three models.
-First we distilled a 24-layer GPT-2 medium (teacher) to a 12-layer GPT-2 medium variant (student) and this student is initialized with all the alternate layers of the teacher.
-This setting is exactly same as discussed in DistillBERT Sanh et al.
-(2019).
-Next we trained two GPT-2 medium variants one from scratch (vanilla training) and the other with $\mathsf{Inheritune}$ recipe.
-Model trained with our recipe beats model trained with distillation.
-We defer a through investigation of distillation compared to $\mathsf{Inheritune}$ to future work.
-How Inheritune addresses Attention Degeneration?
-Recall we have discussed attention degeneration in Section 2 and attention patterns are visualized in Figure 12.
-Following up on our previous discussions in Figure 17 and Figure 18 we demonstrate that models trained with $\mathsf{Inheritune}$ has lesser lazy layers compared to it’s larger counterpart trained form scratch.
-We performed rank analysis for Figure 17 utilizing vanilla 24-layer GPT-2 medium and our 16-layer GPT-2 variant trained using $\mathsf{Inheritune}$.
-Additionally, we performed rank analysis for Figure 18 with a vanilla 48-layer GPT2 xLarge and a 24-layer GPT2 xLarge variant trained using $\mathsf{Inheritune}$.
-Recall we have previously discussed that attention degeneration is connected with vanishing gradients of keys and queries Noci et al.
-(2022).
-The vanishing gradients is caused when the norm of the gradients Bengio et al.
-(1994) are so small that it fails to generate meaningful back-propagation signal.
-Since we are training smaller models intuitively $\lVert W_{Q}\rVert$ and $\lVert W_{K}\rVert$ should be smaller compared to their larger counterparts and hence the norm of gradients in the case of smaller models derived using $\mathsf{Inheritune}$ is higher leading to better training.
+Layers Initialization AvgRank Pre-train Val Loss ($\downarrow$)
+4 rand N/A 3.25
+4 1-4 layers from vanilla GPT2 8.40 3.22
+4 5-8 layers from vanilla GPT2 9.48 3.19
+4 9-12 layers (lazy layers) from GPT2 1.22 3.23
+Table 4: Impact of initialization strategies on GPT-2 small variants.
+We analyzed the rank characteristics of a vanilla GPT2-small model (125M, 12 layers) trained on OpenWebText for 100K steps.
+Four-layer GPT2-small variants were initialized using the first 4 layers [1–4], middle 4 layers [5–8], last 4 layers [9–12], or with random initialization, and then trained for 100K steps on OpenWebText.
+Models initialized with the last 4 layers performed similarly to random initialization, while those initialized with layers exhibiting higher average max ranks achieved the best validation loss, regardless of proximity to the embedding layer.
+The training plots and rank analysis are provided in Figure 2.
+Appendix E Supplementary Experiments
+We provide supplementary tables and plots corresponding to the results discussed in the main paper, along with additional experiments, in this section.
+The final validation losses shown in Figure 2 are presented in Table 4.
+Similarly, the final validation losses from the training curves in Figure 5 and Figure 6 are summarized in Table 5.
+We also include training curves for models trained using various warm-started baselines (i.e., models initialized with learned weights) compared against our method, these results correspond to Table 2 and are visualized in Figure 18.
+Finally, Figure 19 presents the training curves for models used in the ablation study discussed in Section 4.3.
+Distillation vs. Inheritune.
+We conducted an additional experiment to compare $\mathsf{Inheritune}$ with knowledge distillation as a baseline.
+Specifically, we trained GPT-2 Medium variants with 16 layers under three different settings.
+First, we performed logit-based distillation Hinton et al.
+(2015), transferring knowledge from a 24-layer vanilla GPT-2 Medium (teacher) to a 16-layer student model.
+Second, we applied a DistilBERT-style distillation Sanh et al.
+(2019), where the student is initialized with the teacher’s layers.
+Finally, we trained a 16-layer GPT-2 Medium model from scratch using vanilla training.
+Across all comparisons, the model trained with our $\mathsf{Inheritune}$ recipe outperformed both distilled variants, achieving faster convergence and substantially better generalization after 50K training steps.
+We defer a thorough investigation of the relationship between $\mathsf{Inheritune}$ and distillation-based approaches to future work.
+The training configurations are provided in Section F.
+Figure 17: A 16-layer GPT-2 Medium variant derived using $\mathsf{Inheritune}$ converges faster and generalizes better than a same-sized model trained with logit-based distillation baselines.
 Models Layers Initialization Steps Pre-train Val loss ($\downarrow$)
@@
 16 rand init 100K 2.86
+16 rand init 200K 2.83
 12 Ours 100K 2.87
@@
 18 rand init 100K 2.97
+18 rand init 200K 2.84
 18 Ours 100K 2.80
@@
 24 rand init 100K 2.69
+24 rand init 200K 2.62
 24 Ours 100K 2.64
 Table 5: Inheritune achieves superior performance with reduced model size.
-Comparison of $\mathsf{Inheritune}$-trained models (24-layer GPT-2 xLarge, 18-layer GPT-2 Large, and 16-layer GPT-2 Medium) against full-sized counterparts.
-All models are trained for 100K steps, and reported using validation loss on the OpenWebText validation set.
-Note: GPT-2 Large and xLarge use one round of $\mathsf{Inheritune}$; GPT-2 Medium uses three rounds.
-(a) GPT2-Medium
-(b) GPT2-Large
-(c) GPT2-xLarge
-Figure 14: Models derived using Inheritune outperform three zero-shot initialization and efficient training baselines in terms of final validation loss.
+Comparison of $\mathsf{Inheritune}$-trained models (24-layer GPT-2 XLarge, 18-layer GPT-2 Large, and 16-layer GPT-2 Medium) against full-sized counterparts and extended training baselines.
+The training steps of two different baselines are reported in the table, we use validation loss on the OpenWebText validation set.
+Note: GPT-2 Large and XLarge uses one round of $\mathsf{Inheritune}$; GPT-2 Medium uses three rounds.
+The lowest and our corresponding validation losses (lower is better) are highlighted in bold.
+(a) GPT-2 Medium
+(b) GPT-2 Large
+(c) GPT-2 XLarge
+Figure 18: Models derived using Inheritune outperform three warm-started baselines (Baseline-II) in terms of final validation loss.
 Our models demonstrate better convergence and generalization compared to all baselines.
-We trained GPT-2 xlarge, GPT-2 large and GPT-2 medium variants on OpenWebText-9B for 100K steps using baseline model initialization and efficient training techniques and our $\mathsf{Inheritune}$ training recipe.
+All the models are trained with OpenWebText for 100K steps.
+The curves are smoothed for visual clarity.
 Figure 19: Full training curves of 16-layer GPT-2 variants trained during ablations.
@@
 Finally, we perform $\mathsf{Inheritune}$-initialize only the attention and MLP weights with all the respective layer norms.
-Figure 16: A 12-Layer GPT-2 medium variant derived using Inheritune converges faster and generalizes better than a same-sized models trained from scratch and with Logit-based distillation with teacher initialization baseline.
-Three 12-layer GPT-2 medium variants were trained: (1) a distilled model initialized with alternate layers from a 24-layer GPT-2 medium teacher, following the DistillBERT setup Sanh et al.
-(2019); (2) a model trained from scratch (vanilla training); and (3) a model trained using the $\mathsf{Inheritune}$ recipe.
-The model trained with $\mathsf{Inheritune}$ outperforms both the distillation-based model and the one trained from scratch, demonstrating the effectiveness of our approach.
-(a) Vanilla GPT2-Medium 24 layers
-(b) Our GPT2-Medium 16 layer variant trained with $\mathsf{Inheritune}$
-Figure 17: Rank collapse in deeper layers and its mitigation through $\mathsf{Inheritune}$.
-The maximum (max) rank across all attention heads for each layer is plotted, following the methodology in Fig.
-1 (a) Analysis of a 24-layer GPT2 medium model reveals rank-1 attention matrices in later layers (those beyond the halfway point), indicating rank collapse.
-Specifically, 3 out of the last 12 later layers exhibit rank-1 attention matrices (mean rank accross all the 100 runs).
-(b) Our 16-layer GPT2 medium variant, trained with $\mathsf{Inheritune}$, demonstrates improved rank across all layers, highlighting the effectiveness of our approach.
-Notably, none of the later layers in our 16-layer variant exhibit rank-1 attention matrices.
-(a) Vanilla GPT2-xLarge 48 layers
-(b) Our GPT2-xLarge 24 layer variant trained with $\mathsf{Inheritune}$
-Figure 18: Rank collapse worsens for larger LLMs, $\mathsf{Inheritune}$ helps to mitigate rank collapse.
-The maximum (max) rank across all attention heads for each layer is plotted, following the methodology in Fig.
-1 (a) Analysis of a 48-layer GPT2 xLarge model reveals rank-1 attention matrices in later layers (those beyond the halfway point), indicating rank collapse.
-Specifically, 22 out of the last 24 later layers exhibit rank-1 attention matrices (mean rank across all the 100 runs).
-(b) Our 24-layer GPT2 xLarge variant, trained with $\mathsf{Inheritune}$, demonstrates improved rank across all layers, highlighting the effectiveness of our approach.
-Notably, 2 out of 12 of the later layers in our 24-layer variant exhibit rank-1 attention matrices.
-Appendix D Extended Discussion
-D.1 Discussion about Attention Sink
-The term "attention sink" Xiao et al.
-(2024) refers to the phenomenon where the first token in a sequence receives disproportionately high attention scores compared to other tokens in the attention maps.
-While there is some connection with $\mathsf{Inheritune}$, as we have also observed that many attention matrices are not only rank-1 but also single-column (with all attention scores concentrated on the first token), this connection has not been explicitly established in Xiao et al.
-(2024) with respect to rank-1 behavior or poor training of later layers.
-In contrast, as illustrated in Figure 1, we compute the maximum rank of all attention matrices within a layer.
-For instance, consider a layer where only 2 out of 5 attention heads exhibit attention sink behavior.
-This does not make the layer lazy, as attention is computed as a concatenation of activations across all heads.
-A lazy layer, however, has all 5 out of 5 attention heads fully degenerated, with their attention matrices being rank-1.
-We have provided evidence that such lazy layers are indicative of poorly trained layers.
-Layers Initialization Avg max ranks Val Loss ($\downarrow$)
-4 rand n/a 3.25
-4 1-4 layers from vanilla GPT2 8.40 3.22
-4 5-8 layers from vanilla GPT2 9.48 3.19
-4 9-12 layers (lazy layers) from GPT2 1.22 3.23
-Table 4: Impact of initialization strategies on GPT2-small variants.
-We analyzed the rank characteristics of a vanilla GPT2-small model (125M, 12 layers) trained on OpenWebText for 100K steps.
-Four-layer GPT2-small variants were initialized using the first 4 layers [1–4], middle 4 layers [5–8], last 4 layers [9–12], or with random initialization, and then trained for 100K steps on OpenWebText.
-Models initialized with the last 4 layers performed similarly to random initialization, while those initialized with layers exhibiting higher average max ranks achieved the best validation loss, regardless of proximity to the embedding layer.
-The training plots and rank analysis are provided in Figure 2.
-Models Recipe Layers ARCE PIQA SciQ Hellaswag Lambada Average
-(acc) (acc) (acc) (acc norm) (acc)
-GPT-2 Medium rand init 24 51.05 61.81 74.8 30.79 20.28 47.74
-rand init 16 49.92 61.92 73.3 29.56 19.54 46.84
-Ours 16 51.26 61.81 73.8 30.55 23 48.08
-GPT-2 Large rand init 32 52.48 64.58 75.3 32.65 22.2 49.44
-rand init 16 50.34 63.11 75 30.86 21.56 48.17
-Ours 16 52.9 63.55 76.1 32.14 24.06 49.75
-Table 5: Models trained with Inheritune outperforms both their larger and same-size counterparts trained from scratch on average zero-shot downstream performance.
-For evaluation we choose accuracy (acc) and normalized accuracy (acc norm) metrics following Open LLM leaderboard Beeching et al.
+Figure 20: Training curves for the 24-layer GPT-2 Medium model (full model) and three rounds of training following Inheritune recipe (to grow the model).
+We present the training trajectories for all GPT-2 Medium variants trained using the $\mathsf{Inheritune}$ recipe.
+The final model obtained after the third round ($R=3$) with $L=16$ layers matches the final validation loss of the full model.
+All models are trained for 100K steps on the OpenWebText dataset.
+Appendix F Architectural and Training Details
+F.1 GPT-2 Model configurations
+For our main experiments, we focus on three sizes of GPT-2 models Radford et al.
+(2019): GPT-2 XLarge with 1.5B parameters, GPT-2 Large with 770M parameters, and GPT-2 Medium with 355M parameters.
+We developed several variants of these models by adjusting the number of layers, i.e., reducing the depth for vanilla models for to be trained with $\mathsf{Inheritune}$ and baseline mathods.
+In one baseline namely, the half-width variant we modified both the hidden size (and consequently, the number of attention heads) in addition to the depth, as shown in Figure 2.
+The key architectural configurations of the reference, proposed, and baseline models discussed in this paper are summarized in Table 6.
+Model Family Type Layers Hidden Size Heads Notes
+GPT-2 XLarge (1.5B) Reference 48 1600 25 Original architecture
+Variants 24 1600 25 Reduced-depth
+GPT-2 Large (770M) Reference 36 1280 20 Original architecture
+Variants 18 1280 20 Reduced depth
+GPT-2 Large (668M) Reference 32 1280 20 Custom architecture
+Variants 16 1280 20 Reduced depth
+GPT-2 Medium (355M) Reference 24 1024 16 Original architecture
+Variants 16 1024 16 Reduced depth
+GPT-2 Small (125M) Reference 12 768 12 Original architecture
+Variants 4 768 12 Reduced depth
+Table 6: Overview of all GPT-2 models used in this study and their architectural configurations.
+GPT-2 models are Pre-LN based architectures.
+The model configurations employed for the stacking and hybrid stacking baselines are identical to those of our variants.
+For the half-width baseline, we used GPT-2 variants with half the hidden size and number of attention heads.
+F.2 Training details of GPT-2 models
+All GPT-2 models used in this study (unless otherwise stated) were pre-trained on the OpenWebText dataset, which contains approximately 10B tokens.
+We employed a dataloader that samples tokens with replacement, meaning that the tokens used for training are not necessarily unique, following the approach of Liu et al.
 (2023).
-All the models are trained with FineWeb_edu.
-Appendix E Developing a 1.5B Small Base LM in a Low Data Regime with Inheritune
-Figure 19: Performance of our 1.5B base LM derived using 1B data with $\mathsf{Inheritune}$ on an average of 9 different datasets (left) and MMLU benchmark (right) that evaluates commonsense, truthfulness, natural language inference and language understanding.
-We compare our model’s performance with reference model-OpenLLamA-3B (2x size), other small base LMs of size 1B-2B parameters such as MPT-1.3B, OPT-1.3B, Pythia-1.4B (pre-trained from scratch) and ShearLLaMA-1.5B (pruned and continually trained using existing large base LM).
-In this section, we aim to investigate the efficacy of $\mathsf{Inheritune}$ in a data and compute-constrained setting.
-We train a 1.5B parameter small base LM with only 1B tokens using a 3B parameter base LM on a single GPU (A6000) for less than half a day.
-We assume the existence of a pre-trained reference model $\mathcal{M}_{\text{ref}}$, comprising $k$ layers, represented by $\mathsf{W}_{\text{ref}}=\{\mathsf{W}_{0},\mathsf{W}_{1},\ldots,\mathsf{W}_{k-1}\}$ trained with $\mathcal{D}_{\text{train}}$.
-However, this full training data is unavailable, and we only have a random tiny subset $\hat{\mathcal{D}}_{\text{train}}\sim\mathcal{D}_{\text{train}}$.
-We use OpenLLaMA-3B version 1 as the reference model pre-trained with 1T tokens from the RedPajama V1 dataset, which contains data from various domains such as common crawl, C4, Wikipedia, books, arXiv papers, GitHub, and Stack Exchange.
-We take 1B randomly sampled tokens [^3] https://huggingface.co/datasets/togethercomputer/RedPajama-Data-1T-Sample from the RedPajama dataset.
-Training recipe.
-To adapt $\mathsf{Inheritune}$ for this new setting, we perform step 1 and step 2 in Algorithm 1 without growing the model (i.e., we skip step 3).
-We use the first $n=13$ layers from our $k=26$ layer reference model.
-We call our small base LM Ours-1.5B(#tokens).
-We train our model with data repetition for eight epochs (each epoch uses all the 1B tokens) with a batch size of 131K tokens per batch.
-We use 1 A6000 GPU for less than half a day of training.
-The choice of training epochs is based on the analysis provided later in this paper (refer to Figure 22).
-We use the lit-gpt framework for training all small base LMs discussed in this paper.
-Further discussions on the training hyper-parameters can be found in the next Section.
-Baseline models and evaluation.
-We choose similarly sized (1-2B parameter) small base LMs trained with the RedPajama dataset and the reference base LM as primary baselines, as the quality of the pre-training data plays a key role in model development.
-We also include models OPT-1.3B Zhang et al.
-(2022) and Pythia-1.3B Biderman et al.
-(2023) as these models are pre-trained with a dataset similar to the RedPajama dataset.
-Table 7 lists the baseline models with their pre-training data.
-In this study, we use few-shot accuracy, particularly 0-shot and 5-shot accuracy, on ten different downstream tasks to measure the quality of our 1.5B base LM.
-This evaluation of pre-trained base LLMs has been done in several prior works.
-Our evaluation methodology categorizes downstream tasks across four distinct genres: commonsense reasoning, natural language understanding, factuality, and natural language inference.
-We perform 0-shot evaluation for PIQA Bisk et al.
-(2020), BOOLQ Clark et al.
-(2019), WINOGRANDE Sakaguchi et al.
-(2020), WINOGRAD Kocijan et al.
-(2020), LOGIQA Liu et al.
-(2020), TruthfulQA Lin et al.
-(2022), MNLI Bowman et al.
-(2015), QNLI Wang et al.
-(2018) and WNLI Wang et al.
-(2018) datasets.
-Next, we perform a 5-shot evaluation on the massive multitask language understanding benchmark (MMLU) Hendrycks et al.
-(2020).
-We use the lm eval harness framework Gao et al.
-(2024) for the entire evaluation.
-Model Commonsense Reasoning
-Name (# train tokens) Reference Winograd PIQA Boolq WinoGrande Logiqa
-OpenLLaMA-3B (1T) n/a 63.46 74.97 67.18 62.27 28.4
-OPT-1.3B (300B) n/a 38.46 71.82 57.83 59.51 27.04
-Pythia-1.4B (300B) n/a 36.54 70.89 63.12 56.99 27.65
-MPT-1.3B (200B) n/a 63.46 71.44 50.89 58.09 28.26
-Sheared LLaMA-1.3B (50B) LLaMA2-7B 36.54 73.45 62.02 58.17 27.34
-Ours-1.5B (1B) OpenLLaMA-3B 50.96 56.47 61.68 51.69 25.19
-Model Lang. Understanding & Inference Factuality
-Name ( # train tokens) Reference MMLU(5) WNLI QNLI MNLI TruthfulQA
-OpenLLaMA-3B (1T) n/a 27.21 50.7 51.3 37.3 35
-OPT-1.3B (300B) n/a 24.96 42.25 51.29 35.82 38.67
-Pythia-1.4B (300B) n/a 25.56 53.52 49.48 32.76 38.66
-MPT-1.3B (200B) n/a 25.82 40.85 50.52 35.93 38.68
-Sheared LLaMA-1.3B (50B) LLaMA2-7B 25.71 49.3 50.98 37.94 37.14
-Ours-1.5B (1B) OpenLLaMA-3B 25.67 43.66 49.41 34.42 48.61
-Table 6: Our 1.5B model achieves performance comparable to baseline models despite being trained with fewer tokens.
-Comparison of our target model ($\mathcal{M}_{\text{tgt}}$) derived using $\mathsf{Inheritune}$ with the reference model ($\mathcal{M}_{\text{ref}}$) and other baseline models of similar size when pre-trained from scratch and pre-trained with inherited weights and pruning.
-Although trained with fewer tokens, our model achieves performance comparable to the baseline models.
-We have highlighted all the scores in bold where our 1.5B model achieves at least 90% of the score compared to the reference LM or outperforms at least two of the baseline similar-size LMs.
-All the tasks are evaluated using 0-shot except MMLU, which is 5-shot.
-The models against which n/a is mentioned are trained from scratch.
-E.1 Main Results in Low Data Regime
-Table 6 presents a detailed performance evaluation across various tasks.
-Our 1.5B model, developed using $\mathsf{Inheritune}$, excels in 7 out of 10 individual tasks.
-It achieves a score of 90% or higher compared to the reference language model, which is twice its size and trained with 1000 times more data, or it outperforms at least two other base LMs of similar size trained with 50-300 times more data.
-Favorable scores are highlighted in bold.
-Next, we compare our small LM with the MPT-1.3B [^4] https://huggingface.co/mosaicml/mpt-1b-redpajama-200b model trained from scratch with 200B tokens of RedPajama dataset and find that we match 97% accuracy in all nine downstream tasks and the MMLU (5-shot) score.
-Additionally, we compare with OPT-1.3B and Pythia-1.3B models, showing that we outperform both in the MMLU (5-shot) score and perform comparably on the other nine datasets.
-This comparison illustrates that having a large reference base LM and a subset of its pre-training data allows the inherited target size base LM to be trained remarkably more sample-efficiently than training from scratch.
-Extended discussions on comparisons with the ShearedLLaMa model, generated by pruning and continual training from LLaMA2-7B, are provided in the supplementary materials.
-Ablation of Inheritune Across Different Model Sizes with 1B Tokens.
-In the previous section, we considered a single choice of $n=k/2$, i.e., half the layers, for the size of the smaller model.
-Here, we investigate $\mathsf{Inheritune}$ with different choices of $n$, but the same 1B token dataset).
-All models use OpenLLAMA-3B as the large pre-trained reference model, with consistent training hyperparameters, changing only the choice of $n$.
-We developed eight different submodels with $n=\{4,6,8,10,13,16,18,20\}$.
-Figure 20 shows the MMLU (5-shot) score as a function of $n$.
-As expected, the trend line is positive-sloping.
-The submodel with 20 layers slightly decreases performance, potentially due to data overfitting as the model size increases.
-The training details for all these submodels are consistent with the target 1.5B small base LM and are detailed in the appendix.
-A more comprehensive investigation on the choice of $n$—including varying both $n$ and the number of training tokens jointly and evaluating a broader set of tasks is left for future work.
-Figure 20: Inheritune scales across multiple different model sizes.
-Utilizing the OpenLLaMA-3B as a reference large base LM, demonstrates that multiple performant small base LMs of target size can be crafted using $\mathsf{Inheritune}$ with just 1B training tokens.
-The MMLU (5-shot) as a function of the number of submodels.
-E.2 Additional analysis with larger reference LMs and 50B data
-We further analyze $\mathsf{Inheritune}$ to see the impact of it’s performance when more tokens are available.
-Initially for the main results we limited ourselves to 1B (i.e.
-0.1%) tokens from the 1T pre-training data, here we use a 50B subset (i.e.
-5%) of the pre-train data.
-Moreover we also extend this study to include larger base LMs of 7B parameters as reference models, employing OpenLLaMA-7B and LLaMA2-7B as reference models.
-For the purpose of this study we do not repeat the tokens from our 50B subset.
-As shown in Figure 21, we observe that there is clear improvement in overall MMLU (5-shot) score with more data.
-Additionally it is interesting to see that 1.5B (or 1.6B models) developed with $\mathsf{Inheritune}$ using larger reference models show even greater improvements when fed with 50B subset of non repetitive data (i.e fresh tokens).
-We present a Table 9 using Figure 21 to show the best MMLU (5-shot) scores achieved using different reference LMs.
-For developing our small base LMs using larger reference LMs we use $n$=7 (i.e.
-7 layers).
-The training details are discussed in the following section.
-Model Training Data (# tokens)
-OpenLLaMA-3B v1(ref) RedPajama(1T)
-Ours-1.5B* RedPajama (1B)
-Shear-LLaMA-1.3B* RedPajama(50B)
-MPT-1.3B RedPajama(200B)
-Pythia-1.4B The Pile(300B)
-OPT-1.3B Custom data(300B)
-Table 7: Comparison of training data across baseline models.
-Overview of reference and baseline models, including their pre-training datasets and the number of tokens used during training.
-Note the significant variation in training data size, ranging from 1B to 1T tokens.
-Models (# train tokens) GPU Count GPU Type Time (# days)
-MPT-1.3B (200B) 440 A100 half
-Pythia-1.4B (300B) 64 A100 4.6
-TinyLLaMA-1.1B (3T) 16 A100 90
-OPT-1.3B (300B) 992 A100
-Sheared LLaMA-1.3B (50B) 16 A100
-OpenLLaMA-3B (1T) 256 TPU v4 10
-Our-1.5B (1B) 1 A6000 ~half
-Table 8: Computational efficiency of Inheritune versus baseline models.
-Comparison of pre-training compute requirements for publicly available small base LMs and our $\mathsf{Inheritune}$-derived model.
-Metrics include GPU count, GPU type, and training duration, highlighting Inheritune’s significant reduction in computational resources.
-Figure 21: Impact of reference model choice on Inheritune performance.
-MMLU (5-shot) scores for 1.5B base LMs derived using $\mathsf{Inheritune}$, trained on 50B unique tokens.
-Comparison across three reference models: OpenLLaMA-7B, LLaMA2-7B, and OpenLLaMA-3B.
-Results demonstrate Inheritune’s effectiveness with various large language models as references.
-Figure 22: Performance of our 1.5B base LM derived using $\mathsf{Inheritune}$ based on existing OpenLLaMA-3B base model.
-Here we use 1B tokens and perform data repetition (epochs) during training.
-We further evaluate our model on an average of 9 different datasets (left) and MMLU benchmark (right).
-Model (# tokens), ref MMLU(5)
-Ours-1.6B (1B), LLaMA2-7B 24.27
-Ours-1.5B (1B), OpenLLaMA-3B 25.67
-Ours-1.5B (50B), OpenLLaMA-3B 25.71
-Ours-1.6B (50B), LLaMA2-7B 26.07
-Ours-1.6B (50B), OpenLLaMA-7B 26.72
-Table 9: Performance comparison of models on the MMLU (5-shot) task.
-Our models, even when trained with fewer tokens, show competitive performance compared to benchmarks.
-We have highlighted the best MMLU 5-shot score in bold.
-Model (# tokens) Data type MMLU (5-shot)
-Ours-1.5B (1B) 10 epochs 24.95
-Ours-1.5B (50B) 10B fresh 23.62
-Ours-1.5B (1B) 20 epochs 25.46
-Ours-1.5B (50B) 20B fresh 24.96
-Table 10: MMLU (5-shot) scores of Our-1.5B small base LM derived using 1B data for multiple data repetition–10 epochs and 20 epochs compared to the same model trained without data repetition for 10B and 20B fresh tokens.
-We derive all the variants of Our-1.5B small base using $\mathsf{Inheritune}$ with OpenLLaMA-3B as reference model.
-The models featured in this table correspond to those discussed in Figures 21 and 22.
-Ablations with number of epochs.
-We ran ablations (refer Figure 22) to choose the total number of epochs (multiple passes over the data) and observe that repetition when training our 1.5B (or 1.6B) LM is helpful particularly for MMLU.
-We also observe that the for an average of all the 9 other datasets (i.e.
-except MMLU) peaks it’s performance at 5 epochs and then deteriorates.
-Some prior works have studied this phenomenon that the scaling of downstream tasks with data is not always linear Biderman et al.
-(2023).
-To repeat or not to repeat the tokens.
-Next we tackle the question – whether one should re-use 1B tokens for multiple epochs or use the same number of fresh tokens?
-Some prior works have recommended that if you have a reasonably large size dataset one can repeat it upto 4 epochs Muennighoff et al.
-(2023).
-In our study we observe that one can safely re-use 1B tokens upto 10-20 epochs as shown in Table 10.
-We emphasis that this phenomenon needs a through investigation in itself and we defer this to future work.
-The models discussed in Table are saved checkpoints during a single training run and not the final model unless otherwise specified.
-E.3 Implications of Low Data Regime
-In this section, we discuss some of the key implications of our work in low data regime.
-Cheap and easy development of small base LMs.
-Pre-training a small base LM of 1-2B parameters from scratch is extremely expensive.
-For instance mpt-1.3B base LM is pre-trained with 440 A100 GPUs for half a day, while the Pythia-1.4B base LM Biderman et al.
-(2023) utilized 64 A100-40GB GPUs for 4.6 days.
-Similarly, TinyLLaMA-1.1B model Peiyuan Zhang & Lu (2023) was pre-trained using 16 A100 GPUs for 3 months.
-Our 1.5B (1B data variant) LM shows competitive performance despite being trained with 1 A6000 GPU for less than 12 hours.
-The computational details are provided in Table 8, comparing the training resources of the baseline models listed in this paper.
-Typically small base LMs are finetuned for a specific task before deployment and are not used in it’s base form.
-With $\mathsf{Inheritune}$ we present a really easy and cheap way for developing a small base LM to be later finetuned before deployment.
-Naive baseline for pre-training a scaled down variant of large base LMs.
-Typically small variants of large base LMs are pre-trained using the same pre-training data Peiyuan Zhang & Lu (2023); Groeneveld et al.
-(2024).
-Our recipe introduces a new perspective of identifying sufficient depth without losing any generalization on the held out validation set.
-Next, we also show that even with a small fraction of pre-train data (randomly sampled) and few initial layers of the large base LM one can develop a small base LM.
-Therefore our $\mathsf{Inheritune}$ recipe has the potential to become the naive baseline for any pre-training pipeline aiming to develop a smaller variant of a large base LM.
-Appendix F Implementation Details
-F.1 Training details of GPT-2 models
-For our main experiments, we focused on three sizes of GPT-2 models Radford et al.
-(2019): the vanilla GPT-2 xlarge with 1.5B parameters, GPT-2 large with 770M parameters and the vanilla GPT-2 medium with 355M parameters.
-We developed several variants of these models by adjusting the number of layers and hidden size.
-We trained all GPT-2 models with data repetition while using OpenWebText dataset, the trainset has 9B tokens and the validation set has 4.4M tokens.
-The key architectural configurations for the reference models, our models, and baseline models discussed in this paper are summarized in Table 11.
-For all training runs, we used GELU activations, disabled bias terms, and removed dropout, following the nanoGPT codebase and Liu et al.
-(2023).
+For evaluating the pre-trained models, we used the validation split of the same dataset, which contains 4.4M tokens.
+The sole exception to this setup is the GPT-2 models trained on the FineWeb edu with 10B tokens (Figure 7), where we used unique tokens for training by employing a dataloader where we sample without replacement.
 We employed the AdamW optimizer with $\beta_{1}=0.90$ and $\beta_{2}=0.95$.
-The GPT-2 models were trained on a single node with 3 A100 GPUs (each with 40 GB of memory) using distributed data parallelism and gradient accumulation.
-In line with Liu et al.
-(2023), we scaled the attention logits inversely to the layer index across all GPT-2 models.
+All GPT-2 models were trained on a single NVIDIA A100 GPU (40 GB memory) with gradient accumulation.
+For the GPT-2 XLarge and its variants, we utilized an NVIDIA H100 GPU.
 Most hyperparameters were adapted from Liu et al.
-(2023), with key details provided below.
+(2023), with key details discussed in this section.
 Hyper-parameter details of GPT-2 Medium and variants.
-  • Batch size: 50K tokens
-  • Learning rate: $3\times 10^{-4}$,
-  • Warmup steps: 2K,
-  • Scheduler type: cosine decay to $\frac{1}{10}$ of max learning rate,
-  • Weight decay: 0.1,
-  • Gradient clipping value: 1,
+  • Batch size: 394K tokens
+  • Learning rate: $3\times 10^{-4}$
+  • Warmup steps: 2K
+  • Scheduler type: cosine decayed to $1\times 10^{-5}$
+  • Weight decay: 0.1
+  • Gradient clipping value: 1
   • Total training steps: 100K
 Hyper-parameter details of GPT-2 Large and variants.
-  • Batch size: 16K tokens
-  • Learning rate: $2\times 10^{-4}$,
-  • Warmup steps: 2K,
-  • Scheduler type: cosine decayed to $1\times 10^{-5}$,
-  • Weight decay: 0.1,
-  • Gradient clipping value: 1,
+  • Batch size: 128K tokens
+  • Learning rate: $2\times 10^{-4}$
+  • Warmup steps: 2K
+  • Scheduler type: cosine decayed to $1\times 10^{-5}$
+  • Weight decay: 0.1
+  • Gradient clipping value: 1
   • Total training steps: 100K
 Hyper-parameter details of GPT-2 XLarge and variants.
-  • Batch size: 16K tokens
-  • Learning rate: $1.5\times 10^{-4}$,
-  • Warmup steps: 2K,
-  • Scheduler type: cosine decayed to $1\times 10^{-5}$,
-  • Weight decay: 0.1,
-  • Gradient clipping value: 1,
+  • Batch size: 128K tokens
+  • Learning rate: $1.5\times 10^{-4}$
+  • Warmup steps: 2K
+  • Scheduler type: cosine decayed to $1\times 10^{-5}$
+  • Weight decay: 0.1
+  • Gradient clipping value: 1
   • Total training steps: 100K
@@
 We use the below loss for as our distillation based training loss. The validation loss is the student_loss.
+$\text{Total\_loss}=\alpha\cdot\text{student\_loss}+(1-\alpha)\cdot\text{distillation\_loss}$
+  • Model: 16-layer GPT-2 Medium variants
+  • $\alpha$: 0.6
+  • Batch size: 394K tokens
+  • Learning rate: $3\times 10^{-4}$
+  • Warmup steps: 2K
+  • Scheduler type: cosine decay to $\frac{1}{10}$ of max learning rate
+  • Weight decay: 0.1
+  • Gradient clipping value: 1
+  • Total training steps: 50K
+Appendix G Additional Experiments and Discussions
+G.1 Additional Downstream Evaluation
+We extend our evaluation to models trained on OpenWebText, as described in Section 4 (see Table 7).
+For this analysis, we focus on the largest model considered in this work, GPT-2 XLarge, along with its variants.
+Downstream evaluation is performed on all datasets listed in Section 4, with the addition of Winogrande (Sakaguchi et al., 2020), BoolQ (Clark et al., 2019), and Wikitext (Merity et al., 2016).
+Note that the LAMBADA dataset is evaluated under two settings: missing-word prediction (accuracy) and language modeling (perplexity), and is therefore reported twice.
+All evaluations are implemented using the widely adopted lm-eval-harness.
+Task Full model Ours
+Accuracy-based tasks ( $\uparrow$ )
+ARC-E 50.38 $\pm$ 1.03 51.22 $\pm$ 1.03
+PIQA 66.70 $\pm$ 1.10 66.87 $\pm$ 1.10
+SciQ 77.00 $\pm$ 1.33 79.20 $\pm$ 1.28
+HellaSwag 33.65 $\pm$ 0.47 34.20 $\pm$ 0.47
+LAMBADA 39.90 $\pm$ 0.68 43.30 $\pm$ 0.69
+WinoGrande 51.93 $\pm$ 1.40 53.28 $\pm$ 1.40
+BoolQ 57.86 $\pm$ 0.86 60.40 $\pm$ 0.86
+Average 53.92 55.50
+Perplexity-based tasks ( $\downarrow$ )
+Wikitext 25.46 25.52
+LAMBADA 20.24 16.51
+Average 22.85 21.01
+Table 7: Downstream evaluation of GPT-2 XLarge (1.5B) trained from scratch vs.
+a 24-layer model trained with Inheritune (Ours).
+We evaluate both models on 7 accuracy-based tasks (higher is better) and 2 perplexity-based tasks (lower is better).
+All models are trained on OpenWebText.
+Despite using half the depth, the $\mathsf{Inheritune}$ model performs favorably compared to the full-sized counterpart.
+Best average scores are highlighted in bold.
+G.2 Additional Evidence of Attention Collapse in Modern LLMs
+Following the discussion in Section B.1, we extend our analysis to several additional open-weight base LLMs (see Figures 21 and 22), including Falcon-7B [^3] https://huggingface.co/tiiuae/falcon-7b, OLMo-1B [^4] https://huggingface.co/amd/AMD-OLMo, Cerebras-GPT-2.7B [^5] https://huggingface.co/cerebras/Cerebras-GPT-2.7B, and LLaMA-3-3B [^6] https://huggingface.co/meta-llama/Llama-3.2-3B.
+Many of these models incorporate modern architectural components, including Grouped Query Attention (GQA), RoPE positional embeddings, and RMS normalization applied both before and after the attention module and trained with billions-trillions of tokens.
+We visualized the attention heads of each model using heatmaps of head index versus layer index.
+Across all models, we observe a predominance of near rank-1 attention heads, indicating widespread attention collapse.
+Figure 21: Rank analysis of the Falcon-7B model reveals widespread attention collapse.
+We analyzed Falcon-7B (version 1), which comprises 31 layers with 71 attention heads per layer, using the rank metric defined in Section 2.
+The results are visualized as a heatmap with layer index on one axis and head index on the other.
+Potent (non-collapsed) heads are shown in varying shades of green, with higher intensity indicating higher rank, while rank-collapsed (near rank-1) heads are highlighted in red.
+Overall, approximately 75% of attention heads exhibit rank collapse, indicating substantial degeneracy across the model.
+(a) OLMo 1B
+(b) Cerebras-GPT 2.7B
+(c) LLaMA-3 3B
+Figure 22: Rank analysis of large open-weight LLMs reveals widespread attention collapse.
+We analyze three base open-weight language models using the rank metric defined in Section 2.
+(a) OLMo (1B) exhibits rank collapse in approximately 58% of attention heads; (b) Cerebras-GPT (2.7B) shows a similar level of collapse at roughly 58%; and (c) LLaMA-3 (3B) exhibits rank collapse in about 50% of attention heads.
+Each heatmap plots attention head index versus layer index.
+Potent (non-collapsed) heads are shown in varying shades of green, with higher intensity indicating higher rank, while rank-collapsed (near rank-1) heads are highlighted in red.
+G.3 Additional Results on the Functional Ineffectiveness of Lazy Layers
+Following the discussion in Section 2.1 on the poor transferability of lazy layers, we conduct an additional experiment.
+We first train a 24-layer GPT-2 Medium model (reference) on OpenWebText for 100K steps.
+We then train a second model of identical size using the same data and training configuration.
+In this model, the deeper layers (layers 16–24) are initialized with the corresponding weights from the reference model, while the remaining layers are randomly initialized.
+As shown in Figure 23, reusing lazy layers consistently degrades performance relative to training all layers from scratch.
+This result complements our findings in Section 2.1 and provides further evidence that lazy layers do not encode transferable or reusable representations suitable for initialization.
+We used a batch size of 50K tokens with 2 H100 GPUs; all other training details remain unchanged (see Section F.2)
+G.4 Additional Training Results in the Extended Training Regime
+In this section we present a supplementary result where we have extended the training steps from 100K (as discussed in Section 4) to 200K steps to gauge the potential $\mathsf{Inheritune}$ holds for longer training runs.
+As shown in Figure 24 a GPT-2 medium (24 layer full model) is compared against a 18 layer GPT-2 medium variant trained following $\mathsf{Inheritune}$ recipe.
+We have inherited $l=18$ layers based on the previously known best configuration for GPT-2 medium variant where a 16 layer model matches the performance its full sized counterpart.
+We observe that the $\mathsf{Inheritune}$ model with 18 layers achieves a lower final validation loss (2.86) than the 24-layer baseline (2.87), demonstrating that $\mathsf{Inheritune}$ continues to provide benefits even with extended training.
+We used a batch size of 50K tokens with 2 H100 GPUs; all other training details remain unchanged (see Section F.2).
+Figure 23: Initializing deeper layers of GPT-2 Medium with lazy layers degrades performance.
+We first train a full GPT-2 Medium model from scratch on OpenWebText for 100K steps.
+We then construct a second model in which layers 16–24 are initialized using the corresponding lazy layers layers from the trained reference model, while layers 1–15 are randomly initialized.
+Both models are trained using the same optimization and data settings.
+The model whose deeper layers are initialized with lazy layers performs significantly worse than its counterpart trained entirely from scratch.
+Figure 24: Inheritune continues to provide gains under extended training.
+Comparison between a full 24-layer GPT-2 Medium model and an 18-layer GPT-2 Medium variant trained using the $\mathsf{Inheritune}$ recipe, both trained on OpenWebText for 200K steps.
+The $\mathsf{Inheritune}$ model achieves a lower final validation loss (2.86) than the full 24-layer baseline (2.87).
+G.5 Discussion on Lazy Layers
+In this section, we provide a formal characterization of lazy layers through the rank-based criterion.
+Definition G.1 (Lazy Layer (Rank-based)).
+Let $\{X_{n}\}_{n=1}^{N}$ be a collection of $N$ sequences, each consisting of $T$ tokens.
+For each layer $l\in\{1,\dots,L\}$ and head $h\in\{1,\dots,H\}$, let $A_{n,h,l}\in\mathbb{R}^{T\times T}$ denote the attention matrix produced by head $h$ in layer $l$ on input sequence $X_{n}$.
+Let $\sigma_{1}(A_{n,h,l})\geq\cdots\geq\sigma_{T}(A_{n,h,l})\geq 0$ be the singular values of $A_{n,h,l}$. For a threshold $\tau\in(0,1)$, we define the approximate rank as
+$k^{*}_{n,h,l}(\tau)\;=\;\min\left\{k\in\{1,\dots,T\}\;\middle|\;\frac{\sum_{i=1}^{k}\sigma_{i}(A_{n,h,l})^{2}}{\sum_{j=1}^{T}\sigma_{j}(A_{n,h,l})^{2}}\geq\tau\right\}.$ (2)
+We choose $\tau$ via an ablation study; throughout the paper we use $\tau=0.90$ (see Section 4.3).
+Define the head-wise aggregated rank and layer-wise aggregated rank as
+$\mathrm{Rank}(h,l)\;=\;\frac{1}{N}\sum_{n=1}^{N}k^{*}_{n,h,l}(\tau),\qquad\mathrm{MaxRank}(l)\;=\;\max_{h\in\{1,\dots,H\}}\,\mathrm{Rank}(h,l).$ (3)
+For a group of $L$ layers (where $L$ denotes the number of contiguous layers in the group), define
+$\mathrm{AvgRank}\;=\;\frac{1}{L}\sum_{l=1}^{L}\mathrm{MaxRank}(l).$ (4)
+Lazy Layer. A layer $l$ is a Lazy Layer if $\mathrm{MaxRank}(l)=1$. Lazy Layers. A group of contiguous blocks of $L$ layers is termed Lazy if $\lfloor\mathrm{AvgRank}\rfloor=1$.
+Appendix H Theoretical Analysis
+H.1 Causal Attention Rank-1 Collapse and Vanishing Gradients
+We isolate a simple mechanism in decoder-style (causal) self-attention: when a head becomes highly concentrated on a single sink position across time (e.g., the first/BOS token), the resulting attention matrix is approximately rank-$1$, and the corresponding gradients through the softmax saturate, yielding vanishing updates to $W_{Q}$ and $W_{K}$.
+Setup.
+We inherit notations discussed in Section 2. Let $X\in\mathbb{R}^{T\times d}$ denote a sequence of $T$ tokens. Define $Q=XW_{Q},\;K=XW_{K},\;V=XW_{V}\in\mathbb{R}^{T\times d}$ and masked logits
+$\tilde{A}=\frac{1}{\sqrt{d}}QK^{\top}+M,\qquad M_{ij}=\begin{cases}0&i\geq j,\\ -\infty&i<j,\end{cases}$ (5)
 […]
-  • Model: 16-layer and 12-layer GPT-2 medium variants
-  • Softmax temperature: 1
-  • $\alpha$: 0.6
-  • Batch size: 50K tokens
-  • Learning rate: $3\times 10^{-4}$,
-  • Warmup steps: 2K,
-  • Scheduler type: cosine decay to $\frac{1}{10}$ of max learning rate,
-  • Weight decay: 0.1,
-  • Gradient clipping value: 1,
-  • Total training steps: 50K
-Models Layers Hidden Size Heads Variant
+Definition (attention collapse).
+Let $A\in\mathbb{R}^{T\times T}$ be an attention matrix (row-stochastic: $A_{ij}\geq 0$ and $\sum_{j=1}^{T}A_{ij}=1$ for each row $i$).
+Fix an index $j^{\star}\in\{1,\ldots,T\}$ that is unmasked for all rows (e.g., $j^{\star}=1$ for the BOS token).
+We say that $A$ is $\varepsilon$-sink-collapsed to $j^{\star}$ if, for every row $i\in\{1,\ldots,T\}$,
+$\sum_{j\neq j^{\star}}A_{ij}\;\leq\;\varepsilon.$ (6)
+Equivalently, each row places at least $1-\varepsilon$ probability mass on the same column $j^{\star}$, i.e.,
+$A_{ij^{\star}}\;=\;1-\sum_{j\neq j^{\star}}A_{ij}\;\geq\;1-\varepsilon,\qquad\forall i\in\{1,\ldots,T\}.$ (7)
+Theorem H.1 (Attention collapse to rank-1 sink and vanishing gradients).
+Consider a single-head causal self-attention module with $A=\mathrm{softmax}(\frac{1}{\sqrt{d}}QK^{\top}+M)$ and $O=AV$.
+If $A$ is $\varepsilon$-sink-collapsed to some $j^{\star}$ as in equation 6, then:
+(i) Rank-1 collapse. Let $A_{0}\triangleq\mathbf{1}e_{j^{\star}}^{\top}$ (rank-11). Then
+$\|A-A_{0}\|_{F}\leq\varepsilon\sqrt{2T}\;\Rightarrow\;\sigma_{2}(A)\leq\|A-A_{0}\|_{2}\leq\|A-A_{0}\|_{F}\leq\varepsilon\sqrt{2T},$ (8)
+so $A$ is numerically rank-$1$ where $\varepsilon>0$ is small.
+(ii) Vanishing gradients through softmax. For each row $i$, letting $a^{(i)}$ be the entries of that row (restricted to unmasked indices),
+$\frac{\partial\mathcal{L}}{\partial\tilde{a}^{(i)}}=\Bigl(\mathrm{diag}(a^{(i)})-a^{(i)}{a^{(i)}}^{\top}\Bigr)\;\frac{\partial\mathcal{L}}{\partial a^{(i)}},\quad\Bigl\|\frac{\partial\mathcal{L}}{\partial\tilde{a}^{(i)}}\Bigr\|_{2}\leq 2\varepsilon\;\Bigl\|\frac{\partial\mathcal{L}}{\partial a^{(i)}}\Bigr\|_{2}.$ (9)
+Thus, once rows are near one-hot ($\varepsilon\!\to\!0$), the gradient to logits vanishes.
+(iii) Vanishing gradients to $W_{Q}$ and $W_{K}$. Moreover,
+$\Bigl\|\frac{\partial\mathcal{L}}{\partial W_{Q}}\Bigr\|_{F}\;\leq\;\frac{2\varepsilon}{\sqrt{d}}\;\|X\|_{2}\;\|K\|_{2}\;\Bigl\|\frac{\partial\mathcal{L}}{\partial A}\Bigr\|_{F},\quad\Bigl\|\frac{\partial\mathcal{L}}{\partial W_{K}}\Bigr\|_{F}\;\leq\;\frac{2\varepsilon}{\sqrt{d}}\;\|X\|_{2}\;\|Q\|_{2}\;\Bigl\|\frac{\partial\mathcal{L}}{\partial A}\Bigr\|_{F},$ (10)
+so parameter updates to queries/keys shrink linearly with the collapse level $\varepsilon$.
+Proof.
+We prove the three claims sequentially.
+(i) Let $A_{0}\;=\;\mathbf{1}e_{j^{\star}}^{\top}\in\mathbb{R}^{T\times T}$.
+Clearly $A_{0}$ has rank $1$.
+As the Frobenius norm is the $\ell_{2}$ norm aggregated over rows, it follows that $\|A-A_{0}\|_{F}^{2}\;=\;\sum_{i=1}^{T}\bigl\|a^{(i)}-e_{j^{\star}}\bigr\|_{2}^{2}$.
+Fix a row $a\in\mathbb{R}^{T}$ (without loss of generality we omit the superscript $i$), under $\varepsilon$-sink-collapse, $\sum_{j\neq j^{\star}}a_{j}\leq\varepsilon$.
+Since $A$ is an attention matrix (row-stochastic), then $\sum_{j=1}^{T}a_{j}=1$ and $a_{j}\geq 0$, hence $a_{j^{\star}}\;=\;1-\sum_{j\neq j^{\star}}a_{j}\;\geq\;1-\varepsilon.$ Now consider the squared error $\|a-e_{j^{\star}}\|_{2}^{2}\;=\;(a_{j^{\star}}-1)^{2}+\sum_{j\neq j^{\star}}a_{j}^{2}.$ Here, $(a_{j^{\star}}-1)^{2}\;\leq\;\varepsilon^{2}$ (follows from $a_{j^{\star}}\geq 1-\varepsilon$), and $\sum_{j\neq j^{\star}}a_{j}^{2}\;\leq\;\Bigl(\sum_{j\neq j^{\star}}a_{j}\Bigr)^{2}\;\leq\;\varepsilon^{2}$.
+Thus, the squared error $\|a-e_{j^{\star}}\|_{2}^{2}\leq 2\varepsilon^{2}$.
+Summing over $T$ rows, gives
 […]
-GPT2-large(770M) 36 1280 20 Original
-GPT2-large(680M) 32 1280 20 Original
-GPT2-medium(355M) 24 1024 16 Original
-[…]
-GPT2-medium 16 512 8 half width
-[…]
-GPT2-large 18 1280 20 Ours
-GPT2-large 16 1280 20 Ours
-GPT2-medium 16 1024 16 Ours
-Table 11: Overview of all the GPT2 models used in this study and their architectural configurations. The model configurations of stacking and hybrid stacking are same as our variants.
-F.2 Training details of 1.5B OpenLLaMA model
-Small base LMs trained with 1B data
-We present our main results with Our-1.5B model trained with an existing OpenLLaMA version 1 Geng & Liu (2023) and 1 B tokens randomly sampled from 1T redpajama version1 data.
-The hyper-parameters related to this model is provided below.
-It is important to note that our claim that we only use 1 GPU for less than 12 hours to train Our-1.5 B model is specific to models derived using $\mathsf{Inheritune}$ with 1B data.
-Next we also train multiple sub-models as shown in Figure 20 the training details remains consistent with that of the initial model discussed earlier.
-However we observe that increasing the number of layers in a sub-model also increase the training time.
-Hyper-parameter details of our 1.5B base LM derived using OpenLLaMA-3B as refernce LM:
-  • Training tokens: 1B
-  • Training epochs: 8
-  • Training steps: 64K
-  • Learning rate: $3\times 10^{-4}$
-  • Scheduler: Cosine
-  • Weight decay: 0.1
-  • Optimizer: AdamW
-  • Warm up steps: 1000
-  • Batch size: 131K
-  • GPU count: 1
-  • GPU type: A6000
-  • GPU hours: $\sim$ 8 hours
-  • GPU hours/epoch: $\sim$ 54 minutes
-Training details of small base LMs with 50B data.
-We also trained our 1.5B model with larger subsets of data as shown in Figure 21.It is important to note that all the intermediate tokens until 50B are intermediate checkpoints of a single training run.
-Some of the key hyper-parameters of our training runs are discussed below.
-We have also trained three variants of small base LMs utilizing 3 different reference base LMs namely OpenLLaMA-3B, OpenLLaMA-7B and LLaMA2-7B.
-For target LMs developed with OpenLLaMA-3B we use $n$=13 i.e.
-13 layers.
-For target LMs developed using reference LMs of 7B parameters we use $n$=7 i.e.
-7 layers.
-The training hyper-parameters remains consistent across all the models trained with 50B subset of the pre-train data.
-Training hyper-parameters of our target 1.5B and 1.6B small base LMs:
-  • Training tokens: 50B
-  • Training epochs: $\sim$1
-  • Training steps: 191K
-  • Learning rate: $3\times 10^{-4}$
-  • Scheduler: Cosine
-  • Weight decay: 0.1
-  • Optimizer: AdamW
-  • Warm-up steps: 1000
-  • Batch size: 131K tokens
-  • GPU count: 1
-  • GPU type: A100
-  • GPU hours: $\sim$18 hours
+$\text{Hence,}\;\sigma_{2}(A)\leq\|A-A_{0}\|_{2}\leq\|A-A_{0}\|_{F}\leq\varepsilon\sqrt{2T}\quad\text{(following \emph{Weyl's inequality})}.$
+(ii) (Gradients are taken only over unmasked logits; masked positions have zero gradient.) Let $g\triangleq\partial\mathcal{L}/\partial a^{(i)}$ and $\tilde{g}\triangleq\partial\mathcal{L}/\partial\tilde{a}^{(i)}$.
+For the row-wise softmax, the Jacobian is $J=\mathrm{diag}(a)-aa^{\top}$, hence $\tilde{g}=Jg$.
+Since $J\succeq 0$, $\|J\|_{2}\leq\mathrm{tr}(J)$.
+Moreover, $\mathrm{tr}(J)=\sum\limits_{i=1}^{T}a_{i}(1-a_{i})=1-\|a\|_{2}^{2}$.
+Under $\varepsilon$-sink collapse, $a_{j^{\star}}\geq 1-\varepsilon$, so $\|a\|_{2}^{2}\geq(1-\varepsilon)^{2}$ and therefore
+$\|J\|_{2}\leq 1-(1-\varepsilon)^{2}\leq 2\varepsilon.$
+Thus $\|\tilde{g}\|_{2}\leq\|J\|_{2}\|g\|_{2}\leq 2\varepsilon\|g\|_{2}$, proving equation 9.
+(iii) We focus on $W_{Q}$; the argument for $W_{K}$ is analogous.
+By the chain rule, $\frac{\partial\mathcal{L}}{\partial W_{Q}}=X^{\top}\frac{\partial\mathcal{L}}{\partial Q}$.
+Let $\bar{G}\triangleq\frac{\partial\mathcal{L}}{\partial\tilde{A}}$.
+Since $\tilde{A}=\frac{1}{\sqrt{d}}QK^{\top}+M$, we have $\frac{\partial\mathcal{L}}{\partial Q}=\frac{1}{\sqrt{d}}\bar{G}K$.
+Bounding norms and using $\|X^{\top}H\|_{F}\leq\|X\|_{2}\|H\|_{F}$, (by definition)
+$\Bigl\|\frac{\partial\mathcal{L}}{\partial W_{Q}}\Bigr\|_{F}\leq\|X\|_{2}\Bigl\|\frac{\partial\mathcal{L}}{\partial Q}\Bigr\|_{F}\leq\frac{1}{\sqrt{d}}\|X\|_{2}\|\bar{G}\|_{F}\|K\|_{2}.$
+From (ii), row-wise $\|\partial\mathcal{L}/\partial\tilde{a}^{(i)}\|_{2}\leq 2\varepsilon\|\partial\mathcal{L}/\partial a^{(i)}\|_{2}$, and therefore $\|\bar{G}\|_{F}\leq 2\varepsilon\|\frac{\partial\mathcal{L}}{\partial A}\|_{F}$.
+Substituting gives equation 10.
+∎
+Interpretation and training-time implication.
+Theorem H.1 connects rank-1 causal attention collapse (sink behavior) and vanishing gradients: once a head becomes near-deterministic, the softmax Jacobian saturates and the head receives negligible learning signal to recover.
+Consequently, collapse becomes sticky (hard to recover from) after entering the collapsed regime, subsequent gradient updates tend to be too small to move the head back to a higher rank (non-collapsed) state.
+Over training, this stickiness can accumulate across layers especially in deeper layers which are more prone to vanishing gradients leading to increasingly persistent rank collapsed attention patterns and progressively smaller query/key weight updates.